### v1

# Random expansions of trees with bounded height

Abstract.
We consider a sequence $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ of trees $\mathcal{T}_{n}$ where, for some $\Delta\in\mathbb{N}^{+}$ every $\mathcal{T}_{n}$ has height at most $\Delta$ and as $n\to\infty$ the minimal number of children of a nonleaf tends to infinity. We can view every tree as a (first-order) $\tau$-structure where $\tau$ is a signature with one binary relation symbol. For a fixed (arbitrary) finite and relational signature $\sigma\supseteq\tau$ we consider the set $\mathbf{W}_{n}$ of expansions of $\mathcal{T}_{n}$ to $\sigma$ and a probability distribution $\mathbb{P}_{n}$ on $\mathbf{W}_{n}$ which is determined by a (parametrized/lifted) Probabilistic Graphical Model (PGM) $\mathbb{G}$ which can use the information given by $\mathcal{T}_{n}$.
The kind of PGM that we consider uses formulas of a many-valued logic that we call $PLA^{*}$ with truth values in the unit interval $[0,1]$. We also use $PLA^{*}$ to express queries, or events, on $\mathbf{W}_{n}$. With this setup we prove that, under some assumptions on $\mathbf{T}$, $\mathbb{G}$, and a (possibly quite complex) formula $\varphi(x_{1},\ldots,x_{k})$ of $PLA^{*}$, as $n\to\infty$, if $a_{1},\ldots,a_{k}$ are vertices of the tree $\mathcal{T}_{n}$ then the value of $\varphi(a_{1},\ldots,a_{k})$ will, with high probability, be almost the same as the value of $\psi(a_{1},\ldots,a_{k})$, where $\psi(x_{1},\ldots,x_{k})$ is a “simple” formula the value of which can always be computed quickly (without reference to $n$), and $\psi$ itself can be found by using only the information that defines $\mathbf{T}$, $\mathbb{G}$ and $\varphi$. A corollary of this, subject to the same conditions, is a probabilistic convergence law for $PLA^{*}$-formulas.
1. Introduction
When building a model for probabilistic inference we may not only want to take into account properties and relations which are uncertain and described by (conditional) probabilities, but also “background information” that is certain. A common kind of background information is information about a hierarchical structure, and such information can be described by a tree. As an example of a tree that represents a hierarchical structure, consider a tree in which the children of the root represent some communities, where the children of a community represent subcommunities, and the children of a subcommunity represent individuals. The height of this hierarchy, or tree, is 3 (or 4, depending on our conventions). But the number of individuals in even the smallest subcommunity that we consider may be much larger than 3. This article considers hierarchical background information represented by a tree such that the height of the tree is very much smaller than the minimal number of children of a nonleaf of the tree.
On top of the background information, represented by the example tree described above, we may consider information that is given by (conditional) probabilities and is formally represented by (parametrized) 0/1-valued (i.e. false/true-valued) random variables. So we may have random variables (also viewed as atomic logical formulas) $P_{1}(x)$, $P_{2}(x)$, and $P_{3}(x)$, representing some property that a community, subcommunity, respectively, individual, may have. The probability that a subcommunity has the property $P_{2}$ may (for example) depend on whether the community which it is part of has property $P_{1}$. Or alternatively, the probability that a community has $P_{1}$ may depend on the proportion of its subcommunities that have $P_{2}$ (and/or the proportion of individuals in the community that have $P_{3}$). We may also have a random variable $R(x,y)$, representing some relationship between objects, for example the perception (good/bad) that $x$ has of $y$. Now, for an individual $x$, the probability that $R(x,y)$ holds may depend on properties of the (sub)community that $x$ belongs to and on properties of $y$ and/or its (sub)community; it may also depend on whether $x$ and $y$ belong to the same (sub)community.
Such uncertain, or probabilistic, information as exemplified above can be represented by a so-called (parametrized/lifted) Probabilistic Graphical Model (PGM) based on the background knowledge represented by the tree. (For introductions or surveys about PGMs, a tool in machine learning and statistical relational artifical intelligence, see for example [7, 11, 17, 18].) A (parametrized/lifted) PGM consists partly of a directed acyclic graph (DAG), the vertices of which are (parametrized/lifted) random variables. The arcs/arrows of the DAG describes the (in)dependencies between the random variables, by stipulating that a random variable (a vertex of the DAG), say $R$, is independent from all other vertices of the DAG, conditioned on knowing the states (true/false) of the parents of $R$ in the DAG. In the setup of this article a parametrized random variable will be identified with a relation symbol in the sense of formal logic. So in the example above the vertex set of the DAG would be $\{P_{1},P_{2},P_{3},R\}$, where $P_{1},P_{2},P_{3},R$ are relation symbols, and its arcs would describe the (conditional) (in)dependencies between the vertices.
So far, nothing has been said about how to express (conditional) probabilities in a PGM and, indeed, there are different ways of doing it. It turns out that a general way of doing it is to use some sort of “probability logic”, as was first done (as far as we know) in the contex of PGMs by Jaeger in [14]. Here we will use a logic which we call $PLA^{*}$, for probability logic with aggregation functions (Definition 3.4) to describe (conditional) probabilities associated to the relation symbols of the DAG of a PGM. (The ‘*’ in $PLA^{*}$ just indicates that it is a more general variant of earlier versions of $PLA$ considered by Koponen and Weitkämper in [21, 22].) The resulting type of PGM will be called a $PLA^{*}$-network (Definition 6.8).
$PLA^{*}$ is a logic such that the (truth) values of its formulas can be any number in the unit interval $[0,1]$, so it is suitable for expressing probabilities. $PLA^{*}$ achieves this partly by considering every $c\in[0,1]$ as an atomic formula, and partly by using aggregation functions which take finite sequences of reals in $[0,1]$ as input and outputs a real number in $[0,1]$. An example of an aggregation function is the arithmetic mean, or average, which returns the average of a finite sequence of reals in the unit interval. So for example, if $R$ is a unary relation symbol then there is a $PLA^{*}$-formula (without free variables) such that its value is the proportion of elements in the domain which satisfy $R(x)$. Since the aggregation functions maximum and minimum are available there is a $PLA^{*}$ formula which expresses that at least one (alternatively all) element in the domain satisfies $R(x)$. By using the aggregation functions maximum and minimum one can show that every query on finite structures which can be expressed by a first-order formula can be expressed by a $PLA^{*}$-formula (Lemma 3.13). Hence, $PLA^{*}$ subsumes first-order logic. For a formula $\varphi(x_{1},\ldots,x_{k})$ of $PLA^{*}$, finite structure $\mathcal{A}$ (with matching signature/vocabulary), and $a_{1},\ldots,a_{k}\in A$, we let $\mathcal{A}(\varphi(a_{1},\ldots,a_{k}))$ denote the value of $\varphi(a_{1},\ldots,a_{k})$ in $\mathcal{A}$.
The precise way that we will mathematically represent the kind of probabilistic (or uncertain) information described above together with some background information, in the form of a tree, is as follows. We let $\Delta$ be a fixed positive integer and we let $\mathcal{T}_{n}$, where $n$ is a positive integer, be a tree with maximum height $\Delta$ that represents the hierarchical background information. Since we are interested in what happens for (arbitrarily) large trees where nonleaves have many children, we assume that, for every positive integer $k$, if $n$ is large enough then every nonleaf in $\mathcal{T}_{n}$ has at least $k$ children. Each $\mathcal{T}_{n}$ can be represented by a first-order $\tau$-structure where $\tau=\{E\}$ and $E$ is a binary relation symbol of arity 2. Then let $\sigma$ be a finite relational signature that includes $\tau$. In the example above we have $\sigma=\{E,P_{1},P_{2},P_{3},R\}$. Then let $\mathbf{W}_{n}$ be the set of all expansions of $\mathcal{T}_{n}$ to $\sigma$, where we think of each such expansion $\mathcal{A}$ as a “possible world”, and let $\mathbb{P}_{n}$ be the probability distribution on $\mathbf{W}_{n}$ induced by a $PLA^{*}(\sigma)$-network $\mathbb{G}$ as defined in Definition 6.9 .
So given a sequence $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ of trees (of the described kind) and a $PLA^{*}(\sigma)$-network $\mathbb{G}$, we have a precise mathematical setting of computing the probability of a query (event) on $\mathbf{W}_{n}$ that can be expressed by a $PLA^{*}$-formula. The probability of a query depends on both on the hierarchical information given by $\mathcal{T}_{n}$ and by $\mathbb{G}$. However, if we use the brute force way of computing the probability of a query, namely that we compute the probability of every structure in $\mathbf{W}_{n}$ in which the query is true and then add all such probabilities, then the time needed for the computation will be exponential in the number of vertices of the tree $\mathcal{T}_{n}$. In other words, the brute force method of computing the probability of a query does not scale well to large trees $\mathcal{T}_{n}$. This motivates the search for other methods of computing, or estimating, the probability of a query in the described context.
In this article we identify
  1. (1)
    certain conditions on the sequence of trees $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ (the stronger Assumption 6.3 and the milder Assumption 8.3),
  2. (2)
    certain conditions on the $PLA^{*}(\sigma)$-network $\mathbb{G}$ used for defining a probability distribution on $\mathbf{W}_{n}$, roughly meaning that all aggregation functions used in the formulas associated to $\mathbb{G}$ satisfy a continuity property,
such that if
  • (3)
    $\varphi(x_{1},\ldots,x_{k})$ (with free variables $x_{1},\ldots,x_{k}$) is a $PLA^{*}(\sigma)$-formula in which all aggregation functions satisfy the continuity property,
then there is a simpler, “closure-basic”, formula, say $\psi(x_{1},\ldots,x_{k})$, such that, for all $n$ and $a_{1},\ldots,a_{k}\in T_{n}$,
  • (a)
    the truth value of $\psi(a_{1},\ldots,a_{k})$ in every structure $\mathcal{A}\in\mathbf{W}_{n}$ can be evaluated by using only (the fixed information defining) $\mathbf{T}$, $\mathbb{G}$, $\psi$, and the (bounded number of) ancestors of $a_{1},\ldots,a_{k}$ in $\mathcal{T}_{n}$, and
  • (b)
    for every $\varepsilon>0$, if $n$ is large enough, then, with probability at least $1-\varepsilon$ (with the distribution induced by $\mathbb{G}$), the values of $\varphi(a_{1},\ldots,a_{k})$ and $\psi(a_{1},\ldots,a_{k})$ differ by at most $\varepsilon$. (We will say that $\varphi$ and $\psi$ are asymptotically equivalent.)
This intuitively means that, for large enough $n$, with high probability the value of $\varphi(a_{1},\ldots,a_{k})$ is approximated (with as good accuracy as we like) by the value of
$\psi(a_{1},\ldots,a_{k})$ for all vertices $a_{1},\ldots,a_{k}$ in the tree $\mathcal{T}_{n}$. Moreover, under the same conditions (1)–(3), one can find such $\psi$ by using only $\varphi$, $\mathbf{T}$ and $\mathbb{G}$, so the time needed to compute the approximation is independent of $n$, that is, of the size of the tree $\mathcal{T}_{n}$. A corollary of the above result is that (under the same conditions) we get a convergence law for $PLA^{*}$-formulas, which of course has to be suitably formulated in the context of many-valued formulas and trees as “background”, or “base”, structures. All three conditions (1)–(3) are necessary for the conclusions described to hold, as we will show.
The main results, stated in Section 8.1, have the general form described above. We first prove Theorem 8.1. Then we make some observations which allow us to easily obtain some corollaries which are variations of Theorem 8.1. The essence of these corollaries is (i) that if we consider stronger assumptions on the $PLA^{*}(\sigma)$-network $\mathbb{G}$ (i.e. less general probability distributions), then we get results that apply to more queries, and (ii) if we loosen the assumptions on the sequence of trees $\mathbf{T}$ then we get results that apply to less, but still interesting, queries.
Related work. Instead of bounding the height of the trees, as we do in this article, one could consider trees with a bound on the number of children that vertices may have. Such sequences of trees are covered by the (more general) context considered by Koponen in [20], where probability distributions where defined by $PLA^{*}(\sigma)$-networks and queries by $PLA^{*}$-formulas. Other work that we are aware of on logical convergence laws, and related issues, in the setting of expansions of “base structures” uses first-order logic (or $L_{\infty,\omega}^{\omega}$) as the query language. Baldwin [4] and Shelah [27] considered a base sequence where each $\mathcal{B}_{n}$ is a directed path of length $n$. In both cases $\mathbf{W}_{n}$ consists of all expansions of $\mathcal{B}_{n}$ to a new binary relation symbol $R$ interpreted as an irreflexive and symmetric relation. In [4] the probability distribution on $\mathbf{W}_{n}$ is defined by letting an $R(x,y)$ be true with probability $n^{-\alpha}$ for some irrational $\alpha\in(0,1)$, independently of whether it is true for other pairs. In [27] the probability of $R(x,y)$ is $d^{-\alpha}$ where $d$ is the distance in $\mathcal{B}_{n}$ between $x$ and $y$ and $\alpha\in(0,1)$ is irrational. Lynch [25, Corollary 2.16] and later Abu Zaid, Dawar, Grädel and Pakusa [1] and Dawar, Grädel and Hoelzel [10] considered the context where $\mathcal{B}_{n}$ is a product of finite cyclic groups and the probability distribution considered is the uniform one. Ahlman and Koponen [2] considered base structures $\mathcal{B}_{n}$ with a (nicely behaved) pregeometry and a probability distribution defined by a kind of stochastic block model where the “blocks” correspond to subspaces. Lynch [25] formulated a condition ($k$-extendibility) that makes sense for base sequences $\mathbf{B}$ in general and which guarantees that a convergence law holds for first-order logic and the uniform probability distribution.
Besides this article and [20], other studies that use a a PGM for generating a probability distribution and that prove some result about asymptic probabilities do not consider background information in the form of a sequence of nontrivial base structures. But in the context of no background information, when we let $\sigma$ be a finite and relational signature, $\mathbf{W}_{n}$ the set of all $\sigma$-structures with domain $\{1,\ldots,n\}$, $\mathbb{P}_{n}$ a probability distribution induced by a PGM, there are some studies. The most closely related to this one are [21, 22, 23] by Koponen and Weitkämper and [14] by Jaeger. But we also have [19] by Koponen, [28, 29] by Weitkämper, and [9] by Cozman and Maua. The first result about “asymptotic equivalence” between complex first-order formulas and simpler formulas was probably given by Glebskii et. al. [13] and was used to prove the classical zero-one law for first-order logic, also proved independently by somewhat different methods by Fagin [12]. Since the pioneering work in [13] and [12] a large number of studies on logical convergence laws and related issues have been conducted in the fields of finite model theory and probabilistic combinatorics, but those studes consider less flexible ways of generating probability distributions than PGMs (most often the uniform distribution).
Organization. Section 2 specifies some notation and terminology that will be used, as well as a couple of probability theoretic results. Section 3 defines the formal logic $PLA^{*}$, and some related notions, that will be used in PGMs and for defining queries (events). Section 4 describes a general method (introduced in [23]) for “asymptotically eliminating aggregation functions”. Section 5 gives the necessary background about directed acyclic graphs and trees, and some other notions related to trees, that will be used in the rest of the article. Section 6 specifies the assumptions on the sequence of trees $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ that we make and also explains why we make these assumptions. It then defines the notion of a $PLA^{*}$-network, which is the kind of PGM that we use, and how it induces a probability distribution on the set $\mathbf{W}_{n}$ of all expansions to $\sigma$ of $\mathcal{T}_{n}$. In section 7 we prove technical results that imply that the assumptions of Theorem 4.8 in Section 4 are satisfied. Therefore we can, in the last subsection of Section 7, use Theorem 4.8 and the results from earlier subsections of Section 7 to prove a result about “asymptotic elimination of aggregation functions”. In Section 8 we use induction on the height of the $PLA^{*}$-network and the results from Section 7 to prove our main results.
2. Preliminaries
Structures in the sense of first-order logic are denoted by calligraphic letters $\mathcal{A},\mathcal{B},\mathcal{C},\ldots$ and their domains (universes) by the corresponding noncalligraphic letter $A,B,C,\ldots$. Finite sequences (tuples) of objects are denoted by $\bar{a},\bar{b},\ldots,\bar{x},\bar{y},\ldots$. We usually denote logical variables by $x,y,z,u,v,w$. Unless stated otherwise, when $\bar{x}$ is a sequence of variables we assume that $\bar{x}$ does not repeat a variable. But if $\bar{a}$ denotes a sequence of elements from the domain of a structure then repetitions may occur.
We let $\mathbb{N}$ and $\mathbb{N}^{+}$ denote the set of nonnegative integers and the set of positive integers, respectively. For a set $S$, $|S|$ denotes its cardinality, and for a finite sequence $\bar{s}$, $|\bar{s}|$ denotes its length and $\mathrm{rng}(\bar{s})$ denotes the set of elements in $\bar{s}$. For a set $S$, $S^{<\omega}$ denotes the set of finite nonempty sequences (where repetitions are allowed) of elements from $S$, so $S^{<\omega}=\bigcup_{n\in\mathbb{N}^{+}}S^{n}$. In particular, $[0,1]^{<\omega}$ denotes the set of all finite nonempty sequences of reals from the unit interval $[0,1]$.
A signature (vocabulary) is called finite (and) relational if it is finite and contains only relation symbols. Let $\sigma$ be a signature and let $\mathcal{A}$ be a $\sigma$-structure. If $\tau\subseteq\sigma$ then $\mathcal{A}\negthickspace\upharpoonright\negthickspace\tau$ denotes the reduct of $\mathcal{A}$ to $\tau$. If $B\subseteq A$ then $\mathcal{A}\negthickspace\upharpoonright\negthickspace B$ denotes the substructure of $\mathcal{A}$ generated by $B$. If $R\in\sigma$ is a relation symbol then $R^{\mathcal{A}}$ denotes the interpretation of $R$ in $\mathcal{A}$.
A random variable will be called binary if it can only take the value 00 or $1$. The following is a direct consequence of [3, Corollary A.1.14] which in turn follows from the Chernoff bound [8]:
Lemma 2.1.
Let $Z$ be the sum of $n$ independent binary random variables, each one with probability $p$ of having the value 1, where $p>0$. For every $\varepsilon>0$ there is $c_{\varepsilon}>0$, depending only on $\varepsilon$, such that the probability that $|Z-pn|>\varepsilon pn$ is less than $2e^{-c_{\varepsilon}pn}$. (If $p=0$ then the same statement holds if ‘$2e^{-c_{\varepsilon}pn}$’ is replaced by (for example) ‘$e^{-n}$’.)
The following is a straightforward corollary (proved in [22]):
Corollary 2.2.
Let $p\in[0,1]$ and let $\varepsilon>0$. Let $Z$ be the sum of $n$ independent binary random variables $Z_{1},\ldots,Z_{n}$, where for each $i=1,\ldots,n$ the probability that $Z_{i}$ equals 1 belongs to the interval $[p-\varepsilon,p+\varepsilon]$. Then there is $c>0$, depending only on $p$ and $\varepsilon$, such that the probability that $Z>(1+\varepsilon)(p+\varepsilon)n$ or $Z<(1-\varepsilon)(p-\varepsilon)n$ is less than $2e^{-cn}$.
The following lemma follows easily from the definition of conditional probability.
Lemma 2.3.
Suppose that $\mathbb{P}$ is a probability measure on a set $\Omega$. Let $X\subseteq\Omega$ and $Y\subseteq\Omega$ be measurable. Also suppose that $Y=Y_{1}\cup\ldots\cup Y_{k}$, $Y_{i}\cap Y_{j}$ if $i\neq j$, and that each $Y_{i}$ is measurable. If $\alpha\in[0,1]$, $\varepsilon>0$, and $\mathbb{P}(X\ |\ Y_{i})\in[\alpha-\varepsilon,\alpha+\varepsilon]$ for all $i=1,\ldots,k$, then $\mathbb{P}(X\ |\ Y)\in[\alpha-\varepsilon,\alpha+\varepsilon]$.
3. Probability logic with aggregation functions
We consider a logic that we call probability logic with aggregation functions, or $PLA^{*}$, where $PLA^{*}$ is a more general version of $PLA$ and $PLA^{+}$ which where considered in [21] and [22], respectively, and of the probability logic considered by Jaeger in [14, 15]. $PLA^{*}$ is a logic with (truth) values in the unit interval $[0,1]$. With $PLA^{*}$ we can, for example, express the proportion of all elements in a domain that have some property, and the proportion need not be 0 or 1. As a nontrivial example of what $PLA^{*}$ can express, Example 3.12 shows that the PageRank can be expressed with $PLA^{*}$. On finite structures, all queries that can be expressed by first-order logic can also be expressed by $PLA^{*}$ (as stated by Lemma 3.13).
Recall that $[0,1]^{<\omega}$ denotes the set of all finite nonempty sequences of reals in the unit interval $[0,1]$.
Definition 3.1.
Let $k\in\mathbb{N}^{+}$.
(i) A function $\mathsf{C}:[0,1]^{k}\to[0,1]$ will also be called a ($k$-ary) connective.
(ii) A function $F:\big{(}[0,1]^{<\omega}\big{)}^{k}\to[0,1]$ which is symmetric in the following sense will be called a ($k$-ary) aggregation function: if $\bar{p}_{1},\ldots,\bar{p}_{k}\in[0,1]^{<\omega}$ and, for $i=1,\ldots,k$, $\bar{q}_{i}$ is a reordering of the entries in $\bar{p}_{i}$, then $F(\bar{q}_{1},\ldots,\bar{q}_{k})=F(\bar{p}_{1},\ldots,\bar{p}_{k})$.
In the definition below we extend the usual 0/1-valued connectives according to the semantics of Lukasiewicz logic (see for example [5, Section 11.2], or [24]) to get connectives which are continous as functions from $[0,1]$, or from $[0,1]\times[0,1]$, to $[0,1]$.
Definition 3.2.
Let
  1. (1)
    $\neg:[0,1]\to[0,1]$ be defined by $\neg(x)=1-x$,
  2. (2)
    $\wedge:[0,1]^{2}\to[0,1]$ be defined by $\wedge(x,y)=\min(x,y)$,
  3. (3)
    $\vee:[0,1]^{2}\to[0,1]$ be defined by $\vee(x,y)=\max(x,y)$, and
  4. (4)
[…]
The following aggregation functions are useful:
Definition 3.3.
For $\bar{p}=(p_{1},\ldots,p_{n})\in[0,1]^{<\omega}$, define
  1. (1)
    $\max(\bar{p})$ to be the maximum of $p_{1},\ldots,p_{n}$,
  2. (2)
    $\min(\bar{p})$ to be the minimum of $p_{1},\ldots,p_{n}$,
  3. (3)
    $\mathrm{am}(\bar{p})=(p_{1}+\ldots+p_{n})/n$, so ‘am’ is the arithmetic mean, or average,
  4. (4)
    $\mathrm{gm}(\bar{p})=\big{(}\prod_{i=1}^{n}p_{i}\big{)}^{(1/n)}$, so ‘gm’ is the geometric mean, and
  5. (5)
    for every $\beta\in(0,1]$, $\mathrm{length}^{-\beta}(\bar{p})=|\bar{p}|^{-\beta}$.
The aggregation functions above take only one sequence from $[0,1]^{<\omega}$ as input. But there are useful aggregation functions of higher arities, i.e. taking two or more sequences as input, as shown in Examples 5.5 – 5.7 in [21] and in Example 6.4 in [21].
For the rest of this section we fix a finite and relational signature $\sigma$.
Definition 3.4.
(Syntax of $PLA^{*}$) We define the formulas of $PLA^{*}(\sigma)$, as well as the set of free variables of a formula $\varphi$, denoted $Fv(\varphi)$, as follows:
  1. (1)
    For each $c\in[0,1]$, $c\in PLA^{*}(\sigma)$ (i.e. $c$ is a formula) and $Fv(c)=\emptyset$. We also let $\bot$ and $\top$ denote 00 and $1$, respectively.
  2. (2)
    For all variables $x$ and $y$, ‘$x=y$’ belongs to $PLA^{*}(\sigma)$ and $Fv(x=y)=\{x,y\}$.
  3. (3)
    For every $R\in\sigma$, say of arity $r$, and any choice of variables $x_{1},\ldots,x_{r}$, $R(x_{1},\ldots,x_{r})$ belongs to $PLA^{*}(\sigma)$ and $Fv(R(x_{1},\ldots,x_{r}))=\{x_{1},\ldots,x_{r}\}$.
  4. (4)
    If $k\in\mathbb{N}^{+}$, $\varphi_{1},\ldots,\varphi_{k}\in PLA^{*}(\sigma)$ and $\mathsf{C}:[0,1]^{k}\to[0,1]$ is a continuous connective, then $\mathsf{C}(\varphi_{1},\ldots,\varphi_{k})$ is a formula of $PLA^{*}(\sigma)$ and its set of free variables is $Fv(\varphi_{1})\cup\ldots\cup Fv(\varphi_{k})$.
  5. (5)
    Suppose that $\varphi_{1},\ldots,\varphi_{k}\in PLA^{*}(\sigma)$, $\chi_{1},\ldots,\chi_{k}\in PLA^{*}(\sigma)$, $\bar{y}$ is a sequence of distinct variables, and that $F:\big{(}[0,1]^{<\omega}\big{)}^{k}\to[0,1]$ is an aggregation function. Then
    $F(\varphi_{1},\ldots,\varphi_{k}:\bar{y}:\chi_{1},\ldots,\chi_{k})$
    is a formula of $PLA^{*}(\sigma)$ and its set of free variables is
    $\bigg{(}\bigcup_{i=1}^{k}Fv(\varphi_{i})\bigg{)}\setminus\mathrm{rng}(\bar{y}),$
    so this construction binds the variables in $\bar{y}$. The construction $F(\varphi_{1},\ldots,\varphi_{k}:\bar{y}:\chi_{1},\ldots,\chi_{k})$ will be called an aggregation (over $\bar{y}$) and the formulas $\chi_{1},\ldots,\chi_{k}$ are called the conditioning formulas of this aggregation.
Definition 3.5.
(i) A formula in $PLA^{*}(\sigma)$ without free variables is called a sentence.
(ii) In part (4) of Definition 3.4 the formulas $\varphi_{1},\ldots,\varphi_{k}$ are called subformulas of $\mathsf{C}(\varphi_{1},\ldots,\varphi_{k})$.
(iii) In part (5) of Definition 3.4 the formulas $\varphi_{1},\ldots,\varphi_{k}$ and $\chi_{1},\ldots,\chi_{k}$ are called subformulas of $F(\varphi_{1},\ldots,\varphi_{k}:\bar{y}:\chi_{1},\ldots,\chi_{k})$. We also call $\chi_{1},\ldots,\chi_{k}$ conditioning subformulas of the formula $F(\varphi_{1},\ldots,\varphi_{k}:\bar{y}:\chi_{1},\ldots,\chi_{k})$, and we say that (this instance of the aggregation) $F$ is conditioned on $\chi_{1},\ldots,\chi_{k}$.
(iv) We stipulate the following transitivity properties: If $\psi_{1}$ is a subformula of $\psi_{2}$ and $\psi_{2}$ is a subformula of $\psi_{3}$, then $\psi_{1}$ is a subformula of $\psi_{3}$. If $\psi_{1}$ is a conditioning subformula of $\psi_{2}$ and $\psi_{2}$ is a subformula of $\psi_{3}$, then $\psi_{1}$ is a conditioning subformula of $\psi_{3}$.
Notation 3.6.
When denoting a formula in $PLA^{*}(\sigma)$ by for example $\varphi(\bar{x})$ then we assume that $\bar{x}$ is a sequence of different variables and that every free variable in the formula denoted by $\varphi(\bar{x})$ belongs to $\mathrm{rng}(\bar{x})$ (but we do not require that every variable in $\mathrm{rng}(\bar{x})$ actually occurs as a free variable in the formula). If a formula is denoted by $\varphi(\bar{x},\bar{y})$ then we assume that $\mathrm{rng}(\bar{x})\cap\mathrm{rng}(\bar{y})=\emptyset$.
Definition 3.7.
The $PLA^{*}(\sigma)$-formulas described in parts (2) and (3) of Definition 3.4 are called first-order atomic $\sigma$-formulas. A $PLA^{*}(\sigma)$-formula is called a $\sigma$-literal if it has the form $\varphi(\bar{x})$ or $\neg\varphi(\bar{x})$, where $\varphi(\bar{x})$ is a first-order atomic formula and $\neg$ is like in Definition 3.2 (so it corresponds to negation when truth values are restricted to 0 and 1).
Definition 3.8.
(Semantics of $PLA^{*}$) For every $\varphi\in PLA^{*}(\sigma)$ and every sequence of distinct variables $\bar{x}$ such that $Fv(\varphi)\subseteq\mathrm{rng}(\bar{x})$ we associate a mapping from pairs $(\mathcal{A},\bar{a})$, where $\mathcal{A}$ is a finite $\sigma$-structure and $\bar{a}\in A^{|\bar{x}|}$, to $[0,1]$. The number in $[0,1]$ to which $(\mathcal{A},\bar{a})$ is mapped is denoted $\mathcal{A}(\varphi(\bar{a}))$ and is defined by induction on the complexity of formulas, as follows:
  1. (1)
    If $\varphi(\bar{x})$ is a constant $c$ from $[0,1]$, then $\mathcal{A}(\varphi(\bar{a}))=c$.
  2. (2)
    If $\varphi(\bar{x})$ has the form $x_{i}=x_{j}$, then $\mathcal{A}(\varphi(\bar{a}))=1$ if $a_{i}=a_{j}$, and otherwise $\mathcal{A}(\varphi(\bar{a}))=0$.
  3. (3)
    For every $R\in\sigma$, of arity $r$ say, if $\varphi(\bar{x})$ has the form $R(x_{i_{1}},\ldots,x_{i_{r}})$, then $\mathcal{A}(\varphi(\bar{a}))=1$ if $\mathcal{A}\models R(a_{i_{1}},\ldots,a_{i_{r}})$ (where $\models$’ has the usual meaning of first-order logic), and otherwise $\mathcal{A}(\varphi(\bar{a}))=0$.
  4. (4)
    If $\varphi(\bar{x})$ has the form $\mathsf{C}(\varphi_{1}(\bar{x}),\ldots,\varphi_{k}(\bar{x}))$, where $\mathsf{C}:[0,1]^{k}\to[0,1]$ is a continuous connective, then
    $\mathcal{A}\big{(}\varphi(\bar{a})\big{)}\ =\ \mathsf{C}\big{(}\mathcal{A}(\varphi_{1}(\bar{a})),\ldots,\mathcal{A}(\varphi_{k}(\bar{a}))\big{)}.$
  5. (5)
    Suppose that $\varphi(\bar{x})$ has the form
    $F(\varphi_{1}(\bar{x},\bar{y}),\ldots,\varphi_{k}(\bar{x},\bar{y}):\bar{y}:\chi_{1}(\bar{x},\bar{y}),\ldots,\chi_{k}(\bar{x},\bar{y}))$
    where $\bar{x}$ and $\bar{y}$ are sequences of distinct variables, $\mathrm{rng}(\bar{x})\cap\mathrm{rng}(\bar{y})=\emptyset$, and $F:\big{(}[0,1]^{<\omega}\big{)}^{k}\to[0,1]$ is an aggregation function. If, for every $i=1,\ldots,k$, the set $\{\bar{b}\in A^{|\bar{y}|}:\mathcal{A}(\chi_{i}(\bar{a},\bar{b}))=1\}$ is nonempty, then let
    $\bar{p}_{i}=\big{(}\mathcal{A}\big{(}\varphi_{i}(\bar{a},\bar{b})\big{)}:\bar{b}\in A^{|\bar{y}|}\text{ and }\mathcal{A}\big{(}\chi_{i}(\bar{a},\bar{b})\big{)}=1\big{)}$
    and
    $\mathcal{A}\big{(}\varphi(\bar{a})\big{)}=F(\bar{p}_{1},\ldots,\bar{p}_{k}).$
    Otherwise let $\mathcal{A}\big{(}\varphi(\bar{a})\big{)}=0$.
Definition 3.9.
Let $\varphi(\bar{x}),\psi(\bar{x})\in PLA^{*}(\sigma)$. We say that $\varphi$ and $\psi$ are equivalent if for every finite $\sigma$-structure $\mathcal{A}$ and every $\bar{a}\in A^{|\bar{x}|}$, $\mathcal{A}(\varphi(\bar{a}))=\mathcal{A}(\psi(\bar{a}))$.
Notation 3.10.
(i) For any formula $\varphi(\bar{x},\bar{y})\in PLA^{*}(\sigma)$, finite $\sigma$-structure $\mathcal{A}$ and $\bar{a}\in A^{|\bar{x}|}$, let
$\varphi(\bar{a},\mathcal{A})=\{\bar{b}\in A^{|\bar{y}|}:\mathcal{A}(\varphi(\bar{a},\bar{b}))=1\}.$
(ii) Let $\varphi(\bar{x})\in PLA^{*}(\sigma)$, let $\mathcal{A}$ be a finite $\sigma$-structure, and let $\bar{a}\in A^{|\bar{x}|}$. Then ‘$\mathcal{A}\models\varphi(\bar{a})$’ means the same as ‘$\mathcal{A}(\varphi(\bar{a}))=1$’.
(iii) Let $\varphi(\bar{x}),\psi(\bar{x})\in PLA^{*}(\sigma)$. When writing ‘$\varphi(\bar{x})\models\psi(\bar{x})$’ we mean that for every finite $\sigma$-structure $\mathcal{A}$ and all $\bar{a}\in A^{|\bar{x}|}$, if $\mathcal{A}\models\varphi(\bar{a})$ then $\mathcal{A}\models\psi(\bar{a})$.
Definition 3.11.
(i) A formula in $PLA^{*}(\sigma)$ such that no aggregation function occurs in it is called aggregation-free.
(ii) If $\varphi(\bar{x})\in PLA^{*}(\sigma)$ and for every finite $\sigma$-structure $\mathcal{A}$, and every $\bar{a}\in A^{|\bar{x}|}$, $\mathcal{A}(\varphi(\bar{a}))$ is 0 or 1, then we call $\varphi(\bar{x})$ $0/1$-valued.
(iii) If $L\subseteq PLA^{*}(\sigma)$ and every formula in $L$ is 0/1-valued, then we say that $L$ is $0/1$-valued.
(iv) Let $L\subseteq PLA^{*}(\sigma)$. A formula of $PLA^{*}(\sigma)$ is called an $L$-basic formula if it has the form $\bigwedge_{i=1}^{k}\big{(}\varphi_{i}(\bar{x})\to c_{i}\big{)}$ where $\varphi_{i}(\bar{x})\in L$ and $c_{i}\in[0,1]$ for all $i=1,\ldots,k$. (We will only use this notion when $L$ is $0/1$-valued.)
Example 3.12.
We exemplify what can be expressed with $PLA^{*}(\sigma)$, provided that it contains a binary relation symbol, with the notion of PageRank [6]. The PageRank of an internet site can be approximated in “stages” as follows (if we supress the “damping factor” for simplicity), where $IN_{x}$ is the set of sites that link to $x$, and $OUT_{y}$ is the set of sites that $y$ link to:
$PR_{0}(x)=1/N\text{ where $N$ is the number of sites,}$,
$PR_{k+1}(x)=\sum_{y\in IN_{x}}\frac{PR_{k}(y)}{|OUT_{y}|}.$
It is not difficult to prove, by induction on $k$, that for every $k$ the sum of all $PR_{k}(x)$ as $x$ ranges over all sites is 1. Hence the sum in the definition of $PR_{k+1}$ is less or equal to 1 (and this will matter later). Let $E\in\sigma$ be a binary relation symbol representing a link. Then $PR_{0}(x)$ is expressed by the $PLA^{*}(\sigma)$-formula $\mathrm{length}^{-1}(x=x:y:\top)$.
Suppose that $PR_{k}(x)$ is expressed by $\varphi_{k}(x)\in PLA^{*}(\sigma)$. Note that multiplication is a continuous connective from $[0,1]^{2}$ to $[0,1]$ so it can be used in $PLA^{*}(\sigma)$-formulas. Then observe that the quantity $|OUT_{y}|^{-1}$ is expressed by the $PLA^{*}(\sigma)$-formula
$\mathrm{length}^{-1}\big{(}y=y:z:E(y,z)\big{)}$
which we denote by $\psi(y)$. Let $\mathrm{tsum}:[0,1]^{<\omega}\to[0,1]$ be the “truncated sum” defined by letting $\mathrm{tsum}(\bar{p})$ be the sum of all entries in $\bar{p}$ if the sum is at most 1, and otherwise $\mathrm{tsum}(\bar{p})=1$. Then $PR_{k+1}(x)$ is expressed by the $PLA^{*}(\sigma)$-formula
$\mathrm{tsum}\big{(}x=x\wedge(\varphi_{k}(y)\cdot\psi(y)):y:E(y,x)\big{)}.$
With $PLA^{*}(\sigma)$ we can also define all stages of the SimRank [16] in a simpler way than done in [21] with the sublogic $PLA(\sigma)\subseteq PLA^{*}(\sigma)$.
Lemma 3.13.
Suppose that $\varphi(\bar{x})$ is a first-order formula over $\sigma$. Then there is a 0/1-valued $\psi(\bar{x})\in PLA^{*}(\sigma)$ such that for every finite $\sigma$-structure $\mathcal{A}$ and every $\bar{a}\in A^{|\bar{x}|}$, $\mathcal{A}\models\varphi(\bar{a})$ if and only if $\mathcal{A}(\psi(\bar{a}))=1$.
Proof. We argue by induction on the complexity of $\varphi$. If it is atomic then the conclusion follows from parts (2) and (3) of the syntax and semantics of $PLA^{*}$. The inductive step for the connectives of first-order logic follows since in $PLA^{*}$ we can use their extensions to the unit interval as defined in Definition 3.2.
Now suppose that $\varphi(\bar{x})$ has the form $\exists y\varphi(\bar{x},y)$. By the induction hypothesis, there is a 0/1-valued $\psi(\bar{x},y)\in PLA^{*}(\sigma)$ such that for all finite $\sigma$-structures $\mathcal{A}$, all $\bar{a}\in A^{|\bar{x}|}$, and $b\in A$, $\mathcal{A}\models\varphi(\bar{a},b)$ if and only if $\mathcal{A}(\psi(\bar{a},b))=1$. Let $\mathcal{A}$ be a finite $\sigma$-structure and $\bar{a}\in A^{|\bar{x}|}$. Then
$\mathcal{A}\models\exists y\varphi(\bar{a},y)\text{ if and only if }\mathcal{A}\big{(}\max(\psi(\bar{a},y):y:\top)\big{)}=1$
where $\max(\psi(\bar{a},y):y:\top)$ is 0/1-valued. In a similar way, the aggregation function min can play the role of $\forall$. $\square$
Notation 3.14.
(Using $\exists$ and $\forall$ as abbreviations) Due to Lemma 3.13, if $\varphi(\bar{x},\bar{y})\in PLA^{*}(\sigma)$ is a 0/1-valued formula then we will often write ‘$\exists\bar{y}\varphi(\bar{x},\bar{y})$’ to mean the same as ‘$\max(\varphi(\bar{x},\bar{y}):\bar{y}:\top)$’, and ‘$\forall\bar{y}\varphi(\bar{x},\bar{y})$’ to mean the same as ‘$\min(\varphi(\bar{x},\bar{y}):\bar{y}:\top)$’.
The next basic lemma has analogue in first-order logic and is proved straightforwardly by induction on the complexity of $PLA^{*}$-formulas.
Lemma 3.15.
Suppose that $\sigma^{\prime}\subseteq\sigma$, $\varphi(\bar{x})\in PLA^{*}(\sigma^{\prime})$, $\mathcal{A}$ is a finite $\sigma$-structure, $\mathcal{A}^{\prime}=\mathcal{A}\negthickspace\upharpoonright\negthickspace\sigma^{\prime}$, and $\bar{a}\in A^{|\bar{x}|}$. Then $\mathcal{A}(\varphi(\bar{a}))=\mathcal{A}^{\prime}(\varphi(\bar{a}))$.
4. A general method for asymptotic elimination of aggregation functions
Let $\sigma$ be a finite and relational signature. In all of this section let $D_{n}$, $n\in\mathbb{N}^{+}$, be finite sets such that $\lim_{n\to\infty}|D_{n}|=\infty$, and let $\mathbf{W}_{n}$ be a set of $\sigma$-structures with domain $D_{n}$. We will describe a method for “asymptotically eliminating aggregation functions” if
  • the aggregation functions in question are “sufficiently” continuous, and
  • if there are 0/1-valued formulas with certain properties, including that the proportions of some definable sets “stabilize” as the domain size tends to infinity.
This method is studied in more detail by Koponen and Weitkämper in [23] and the definitions below and Theorem 4.8 come from [23]. Theorem 4.8 will be used later (in Section 7.5) to prove our main results.
Definition 4.1.
By a sequence of probability distributions (for $(\mathbf{W}_{n}:n\in\mathbb{N}^{+})$) we mean a sequence $\mathbb{P}=(\mathbb{P}_{n}:n\in\mathbb{N}^{+})$ such that for every $n$, $\mathbb{P}_{n}$ is a probability distribution on $\mathbf{W}_{n}$.
Definition 4.2.
Let $\varphi(\bar{x}),\psi(\bar{x})\in PLA^{*}(\sigma)$ and let $\mathbb{P}=(\mathbb{P}_{n}:n\in\mathbb{N}^{+})$ be a sequence of probability distributions. We say that $\varphi(\bar{x})$ and $\psi(\bar{x})$ are asymptotically equivalent (with respect to $\mathbb{P}$) if for all $\varepsilon>0$
$\lim_{n\to\infty}\mathbb{P}_{n}\Big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\text{ for all $\bar{a}\in(B_{n})^{|\bar{x}|}$, $|\mathcal{A}(\varphi(\bar{a}))-\mathcal{A}(\psi(\bar{a}))|\leq\varepsilon$}\big{\}}\Big{)}\ =\ 1.$
For the rest of this section we fix a sequence $\mathbb{P}=(\mathbb{P}_{n}:n\in\mathbb{N}^{+})$ of probability distributions.
To define the notions of continuity that we will use we need the notion of convergence testing sequence. It generalizes a similar notion used by Jaeger in [14]. The intuition is that a sequence $\bar{p}_{n}\in[0,1]^{<\omega}$, $n\in\mathbb{N}^{+}$, is convergence testing for parameters $c_{1},\ldots,c_{k}\in[0,1]$ and $\alpha_{1},\ldots\alpha_{k}\in[0,1]$ if the length of $\bar{p}_{n}$ tends to infinity as $n\to\infty$ and, as $n\to\infty$, all entries in $\bar{p}_{n}$ congregate ever closer to the “convergence points” in $\{c_{1},\ldots,c_{k}\}$, and the proportion of entries in $\bar{p}$ which congregate to $c_{i}$ converges to $\alpha_{i}$.
Definition 4.3.
(i) A sequence $\bar{p}_{n}\in[0,1]^{<\omega}$, $n\in\mathbb{N}$, is called convergence testing for parameters $c_{1},\ldots,c_{k}\in[0,1]$ and $\alpha_{1},\ldots\alpha_{k}\in[0,1]$ if the following hold, where $p_{n,i}$ denotes the $i$th entry of $\bar{p}_{n}$:
  1. (1)
    $|\bar{p}_{n}|<|\bar{p}_{n+1}|$ for all $n\in\mathbb{N}$.
  2. (2)
    For every disjoint family of open (with respect to the induced topology on $[0,1]$) intervals $I_{1},\ldots I_{k}\subseteq[0,1]$ such that $c_{i}\in I_{i}$ for each $i$, there is an $N\in\mathbb{N}$ such that $\mathrm{rng}(\bar{p}_{n})\subseteq\bigcup\limits_{j=1}^{k}I_{j}$ for all $n\geq N$, and for every $j\in\{1,\ldots,k\}$,
    $\lim\limits_{n\rightarrow\infty}\frac{\left|\{i:p_{n,i}\in I_{j}\}\right|}{|\bar{p}_{n}|}=\alpha_{j}.$
(ii) More generally, a sequence of $m$-tuples $(\bar{p}_{1,n},\ldots,\bar{p}_{m,n})\in\big{(}[0,1]^{<\omega}\big{)}^{m}$, $n\in\mathbb{N}$, is called convergence testing for parameters $c_{i,j}\in[0,1]$ and $\alpha_{i,j}\in[0,1]$, where $i\in\{1,\ldots,m\}$, $j\in\{1,\ldots,k_{i}\}$ and $k_{1},\ldots k_{m}\in\mathbb{N}^{+}$, if for every fixed $i\in\{1,\ldots,m\}$ the sequence $\bar{p}_{i,n}$, $n\in\mathbb{N}$, is convergence testing for $c_{i,1},\ldots,c_{i,k_{i}}$, and $\alpha_{i,1},\ldots,\alpha_{i,k_{i}}$.
Definition 4.4.
An aggregation function
$F:\big{(}[0,1]^{<\omega}\big{)}^{m}\to[0,1]$ is called ct-continuous (convergence test continuous) with respect to the sequence of parameters $c_{i,j},\alpha_{i,j}\in[0,1]$, $i=1,\ldots,m$, $j=1,\ldots,k_{i}$, if the following condition holds:
  1. For all convergence testing sequences of $m$-tuples $(\bar{p}_{1,n},\ldots,\bar{p}_{m,n})\in\big{(}[0,1]^{<\omega}\big{)}^{m}$, $n\in\mathbb{N}$, and $(\bar{q}_{1,n},\ldots,\bar{q}_{m,n})\in\big{(}[0,1]^{<\omega}\big{)}^{m}$, $n\in\mathbb{N}$, with the same parameters $c_{i,j},\alpha_{i,j}\in[0,1]$, $\underset{n\rightarrow\infty}{\lim}|F(\bar{p}_{1,n},\ldots,\bar{p}_{m,n})-F(\bar{q}_{1,n},\ldots,\bar{q}_{m,n})|=0$.
Definition 4.5.
Let $F:\big{(}[0,1]^{<\omega}\big{)}^{m}\to[0,1]$.
(i) We call $F$ continuous (or strongly admissible) if $F$ is ct-continuous with respect to every choice of parameters $c_{i,j},\alpha_{i,j}\in[0,1]$, $i=1,\ldots,m$ and $j=1,\ldots,k_{i}$ (for arbitrary $m$ and $k_{i}$).
(ii) We call $F$ admissible if $F$ is ct-continuous with respect to every choice of parameters $c_{i,j},\alpha_{i,j}\in[0,1]$, $i=1,\ldots,m$ and $j=1,\ldots,k_{i}$ (for arbitrary $m$ and $k_{i}$) such that $\alpha_{i,j}>0$ for all $i$ and $j$.
Example 4.6.
[…]
The method that we consider for asymptotically eliminating continuous or admissible aggregation functions can be applied if one can find sets $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ of 0/1-valued formulas that satisfy the conditions of the following assumption:
Assumption 4.7.
Suppose that $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ are $0/1$-valued and that the following conditions hold:
  1. (1)
    For every aggregation-free $\varphi(\bar{x})\in PLA^{*}(\sigma)$ there is an $L_{0}$-basic formula $\varphi^{\prime}(\bar{x})$ which is asymptotically equivalent to $\varphi(\bar{x})$ with respect to $\mathbb{P}$.
  2. (2)
    For every $m\in\mathbb{N}^{+}$ and all $\varphi_{1}(\bar{x},\bar{y}),\ldots,\varphi_{m}(\bar{x},\bar{y})\in L_{0}$, there are $L_{\varphi_{1}(\bar{x},\bar{y})},\ldots,L_{\varphi_{m}(\bar{x},\bar{y})}\subseteq L_{1}$ such that if $\chi_{j}(\bar{x},\bar{y})\in L_{\varphi_{j}(\bar{x},\bar{y})}$ for $j=1,\ldots,m$, then there are $s,t\in\mathbb{N}^{+}$, $\varphi^{\prime}_{i}(\bar{x})\in L_{0}$, $\alpha_{i,j}\in[0,1]$, for $i=1,\ldots,s$, $j=1,\ldots,m$, and $\chi^{\prime}_{i}(\bar{x})\in L_{0}$, for $i=1,\ldots,t$, such that for every $\varepsilon>0$ and $n$ there is $\mathbf{Y}^{\varepsilon}_{n}\subseteq\mathbf{W}_{n}$ such that $\lim_{n\to\infty}\mathbb{P}_{n}(\mathbf{Y}^{\varepsilon}_{n})=1$ and for every $\mathcal{A}\in\mathbf{Y}^{\varepsilon}_{n}$ the following hold:
    $(a)\ \mathcal{A}\models\forall\bar{x}\bigvee_{i=1}^{s}\varphi^{\prime}_{i}(\bar{x}),$
    $(b)\ \text{if $i\neq j$ then }\mathcal{A}\models\forall\bar{x}\neg(\varphi^{\prime}_{i}(\bar{x})\wedge\varphi^{\prime}_{j}(\bar{x})),$
    $(c)\ \mathcal{A}\models\forall\bar{x}\Big{(}\Big{(}\bigvee_{i=1}^{m}\neg\exists\bar{y}\chi_{i}(\bar{x},\bar{y})\Big{)}\leftrightarrow\Big{(}\bigvee_{i=1}^{t}\chi^{\prime}_{i}(\bar{x})\Big{)}\Big{)},\text{ and}$
    $(d)\ \text{for all $i=1,\ldots,s$ and $j=1,\ldots,m$, if $\bar{a}\in(D_{n})^{|\bar{x}|}$, and $\mathcal{A}\models\varphi^{\prime}_{i}(\bar{a})$,}$
[…]
In Section 7.5 we will use the following result, which is a less detailed version of Theorem 5.9 in [23].
Theorem 4.8.
[23] Suppose that $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ are 0/1-valued and that Assumption 4.7 holds. Let $F:\big{(}[0,1]^{<\omega}\big{)}^{m}\to[0,1]$, let $\psi_{i}(\bar{x},\bar{y})\in PLA^{*}(\sigma)$, for $i=1,\ldots,m$, and suppose that each $\psi_{i}(\bar{x},\bar{y})$ is asymptotically equivalent to an $L_{0}$-basic formula
$\bigwedge_{k=1}^{s_{i}}(\psi_{i,k}(\bar{x},\bar{y})\to c_{i,k})\quad\text{(so $\psi_{i,k}\in L_{0}$ for all $i$ and $k$).}$.
Suppose that for $i=1,\ldots,m$, $\chi_{i}(\bar{x},\bar{y})\in\bigcap_{k=1}^{s_{i}}L_{\psi_{i,k}(\bar{x},\bar{y})}$. Let $\varphi(\bar{x})$ denote the $PLA^{*}(\sigma)$-formula
$F\big{(}\psi_{1}(\bar{x},\bar{y}),\ldots,\psi_{m}(\bar{x},\bar{y}):\bar{y}:\chi_{1}(\bar{x},\bar{y}),\ldots,\chi_{m}(\bar{x},\bar{y})\big{)}.$
(i) If $F$ is continuous then $\varphi(\bar{x})$ is asymptotically equivalent to an $L_{0}$-basic formula with respect to $\mathbb{P}$.
(ii) Suppose, in addition, that if $\varphi_{j}(\bar{x},\bar{y})$, $\chi_{j}(\bar{x},\bar{y})$, $\varphi^{\prime}_{i}(\bar{x})$, $\mathbf{Y}^{\varepsilon}_{n}$ and $\alpha_{i,j}$ are like in part 2 of Asumption 4.7 and $\alpha_{i,j}=0$, then, for all sufficiently large $n$, all $\bar{a}\in(D_{n})^{|\bar{x}|}$ and all $\mathcal{A}\in\mathbf{Y}^{\varepsilon}_{n}$, if $\mathcal{A}\models\varphi^{\prime}_{i}(\bar{a})$ then $\varphi_{j}(\bar{a},\mathcal{A})\cap\chi_{j}(\bar{a},\mathcal{A})=\emptyset$. If $F$ is admissible then $\varphi(\bar{x})$ is asymptotically equivalent to an $L_{0}$-basic formula with respect to $\mathbb{P}$.
5. Directed acyclic graphs, trees and closure types
For the rest of this article we let $\tau=\{E\}$ where $E$ is a binary relation symbol. Moreover, $\sigma$ will always denote a finite relational signature such that $\tau\subseteq\sigma$, although this later assumption will be repeated.
Definition 5.1.
(i) By a directed acyclic graph (DAG) we mean a finite $\tau$-structure $\mathcal{G}$ such that
  1. (1)
    $\mathcal{G}\models\forall x,y\big{(}E(x,y)\rightarrow(x\neq y\wedge\neg E(y,x))\big{)}$, and
  2. (2)
    there do not exist $k\in\mathbb{N}^{+}$ and distinct $a_{0},\ldots,a_{k}\in G$ such that $\mathcal{G}\models E(a_{i},a_{i+1})$ for all $i=0,\ldots,k-1$ and $\mathcal{G}\models E(a_{k},a_{0})$.
Note that we allow the domain of a DAG $\mathcal{G}$ to be empty.
Definition 5.2.
Suppose that $\mathcal{G}$ is a DAG.
(i) If $a,b\in G$ and $\mathcal{G}\models E(b,a)$ then we call $b$ a parent of $a$ and $a$ a child of $b$, and the set of parents of $a$ is denoted $\mathrm{par}(a)$.
(ii) If $\mathcal{G}$ is a DAG and $a\in G$ then $a$ is called a root if $\mathcal{G}\models\neg\exists xE(x,a)$.
(iii) Let $a,b\in G$. A directed path (of length $l$) from $a$ to $b$ is a sequence $c_{0},c_{1},\ldots,c_{l}\in T$ such that $l\geq 1$, $\mathcal{T}\models E(c_{i},c_{i+1})$ for all $i=0,\ldots,l-1$, $a=c_{0}$ and $b=c_{l}$. If there is a directed path from $a$ to $b$ then $a$ is called an ancestor of $b$ and $b$ is called a successor of $a$.
(iv) The level 00 of $\mathcal{G}$ is the set of all roots of $\mathcal{G}$. An element $a\in G$ belongs to level $l+1$ (of $\mathcal{G}$) if $a$ has a parent in level $l$.
(v) If the domain of $\mathcal{G}$ is nonempty, then the height of $\mathcal{G}$ is the largest $l$ such that level $l$ of $\mathcal{G}$ is nonempty. We also adopt the convention that if the domain of $\mathcal{G}$ is empty then its height is $-1$.
Definition 5.3.
By a tree we mean a finite $\tau$-structure $\mathcal{T}$ such that
  1. (1)
    $\mathcal{T}\models\forall x,y\big{(}E(x,y)\rightarrow(x\neq y\wedge\neg E(y,x))\big{)}$,
  2. (2)
    there is a unique element $a\in T$, called the root, such that $\mathcal{T}\models\neg\exists xE(x,a)$, and
  3. (3)
    for all $a\in T$, if $a$ is not the root then there is a unique $b\in T$, called the parent of $a$, such that $\mathcal{T}\models E(b,a)$, and in this case $a$ is called a child of $b$.
Note that every tree is a DAG.
Definition 5.4.
Let $\mathcal{T}$ be a tree.
(i) A tree $\mathcal{T}^{\prime}$ is a subtree of $\mathcal{T}$ if it is a substructure of $\mathcal{T}$ in the model theoretic sense.
(ii) A subtree $\mathcal{T}^{\prime}$ of $\mathcal{T}$ is rooted in $a\in T$ if $a$ is the root of $\mathcal{T}^{\prime}$.
(iii) Let $\mathcal{T}^{\prime}$ be a tree and $a\in T$. Then $\mathsf{N}_{\mathcal{T}}(a,\mathcal{T}^{\prime})$ denotes the number of subtrees of $\mathcal{T}$ that are rooted in $a$ and isomorphic to $\mathcal{T}^{\prime}$. If $\mathcal{T}$ is clear from the context we may just write $\mathsf{N}(a,\mathcal{T}^{\prime})$.
Definition 5.5.
Let $\mathcal{T}$ be a tree.
(i) Two elements of the tree with a common parent are called siblings.
(ii) If $B\subseteq T$ then the closure of $B$ (in $\mathcal{T}$), denoted $\mathrm{cl}_{\mathcal{T}}(B)$ or just $\mathrm{cl}(B)$, is defined by
$\mathrm{cl}_{\mathcal{T}}(B)=\big{\{}a\in T:\text{ $a\in B$, or $a$ is the root, or $a$ is an ancestor of some element in $B$}\big{\}}.$
If $\bar{a}$ is a sequence of elements from $T$ then we define $\mathrm{cl}_{\mathcal{T}}(\bar{a})=\mathrm{cl}_{\mathcal{T}}(\mathrm{rng}(\bar{a}))$.
(iii) A set $B\subseteq T$ (sequence $\bar{b}$) is closed (in $\mathcal{T}$) if $\mathrm{cl}_{\mathcal{T}}(B)=B$ ($\mathrm{cl}_{\mathcal{T}}(\bar{b})=\mathrm{rng}(\bar{b})$).
(iv) Let $\mathcal{A}$ be a $\sigma$-structure such that $\mathcal{T}=\mathcal{A}\negthickspace\upharpoonright\negthickspace\tau$ is a tree. If $B\subseteq A$ then the closure of $B$ (in $\mathcal{A}$), denoted $\mathrm{cl}_{\mathcal{A}}(B)$ or just $\mathrm{cl}(B)$, is defined by $\mathrm{cl}_{\mathcal{A}}(B)=cl_{\mathcal{T}}(B)$, and similarly for sequences of elements from $A$. We say that $B\subseteq A$ is closed in $\mathcal{A}$ if it is closed in $\mathcal{T}$.
(v) If $\mathcal{A}$ is as in the previous part then we may use notions such as root, child, level, etcetera, with reference to the underlying tree $\mathcal{A}\negthickspace\upharpoonright\negthickspace\tau$.
Remark 5.6.
Observe that for every tree $\mathcal{T}$ and every $B\subseteq T$, $\mathrm{cl}_{\mathcal{T}}(\mathrm{cl}_{\mathcal{T}}(B))=\mathrm{cl}_{\mathcal{T}}(B)$. Note also that the statement “$\{x_{1},\ldots,x_{k}\}$ is closed” is expressed, in every tree, by the formula
$\forall z\bigg{(}\Big{(}\bigvee_{i=1}^{k}E(z,x_{i})\Big{)}\rightarrow\Big{(}\bigvee_{i=1}^{k}z=x_{i}\Big{)}\bigg{)}.$
Suppose that $\mathcal{T}$ is a tree, $a\in T$, $\{b_{1},\ldots,b_{k}\}\subseteq T$ and $a\in\mathrm{cl}_{\mathcal{T}}(b_{1},\ldots,b_{k})$. Then either $a=b_{i}$ for some $i$, or for some $l\in\mathbb{N}^{+}$ and $i$ there is a directed path from $a$ to some $b_{i}$ of length $l$. The statement “there is a directed path from $x$ to $y$ of length $l$” can be expressed by a first-order formula, say $\xi_{l}(x,y)$. We will use the expression ‘$x\in\mathrm{cl}(y)$’ to denote a formula $\xi_{l}(x,y)$ for some unspecified $l$, and more generally, the expression ‘$x\in\mathrm{cl}(y_{1},\ldots,y_{k})$’ means $\bigvee_{i=1}^{k}x\in\mathrm{cl}(y_{i})$.
Finally, observe that if $\mathcal{T}$ is a tree and $B\subseteq T$ is closed, then the substructure $\mathcal{T}\negthickspace\upharpoonright\negthickspace B$ is also a tree.
Definition 5.7.
Let $\tau\subseteq\sigma$ and let $\mathcal{T}$ be a tree. A formula $\varphi(x_{1},\ldots,x_{k})$ is called an atomic type over $\sigma$ with respect to $\mathcal{T}$ if
  1. (1)
    there is a $\sigma$-structure $\mathcal{A}$ that expands $\mathcal{T}$ and $\mathcal{A}\models\exists x_{1},\ldots,x_{k}\varphi(x_{1},\ldots,x_{k})$,
  2. (2)
    $\varphi(x_{1},\ldots,x_{k})$ is a conjunction of $\sigma$-literals,
  3. (3)
    for all $i,j\in\{1,\ldots,k\}$, either $E(x_{i},x_{j})$ or $\neg E(x_{i},x_{j})$ is a conjunct of $\varphi(x_{1},\ldots,x_{k})$, and
  4. (4)
    for all different $i$ and $j$ in $\{1,\ldots,k\}$, $x_{i}\neq x_{j}$ is a conjunct of $\varphi(x_{1},\ldots,x_{k})$.
If, moreover, for every $R\in\sigma$, or arity $r$ say, and all $i_{1},\ldots,i_{r}\in\{1,\ldots,k\}$, either $R(x_{i_{1}},\ldots,x_{i_{r}})$ or $\neg R(x_{i_{1}},\ldots,x_{i_{r}})$ is a conjunct of $\varphi(x_{1},\ldots,x_{k})$, then we call $\varphi(x_{1},\ldots,x_{k})$ a complete atomic type over $\sigma$ with respect to $\mathcal{A}$. If the particular tree $\mathcal{T}$ is not important in the context we may omit the phrase ‘with respect to $\mathcal{T}$’.
Note that every atomic type over $\tau$ (with respect to some tree $\mathcal{T}$) is a complete atomic type over $\tau$, but this implication does not (in general) hold if $\tau$ is replaced by a proper expansion $\sigma\supset\tau$.
Definition 5.8.
Let $\tau\subseteq\sigma$ and let $\mathcal{T}$ be a tree. A formula $\psi(x_{1},\ldots,x_{k})$ is called a closure type over $\sigma$ with respect to $\mathcal{T}$ if it is equivalent to a formula of the form
$\exists y_{1},\ldots,y_{m}\Big{(}\varphi(x_{1},\ldots,x_{k},y_{1},\ldots,y_{m})\ \wedge\ \text{ `$\{x_{1},\ldots,x_{k},y_{1},\ldots,y_{m}\}$ isclosed'}\Big{)}.$’ ) .
where $\varphi(x_{1},\ldots,x_{k},y_{1},\ldots,y_{m})$ is an atomic type over $\sigma$ with respect to $\mathcal{T}$ and
$\varphi(x_{1},\ldots,x_{k},y_{1},\ldots,y_{m})\models\bigwedge_{i=1}^{m}y_{i}\in\mathrm{cl}(x_{1},\ldots,x_{k}).$
If, in addition, the formula $\varphi$ is a complete atomic type over $\sigma$, then we call $\psi$ a complete closure type over $\sigma$ with respect to $\mathcal{T}$. If the particular tree $\mathcal{T}$ is not important in the context we may omit the phrase ‘with respect to $\mathcal{T}$’.
We allow the (important special) case when the quantifier prefix ‘$\exists y_{1},\ldots,y_{m}$’ is empty and the variables ‘$y_{1},\ldots,y_{m}$’ do not occur. We also allow the case when the sequence $x_{1},\ldots,x_{k}$ is empty, and in this case $\psi$ is a sentence, $m=1$, $y_{1}$ denotes the root, and $\psi$ expresses which relations the root satisfies.
Observe that every closure type over $\tau$ (with respect to some tree $\mathcal{T}$) is a complete closure type over $\tau$, but this implication does not (in general) hold if $\tau$ is replaced by a proper expansion $\sigma\supset\tau$. It is easy to see that if $\varphi(\bar{x})$ is a closure-type over $\sigma\supseteq\tau$ with respect to a tree $\mathcal{T}$, then there are infinitely many trees $\mathcal{T}^{\prime}$ such that $\mathcal{T}^{\prime}$ has the same height as $\mathcal{T}$, $\mathcal{T}$ is a subtree (that is, substructure) of $\mathcal{T}^{\prime}$, and $\varphi(\bar{x})$ is a closure-type over $\sigma$ with respect to $\mathcal{T}^{\prime}$.
Definition 5.9.
Let $\tau\subseteq\sigma^{\prime}\subseteq\sigma$ and let $p(\bar{x})$ be a (complete) closure type over $\sigma$ with respect to a tree $\mathcal{T}$. Also let $\bar{y}$ be a subsequence of $\bar{x}$.
(i) The restriction of $p(\bar{x})$ to $\sigma^{\prime}$, denoted $p\negthickspace\upharpoonright\negthickspace\sigma^{\prime}$, is a closure-type $p^{\prime}(\bar{x})$ over $\sigma^{\prime}$ (with respect to $\mathcal{T}$) such that $p(\bar{x})\models p^{\prime}(\bar{x})$, and for every closure-type $p^{*}(\bar{x})$ over $\sigma^{\prime}$, if $p(\bar{x})\models p^{*}(\bar{x})$ then $p^{\prime}(\bar{x})\models p^{*}(\bar{x})$.
(ii) The restriction of $p(\bar{x})$ to $\bar{y}$, denoted $p\negthickspace\upharpoonright\negthickspace\bar{y}$, is a closure type over $\sigma$ (with respect to $\mathcal{T}$) $p^{\prime}(\bar{y})$ such that $p(\bar{x})\models p^{\prime}(\bar{y})$, and for every closure-type $p^{*}(\bar{y})$ over $\sigma$, if $p(\bar{x})\models p^{*}(\bar{y})$ then $p^{\prime}(\bar{y})\models p^{*}(\bar{y})$.
It is not hard to see that, under the same assumptions as in the above definition, if $p(\bar{x})$ is a closure type over $\sigma$ with respect to the tree $\mathcal{T}$, then $p\negthickspace\upharpoonright\negthickspace\sigma^{\prime}$ is a closure-type over $\sigma^{\prime}$ with respect to $\mathcal{T}$, and $p\negthickspace\upharpoonright\negthickspace\bar{y}$ are closure type over $\sigma$ with respect to $\mathcal{T}$. It will be convenient to use ‘$\models_{tree}$’ to denote consequence restricted to expansions of trees, as defined below.
Definition 5.10.
Suppose that $\tau\subseteq\sigma$ and let $\varphi(\bar{x}),\psi(\bar{x})\in PLA^{*}(\sigma)$ are $0/1$-valued formulas. By the expression $\varphi(\bar{x})\models_{tree}\psi(\bar{x})$ we mean that if $\mathcal{A}$ is a $\sigma$-structure that expands a tree, $\bar{a}\in(A)^{|\bar{x}|}$, and $\mathcal{A}\models\varphi(\bar{a})$, then $\mathcal{A}\models\psi(\bar{a})$. The expression $\models_{tree}\psi(\bar{x})$ means the same as $\top\models_{tree}\psi(\bar{x})$, or equivalently, that $\forall\bar{x}\psi(\bar{x})$ is true in all $\sigma$-structures that expand a tree.
Definition 5.11.
Let $\tau\subseteq\sigma$ and let $p(\bar{x})$ be a closure type over $\sigma$ (with respect to some tree $\mathcal{T}$) where $\bar{x}=(x_{1},\ldots,x_{k})$. We call $p(\bar{x})$ self-contained if $p(\bar{x})$ implies that $\bar{x}$ is closed, or more formally, if
$p(\bar{x})\ \models\ \forall z\bigg{(}\Big{(}\bigvee_{i=1}^{k}E(z,x_{i})\Big{)}\rightarrow\Big{(}\bigvee_{i=1}^{k}z=x_{i}\Big{)}\bigg{)}.$
Note that if $p(\bar{x})$ is a self-contained closure type over $\sigma\supseteq\tau$, then $p(\bar{x})$ is equivalent to a formula of the form $p^{*}(\bar{x})\wedge\text{ `$\bar{x}$ is closed'}$’, where $p^{*}(\bar{x})$ is an atomic type over $\sigma$.
Lemma 5.12.
Let $\tau\subseteq\sigma$ and let $p(\bar{x})$ be a closure type over $\sigma$. Then there is a self-contained closure type over $\sigma$ $p^{*}(\bar{x},\bar{y})$ such that if $\mathcal{A}$ is a $\sigma$-structure that expands a tree $\mathcal{T}$, then
$\models_{tree}\forall\bar{x}\big{(}p(\bar{x})\leftrightarrow\exists!\bar{y}p^{*}(\bar{x},\bar{y})\big{)}$
where ‘$\exists!\bar{y}$’ means ‘there exists unique $\bar{y}$’. It follows that $|p(\mathcal{A})|=|p^{*}(\mathcal{A})|$ and if $\bar{x}=\bar{u}\bar{v}$ and $\bar{a}\in A^{|\bar{u}|}$, then $|p(\bar{a},\mathcal{A})|=|p^{*}(\bar{a},\mathcal{A})|$.
Proof. Let $p(\bar{x})$ be a closure type over $\sigma$. Then $p(\bar{x})$ is equivalent to a formula of the form
$\exists\bar{y}\big{(}\varphi(\bar{x},\bar{y})\ \wedge\ `\bar{x}\bar{y}\text{ is closed'}\big{)}$’ )
where $\varphi(\bar{x},\bar{y})$ is an atomic type over $\sigma$, $\bar{y}=(y_{1},\ldots,y_{m})$, and
$\varphi(\bar{x},\bar{y})\models_{tree}\bigwedge_{i=1}^{m}y_{i}\in\mathrm{cl}(\bar{x}).$
Let $p^{*}(\bar{x},\bar{y})$ be the formula
$\varphi(\bar{x},\bar{y})\ \wedge\ `\bar{x}\bar{y}\text{ is closed'}.$’ .
Then $p^{*}(\bar{x},\bar{y})$ is a closure type over $\sigma$ with empty sequence of quantifiers in the beginning. Since $p^{*}(\bar{x},\bar{y})$ implies that ‘$\bar{x}\bar{y}$ is closed’ it follows that $p^{*}$ is self-contained. It is evident that if $\mathcal{A}$ is a $\sigma$-structure then $\mathcal{A}\models\forall\bar{x}\big{(}p(\bar{x})\leftrightarrow\exists\bar{y}p^{*}(\bar{x},\bar{y})\big{)}$. Suppose, moreover, that $\mathcal{A}$ expands a tree and $\mathcal{A}\models p^{*}(\bar{a},\bar{b})$. Then, for every $b\in\mathrm{rng}(\bar{b})$, either $b\in\mathrm{rng}(\bar{a})$ or there is a directed path of a particular length, say $l_{b}$, from $b$ to some $a_{b}\in\mathrm{rng}(\bar{a})$. In the later case, since $\mathcal{A}\negthickspace\upharpoonright\negthickspace\tau$ is a tree, it follows that $b$ is the unique element such that there is a directed path from $b$ to $a_{b}$ of length $l_{b}$. Therefore $\bar{b}$ is the unique tuple such that $\mathcal{A}\models p^{*}(\bar{a},\bar{b})$. $\square$
The following lemma follows straightforwardly from the definition of closure-type, so we omit the proof.
Lemma 5.13.
Suppose that $\tau\subseteq\sigma$ and $p(\bar{x})$ is a closure-type over $\sigma$ where $\bar{x}=(x_{1},\ldots,k)$.
(i) For all distinct $i,j\in\{1,\ldots,k\}$ either $p(\bar{x})\models_{tree}$$x_{j}$ is an ancestor of $x_{i}$”, or $p(\bar{x})\models_{tree}$$x_{j}$ is not an ancestor of $x_{i}$”.
(ii) For every subsequence $\bar{y}$ of $\bar{x}$ and every $x_{i}\in\mathrm{rng}(\bar{x})\setminus\mathrm{rng}(\bar{y})$, either $p(\bar{x})\models_{tree}x_{i}\in\mathrm{cl}(\bar{y})$, or $p(\bar{x})\models_{tree}x_{i}\notin\mathrm{cl}(\bar{y})$.
With the above lemma the following definition makes sense.
Definition 5.14.
Let $\tau\subseteq\sigma$ and let $p(\bar{x},\bar{y})$ be a closure type over $\sigma$, where $\bar{y}=(y_{1},\ldots,y_{k})$. We call $p(\bar{x},\bar{y})$ $\bar{y}$-independent if, for all $i=1,\ldots,k$, $p(\bar{x},\bar{y})\models_{tree}y_{i}\notin\mathrm{cl}(\bar{x})$.
In the next lemma, and later, the expression ‘$\mathrm{rng}(\bar{y})\subseteq\mathrm{cl}(\bar{x})$’ means that ‘for all $y_{i}\in\mathrm{rng}(\bar{y})$, $y_{i}\in\mathrm{cl}(\bar{x})$’.
Lemma 5.15.
Suppose that $\tau\subseteq\sigma$ and that $p(\bar{x},\bar{y},\bar{z})$ is a closure type over $\sigma$ such that $p(\bar{x},\bar{y},\bar{z})\models\mathrm{rng}(\bar{y})\subseteq\mathrm{cl}(\bar{x})$ and $\bar{z}$ is nonempty. Suppose that $\mathcal{A}$ is a $\sigma$-structure that expands a tree and $\mathcal{A}\models p(\bar{a},\bar{b},\bar{c})\wedge p(\bar{a},\bar{b}^{\prime},\bar{c}^{\prime})$ where $\bar{a}\in A^{|\bar{x}|}$, $\bar{b},\bar{b}^{\prime}\in A^{|\bar{y}|}$, and $\bar{c},\bar{c}^{\prime}\in A^{|\bar{z}|}$. Then $\bar{b}=\bar{b}^{\prime}$ and hence $|p(\bar{a},\mathcal{A})|=|p(\bar{a},\bar{b},\mathcal{A})|$.
Proof. Suppose that $\mathcal{A}\models p(\bar{a},\bar{b},\bar{c})\wedge p(\bar{a},\bar{b}^{\prime},\bar{c}^{\prime})$, where $\bar{b}=(b_{1},\ldots,b_{k})$ and $\bar{b}^{\prime}=(b^{\prime}_{1},\ldots,b^{\prime}_{k})$, and $p(\bar{x},\bar{y},\bar{z})\models\mathrm{rng}(\bar{y})\subseteq\mathrm{cl}(\bar{x})$. The last assumption implies that for each $i=1,\ldots,k$ there are $j_{i}$ and $l_{i}$ such that there is a directed path from $b_{i}$ to $a_{j_{i}}$ of length $l_{i}$ and a directed path from $b^{\prime}_{i}$ to $a_{j_{i}}$ of length $l_{i}$. As $\mathcal{A}\negthickspace\upharpoonright\negthickspace\tau$ is a tree it follows that $b_{i}=b^{\prime}_{i}$. $\square$
Remark 5.16.
Suppose that $p(\bar{x},\bar{y})$ is a closure-type over $\tau$. By Lemma 5.12, there is a self-contained closure type $p^{*}(\bar{x},\bar{y},\bar{z})$ over $\tau$ such that
$\models_{tree}\forall\bar{x},\bar{y}\big{(}p(\bar{x},\bar{y})\leftrightarrow\exists!\bar{z}p^{*}(\bar{x},\bar{y},\bar{z})\big{)}.$
Recall that by the definition of closure-type $p^{*}\models u\neq v$ whenever $u$ and $v$ are different entries from the sequence of variables $\bar{x}\bar{y}\bar{z}$. Let $\bar{y}=(y_{1},\ldots,y_{k})$ and $\bar{z}=(z_{1},\ldots,z_{l})$. Lemma 5.13 implies that, by reordering $\bar{y}$ and $\bar{z}$ if necessary we may (without loss of generality) assume that, for some $1\leq m_{0}\leq k$, $1\leq m_{1}\leq l$ and $m_{1}\leq m_{2}\leq l$, $p^{*}(\bar{x},\bar{y},\bar{z})\models_{tree}y_{i}\notin\mathrm{cl}(\bar{x})$ if and only if $i>m_{0}$, $p^{*}(\bar{x},\bar{y},\bar{z})\models_{tree}z_{i}\in\mathrm{cl}(\bar{y})$ if and only if $i>m_{1}$, and for $i>m_{1}$, $p^{*}(\bar{x},\bar{y},\bar{z})\models_{tree}z_{i}\notin\mathrm{cl}(\bar{x})$ if and only if $i>m_{2}$.
It follows that for every tree $\mathcal{T}$ and all $\bar{a}\in T^{|\bar{x}|}$ and $\bar{b}\in T^{|\bar{y}|}$, if $\mathcal{T}\models p(\bar{a},\bar{b})$ (so $\mathcal{T}\models p^{*}(\bar{a},\bar{b},\bar{c})$ for some unique $\bar{c}\in T^{|\bar{z}|}$), then $|\mathrm{cl}_{\mathcal{T}}(\bar{b})\setminus\mathrm{cl}_{\mathcal{T}}(\bar{a})|=(k-m_{0})+(l-m_{2})$. Thus the number $|\mathrm{cl}_{\mathcal{T}}(\bar{b})\setminus\mathrm{cl}_{\mathcal{T}}(\bar{a})|$ depends only on $p(\bar{x},\bar{y})$.
Definition 5.17.
Let $\tau\subseteq\sigma$, let $p(\bar{x},\bar{y})$ be a closure type over $\sigma$, and let $p_{\tau}(\bar{x},\bar{y})$ be the restriction of $p$ to $\tau$. We define the $\bar{y}$-rank of p, denoted $\mathrm{rank}_{\bar{y}}(p)$, to be the number
$\mathrm{rank}_{\bar{y}}(p)=|\mathrm{cl}_{\mathcal{T}}(\bar{b})\setminus\mathrm{cl}_{\mathcal{T}}(\bar{a})|$
where $\mathcal{T}$ is any tree and $\bar{a}\in T^{|\bar{x}|}$ and $\bar{b}\in T^{|\bar{y}|}$ are any tuples (of the specified lengths) such that $\mathcal{T}\models p_{\tau}(\bar{a},\bar{b})$. By Remark 5.16, this definition depends only on $p_{\tau}(\bar{x},\bar{y})$.
Remark 5.18.
It follows from Remark 5.16 and Definition 5.17 that if $p(\bar{x},\bar{y})$ is a closure type over $\sigma$ (where $\tau\subseteq\sigma$) with respect to a tree $\mathcal{T}$, then $\mathrm{rank}_{\bar{y}}(p)=0$ if and only if $p(\bar{x},\bar{y})\models_{tree}\mathrm{rng}(\bar{y})\subseteq\mathrm{cl}(\bar{x})$.
Remark 5.19.
(The relevance of self-contained closure types) Suppose that $\tau\subseteq\sigma$ and that $p(\bar{x},\bar{y})$ is a closure type over $\sigma$ such that $\mathrm{rank}_{\bar{y}}(p)>0$. By Lemma 5.12 there is a sequence of variables $\bar{z}$ which extends $\bar{y}$ and a self-contained closure type over $\sigma$ $p^{*}(\bar{x},\bar{z})$ such that $|p(\bar{a},\mathcal{A})|=|p^{*}(\bar{a},\mathcal{A})|$ for all $\sigma$-structures $\mathcal{A}$ such that $\mathcal{A}\negthickspace\upharpoonright\negthickspace\tau$ is a tree and all $\bar{a}\in A^{|\bar{x}|}$. By Lemma 5.15 we can write $\bar{z}=\bar{u}\bar{v}$ (if $\bar{z}$ is reordered if necessary) so that $p^{*}$ is $\bar{v}$-independent and, for all $\bar{a}\in A^{|\bar{x}|}$ and $\bar{b}\in A^{|\bar{u}|}$ such that $p^{*}(\bar{a},\bar{b},\mathcal{A})\neq\emptyset$, $|p^{*}(\bar{a},\mathcal{A})|=|p^{*}(\bar{a},\bar{b},\mathcal{A})|$ and hence $|p(\bar{a},\mathcal{A})|=|p^{*}(\bar{a},\bar{b},\mathcal{A})|$. (That $\bar{v}$ is nonempty follows from the assumption that $\mathrm{rank}_{\bar{y}}(p)>0$.) Note that the root of $\mathcal{A}\negthickspace\upharpoonright\negthickspace\tau$ belongs to $\mathrm{rng}(\bar{a}\bar{b})$, so $\bar{x}\bar{z}$ is a nonempty sequence (even if $\bar{x}$ is empty). Also observe that if $p(\bar{x},\bar{y})$ is a complete closure type over $\sigma$ then $p^{*}(\bar{x},\bar{z})$ will be a complete closure type over $\sigma$. This justifies that we will, in some technical proofs that will follow, work only with closure types over $\sigma$ $p(\bar{x},\bar{y})$ that are self-contained and $\bar{y}$-independent and where $\bar{x}$ is nonempty.
Remark 5.20.
(Decomposing a closure type over $\sigma$ of rank $\geq 2$) Let $\tau\subseteq\sigma$ and suppose that $p(\bar{x},\bar{y})$ is a self-contained and $\bar{y}$-independent closure type over $\sigma$ such that $\bar{x}$ is nonempty and $\mathrm{rank}_{\bar{y}}(p)=r+1$ where $r\geq 1$. Then $|\bar{y}|=r+1$, $p(\bar{x},\bar{y})\models_{tree}$$\bar{x}$ contains the root”, and there is $u\in\mathrm{rng}(\bar{y})$ such that
$p(\bar{x},\bar{y})\models_{tree}\text{ ``$u$ is the child of a member of $\bar{x}$''.}$ ”.
Let $\bar{v}$ be the sequence obtained from $\bar{y}$ by removing $u$. By reordering $\bar{y}$ if necessary we may assume that $\bar{y}=u\bar{v}$. Let $q(\bar{x},u)=p\negthickspace\upharpoonright\negthickspace\bar{x}u$. By the choice of $u$ if follows that $q(\bar{x},u)$ is self-contained and $u$-independent, and that $q(\bar{x},u)\models u\notin\mathrm{cl}(\bar{x})$. Hence $\mathrm{rank}_{u}(q)=1$. By the choice of $u$, $p(\bar{x},u,\bar{v})\models$$\mathrm{cl}(u)\cap\mathrm{rng}(\bar{v})=\emptyset$”. Since $p$ is $\bar{y}$-independent and $\bar{v}$ is a subsequence of $\bar{y}$ it follows that $p$ is $\bar{v}$-independent.
Suppose that $\mathcal{T}$ is a tree and $\mathcal{T}\models p(\bar{a},b,\bar{c})$ where $\bar{a}\in T^{|\bar{x}|}$, $b\in T$ and $\bar{c}\in T^{|\bar{v}|}$. Then, by the choice of $u$ and since $p$ is self-contained and $\bar{y}$-independent, $\mathrm{cl}_{\mathcal{T}}(\bar{c})\setminus\mathrm{cl}_{\mathcal{T}}(\bar{a}b)=\mathrm{rng}(\bar{c})$ where $|\mathrm{rng}(\bar{c})|=|\bar{c}|$. Hence $\mathrm{rank}_{\bar{v}}(p)=|\bar{c}|=|\bar{v}|=r$.
Our main results will show that every $PLA^{*}$-formula that satisfies certain conditions is asymptotically equivalent to a closure-basic formula as defined below.
Definition 5.21.
Let $\tau\subseteq\sigma$. A closure-basic formula over $\sigma$ is a formula of the form
$\bigwedge_{i=1}^{k}(\varphi_{i}(\bar{x})\rightarrow c_{i})$ where for each $i=1,\ldots,k$, $\varphi_{i}(\bar{x})$ is a complete closure type over $\sigma$ and $c_{i}\in[0,1]$.
At one point of the proof of the main results we will use induction on the complexity on formulas, and the base case of the induction uses the following result.
Lemma 5.22.
(i) Suppose that $\varphi_{1}(\bar{x}),\ldots,\varphi_{k}(\bar{x})$ are closure-basic formulas over $\sigma$ and that $\mathsf{C}:[0,1]^{k}\to[0,1]$. Then the formula $\mathsf{C}(\varphi_{1}(\bar{x}),\ldots,\varphi_{k}(\bar{x}))$ is equivalent to a closure-basic formula over $\sigma$.
(ii) If $\varphi(\bar{x})\in PLA^{*}(\sigma)$ is aggregation-free then it is equivalent to a closure-basic formula.
Proof. (i) Suppose that $\varphi_{i}(\bar{x})$, $i=1,\ldots,k$, is a closure-basic formula over $\sigma$. Let $q_{1}(\bar{x}),\ldots,q_{m}(\bar{x})$ enumerate, up to logical equivalence, all complete closure-basic types over $\sigma$ in the free variables $\bar{x}$. Suppose that $\mathcal{A}$ is a finite $\sigma$-structure abd $\bar{a}\in A^{|\bar{x}|}$. Observe that for each $i$ the value $\mathcal{A}(\varphi_{i}(\bar{a}))$ depends only on which $q_{j}(\bar{x})$ the sequence $\bar{a}$ satisfies. So let $c_{i,j}=\mathcal{A}(\varphi_{i}(\bar{a}))$ if $\mathcal{A}\models q_{j}(\bar{a})$. Then let $d_{j}=\mathsf{C}(c_{1,j},\ldots,c_{k,j})$ for $j=1,\ldots,m$. Now $\varphi(\bar{x})$ is equivalent in to the closure-basic formula $\bigwedge_{j=1}^{m}(q_{j}(\bar{x})\to d_{j})$.
(ii) Let $\varphi(\bar{x})\in PLA^{*}(\sigma)$ be aggregation-free. The proof proceeds by induction on the number of connectives in $\varphi$. If the number of connectives is 0 then $\varphi(\bar{x})$ can be a constant from $[0,1]$, or it can have the form $R(\bar{x}^{\prime})$ for some $R\in\sigma$ and subsequence $\bar{x}^{\prime}$ of $\bar{x}$, or it can have the form $u=v$ for some $u,v\in\mathrm{rng}(\bar{x})$. It is easy to verify that in each case $\varphi(\bar{x})$ is equivalent to a closure-basic formula. The inductive step follows from part (i) of this lemma. $\square$
6. The base sequence of trees, expansions of them, and probabilities
We adopt the assumptions from the previous section, so $\sigma$ is a finite relational signature, $\tau\subseteq\sigma$, and $\tau=\{E\}$ where $E$ is a binary relation symbol. We consider a base sequence $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ where each $\mathcal{T}_{n}$ is a tree, so in particular $\mathcal{T}_{n}$ is a $\tau$-structure, and $|T_{n}|\to\infty$ as $n\to\infty$. Then $\mathbf{W}_{n}$ will be the set of all expansions of $\mathcal{T}_{n}$ to $\sigma$, and a probability distribution $\mathbb{P}_{n}$ will be defined on $\mathbf{W}_{n}$, via the concept of $PLA^{*}(\sigma)$-network. Our goal is to identify some constraints on $\mathbf{T}$ and on the $PLA^{*}(\sigma)$-network such that under these constraints we can prove results about asymptotic equivalence of complex formulas to simpler (closure basic) formulas, and derive convergence results.
For example, we could fix an integer $\Delta\geq 1$ and let $\mathcal{T}_{n}$ be a tree in which every element (that is, vertex) has at most $\Delta$ children. However, this case is covered by the context studied by Koponen in [20]. So we will allow, in fact require, that there is a function, say $g_{1}$, such that $\lim_{n\to\infty}g_{1}(n)=\infty$ and every nonleaf of $\mathcal{T}_{n}$ has at least $g_{1}(n)$ children. This condition alone does not imply the kind of results that we are looking for. So, in addition, we will assume that there is an integer $\Delta\geq 1$ such that, for all $n$, the height of $\mathcal{T}_{n}$ is at most $\Delta$. We justify this condition by the fact that such trees can represent hierarchies consisting of a fixed number of levels and where each “group” of objects in such a hierarchy has many more members than the number of levels. Moreover, such hierarchies are present in many situations and may influence relations which are modelled/expressed by probabilistic methods.
The perhaps simplest sequence $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ for which the main results of this article hold is obtained by fixing some integer $\Delta\geq 1$ and letting $\mathcal{T}_{n}$ be a tree of height $\Delta$ such that every leaf is on level $\Delta$ and all nonleaves have exactly $n$ children. But in order to increase the applicability of our results we wish prove results for more general sequences $\mathbf{T}$. For example, we want to allow different elements (that is, vertices) of $\mathcal{T}_{n}$ to have different numbers of children. However, if we put no constraints on the relative number of children that different elements of a tree can have, then we do not get the results that we aim for. The next couple of examples illustrate this and also motivate the constraints that we will impose on the sequence $\mathbf{T}$ of trees.
Example 6.1.
Let $\sigma=\tau\cup\{R\}$ where $R$ has arity 1. For even $n\in\mathbb{N}^{+}$ let $\mathcal{T}_{n}$ be a tree of height $2$ such that the root has $n$ children, half of the children of the root have $n$ children, and the rest of the children of the root have no child at all. For odd $n\in\mathbb{N}^{+}$ let $\mathcal{T}$ be a tree of height $2$ such that the root has $n$ children, $\lfloor n/3\rfloor$ children of the root have $n$ children, and the rest of the children of the root have no child at all. Let $\mathbf{W}_{n}$ the set of all $\sigma$-structures that expand $\mathcal{T}_{n}$ and let $\mathbb{P}_{n}$ be the uniform probability distribution on $\mathbf{W}_{n}$, or equivalently, let $\mathbb{P}_{n}$ be such that for every $a\in T_{n}$, the probability that $\mathcal{A}\models R(a)$ for a random $\mathcal{A}\in\mathbf{W}_{n}$ is $1/2$, independently of whether $\mathcal{A}\models R(b)$ for $b\neq a$.
Let $q(x)$ be a formula which expresses that “$x$ is a child of the root”, let $p(x,y)$ be a formula which expresses that “$q(x)$ and $y$ is a child of $x$”, and let $\varphi(x)$ be the formula
$\mathrm{am}\big{(}p(x,y)\wedge R(x)\wedge R(y):y:p(x,y)\big{)}.$
Suppose that $n$ is a large and even. By Lemma 2.1 and since $|T_{n}|$ is bounded by a polynomial in $n$ we can argue (somewhat informally) like this. Let $a\in T_{n}$ be a child of the root such that such that $a$ is a child of the root with a child, hence $n$ children.
With high probability (approaching 1 as $n\to\infty$), if $\mathcal{A}\in\mathbf{W}_{n}$ is chosen at random and if we condition on $\mathcal{A}\models R(a)$, then $\mathcal{A}(\varphi(a))\approx 1/2$. If we condition on $\mathcal{A}\not\models R(a)$ then $\mathcal{A}(\varphi(a))=0$. If $a$ is a child of the root without any child, then (by the semantics of $PLA^{*}$) $\mathcal{A}(\varphi(a))=0$ (no matter whether $\mathcal{A}\models R(a)$ or not). The probability that $\mathcal{A}\models R(a)$ is $1/2$. If we put this together we get $\mathcal{A}\big{(}\mathrm{am}(\varphi(x):x:q(x))\big{)}\approx(\frac{n}{2}\cdot\frac{1}{2}+\frac{n}{2}\cdot 0)/n=1/4$. So if $\chi=\mathrm{am}(\varphi(x):x:q(x))$ then, with high likelihood, $\mathcal{A}(\chi)\approx 1/4$.
By a similar argument, for large odd $n$, if $\mathcal{A}\in\mathbf{W}_{n}$ is chosen at random, then, with high likelihood, $\mathcal{A}(\chi)\approx 1/6$. It follows that, under the given assumptions, we do not have a convergence result as in part (ii) of Theorem 8.1. It is not too difficult to show that under the same assumptions part (i) of Theorem 8.1 does not hold either. In this example we can blame the failure on the fact that we allow that different leaves in $\mathcal{T}_{n}$ are on different levels. So in the next example we stipulate that all leaves are on the same level (which turns out to be insufficient for convergence).
Example 6.2.
Let $\sigma$ be as in Example 6.1. For odd $n\in\mathbb{N}^{+}$ let $\mathcal{T}_{n}$ be a tree of height $2$ such that the root has $n$ children, one child of the root has $2n^{3}$ children, and all other children of the root have $n$ children. For even $n\in\mathbb{N}^{+}$ let $\mathcal{T}_{n}$ be a tree of height $2$ such that the root has $n$ children, two children of the root has $n^{3}$ children, and all other children of the root have $n$ children. Let $\mathbf{W}_{n}$ be the set of all expansions of $\mathcal{T}_{n}$ to $\sigma$ and let $\mathbb{P}_{n}$ be the uniform probability distribution on $\mathbf{W}_{n}$.
Suppose that $n$ is odd and large. Let $b_{n}\in T_{n}$ be the unique child of the root that has $2n^{3}$ children. Note that $\mathcal{T}_{n}$ has $(n-1)n+2n^{3}$ leaves and that $2n^{3}$ of these leaves are children of $b_{n}$. As $n$ is large, almost all leaves are children of $b_{n}$. Let $q(\bar{x})$ and $p(\bar{x},\bar{y})$ be as in Example 6.1. Hence $\exists xp(x,y)$ expresses, in every $\mathcal{T}_{n}$, that “$y$ is a leaf”. Let $\varphi$ be
$\mathrm{am}\big{(}\exists x(p(x,y)\wedge R(x)\wedge R(y)):y:\exists xp(x,y)\big{)}.$
For a random $\mathcal{A}\in\mathbf{W}_{n}$ the probability that $\mathcal{A}\models R(b_{n})$ is $1/2$, and conditioned on $\mathcal{A}\models R(b_{n})$ then, with high probability, $\mathcal{A}(\varphi)\approx 1/2$, and conditioned on $\mathcal{A}\not\models R(b_{n})$, $\mathcal{A}(\varphi)\approx 0$. So if $n$ is odd and large then, with probability roughly $1/2$ we have $\mathcal{A}(\varphi)\approx 1/2$, and with probability roughly $1/2$ we have $\mathcal{A}(\varphi)\approx 0$.
By reasoning in a similar way it also follows that if $n$ is even and large then, with probability roughly $1/4$ we have $\mathcal{A}(\varphi)\approx 1/2$, with probability roughly $1/2$ we have $\mathcal{A}(\varphi)\approx 1/4$, and with probability roughly $1/4$ we have $\mathcal{A}(\varphi)\approx 0$. It follows that we do not have a convergence result as in part (ii) of Theorem 8.1. One can also show that neither does part (i) of the same theorem hold.
In Example 6.2 an obstacle to convergence is that we have a varying and bounded number (either one or two) of children of the root with many more vertices than the other children of the root. If we have trees $\mathcal{T}_{n}$ such that all leaves are on level 2, the root has $n$ children, and, for every child $a$ of the root of $\mathcal{T}_{n}$, there is $k_{n}$, where $\lim_{n\to\infty}k_{n}=\infty$ but not too slow, and there are at least $k_{n}$ children $a^{\prime}$ of the root such that the number of children of $a^{\prime}$ is not too different from the number of children of $a$, then we do not encounter the same problem.
However, in general it is not sufficient to just have some (not too small) lower bound on the number of siblings in $\mathcal{T}_{n}$ with a similar number of children. But we actually need, for every fixed tree $\mathcal{T}^{\prime}$, to put a (not too large) lower bound on the number of siblings $a^{\prime}$ of (any vertex) $a$ such that $\mathsf{N}_{\mathcal{T}_{n}}(a^{\prime},\mathcal{T}^{\prime})$ is not too different from $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})$, where we recall that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})$ is the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $a$ and isomorphic to $\mathcal{T}^{\prime}$. (It is not hard to modify Example 6.2, by considering trees in which all leaves are on level 3, so that it shows that this stronger constraint is necessary.) Also, we will assume that there is a polynomial upper bound to the number of children that any member of $\mathcal{T}_{n}$ can have. This will allow us to use Corollary 2.2 in such a way that the main results follow.
The above considerations motivate the following assumption:
Assumption 6.3.
(The base sequence of trees)
  1. (1)
    $\Delta\in\mathbb{N}^{+}$,
  2. (2)
    $g_{1},g_{2},g_{3},g_{4}$ are functions from $\mathbb{N}$ to the positive reals,
    1. (a)
      for $i=1,2,3,4$, $\lim_{n\to\infty}g_{i}(n)=\infty$,
    2. (b)
      for every $\alpha\in\mathbb{R}$, $\lim_{n\to\infty}(g_{3}(n)-\alpha\ln(n))=\lim_{n\to\infty}(g_{1}(n)-g_{3}(n))=\infty$,
    3. (c)
      $\lim_{n\to\infty}\frac{g_{2}(n)}{g_{1}(n)}=0$, and
    4. (d)
      $g_{4}$ is a polynomial,
  3. (3)
    $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ and each $\mathcal{T}_{n}$ is a tree such that
    1. (a)
      the height of $\mathcal{T}_{n}$ is $\Delta$ and all leaves of $\mathcal{T}$ are on level $\Delta$,
    2. (b)
      every nonleaf has at least $g_{1}(n)$ children and at most $g_{4}(n)$ children, and
    3. (c)
      for every tree $\mathcal{T}^{\prime}$ of height at least 1, if $n$ is sufficently large then the following holds: if $a\in T_{n}$ is not the root and $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})>0$, then $a$ has at least $g_{3}(n)$ siblings $b$ such that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})-g_{2}(n)<\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})+g_{2}(n)$.
From now until and including Section 7 let $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ be a sequence of trees that satisfies Assumption 6.3.
Remark 6.4.
(i) Note that (by Assumption 6.3) $|T_{n}|\leq g_{4}(n)^{\Delta}$ for all $n$, where $g_{4}(x)^{\Delta}$ is a polynomial function.
(ii) Suppose that $\mathcal{T}^{\prime}$ is a tree of heigt $\geq 1$. Let $a\in T_{n}$ and suppose that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})>0$. Since every nonleaf of $\mathcal{T}_{n}$ has at least $g_{1}(n)$ children it follows that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})\geq g_{1}(n)$ if $n$ is large enough. Moreover, if $n$ is large enough, then whether $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})>0$ or not depends only on which level $a$ belongs to, because all leaves of $\mathcal{T}_{n}$ are on level $\Delta$ and all nonleaves have at least $g_{1}(n)$ children where $g_{1}(n)\to\infty$.
One part of part 2(b) of Assumption 6.3 will be used in the form of the following lemma.
Lemma 6.5.
Let $k\in\mathbb{N}^{+}$. For every $\alpha>0$, if $n$ is sufficiently large then $\frac{n^{k}}{e^{\alpha g_{3}(n)}}\leq e^{-\frac{1}{2}\alpha g_{3}(n)}$.
Proof. Let $k\in\mathbb{N}^{+}$. We first show that for every $\alpha>0$, $\lim_{n\to\infty}\frac{n^{k}}{e^{\alpha g_{3}(n)}}=0$. Since $\ln(x)$ is bounded on the interval $[a,b]$ for all $a,b\in\mathbb{R}^{+}$ such that $a<b$, and $\lim_{x\to\infty}\ln(x)=\infty$ it suffices to show that, for all $k\in\mathbb{N}^{+}$ and $\alpha>0$, $\lim_{n\to\infty}\ln\frac{e^{\alpha g_{3}(n)}}{n^{k}}=\infty$. We have $\lim_{n\to\infty}\ln\frac{e^{\alpha g_{3}(n)}}{n^{k}}=\lim_{n\to\infty}(\ln e^{\alpha g_{3}(n)}-\ln(n^{k}))=\lim_{n\to\infty}(\alpha g_{3}(n)-k\ln(n))=\lim_{n\to\infty}\alpha(g_{3}(n)-\frac{k}{\alpha}\ln(n))=\infty,$ where the last identity follows from Assumption 6.3. For all $k\in\mathbb{N}^{+}$ and $\alpha>0$ we now get $\frac{n^{k}}{e^{\alpha g_{3}(n)}}\leq\frac{n^{2k}}{e^{\alpha g_{3}(n)}}=\frac{n^{2k}}{e^{\frac{1}{2}\alpha g_{3}(n)}}\cdot\frac{1}{e^{\frac{1}{2}\alpha g_{3}(n)}}\leq e^{-\frac{1}{2}\alpha g_{3}(n)}$ if $n$ is large enough, by what we just proved. $\square$
Definition 6.6.
For all $n\in\mathbb{N}^{+}$ let $\mathbf{W}_{n}$ be the set of all $\sigma$-structures $\mathcal{A}$ such that $\mathcal{A}\negthickspace\upharpoonright\negthickspace\tau=\mathcal{T}_{n}$.
Note that if $\sigma=\tau$ then $\mathbf{W}_{n}=\{\mathcal{T}_{n}\}$.
Definition 6.7.
Let $\varphi(\bar{x})\in PLA^{*}(\sigma)$. We call $\varphi(\bar{x})$ cofinally satisfiable if there are infinitely many $n\in\mathbb{N}^{+}$ such that there is $\bar{a}\in(T_{n})^{|\bar{x}|}$ and $\mathcal{A}\in\mathbf{W}_{n}$ such that $\mathcal{A}(\varphi(\bar{a}))=1$.
Suppose that $p(\bar{x})$ is a closure-type over $\sigma$ and let $p_{\tau}(\bar{x})=p\negthickspace\upharpoonright\negthickspace\tau$. If $p_{\tau}(\bar{x})$ is satisfied in $\mathcal{T}_{n}$ for some $n$, then (because of the assumptions on $\mathbf{T}$) it is cofinally satisfiable. Since $\mathbf{W}_{n}$ contains all expansions to $\sigma$ of $\mathcal{T}_{n}$ it follows that if $p_{\tau}(\bar{x})$ is satisified in $\mathcal{T}_{n}$ for some $n$, then $p(\bar{x})$ is cofinally satisfiable.
Definition 6.8.
(i) A $PLA^{*}(\sigma)$-network based on $\tau$ is specified by the following two parts:
  1. (1)
    A DAG $\mathbb{G}$ with vertex set (or domain) $\sigma\setminus\tau$.
  2. (2)
    To each relation symbol $R\in\sigma\setminus\tau$ a formula $\theta_{R}(\bar{x})\in PLA^{*}(\mathrm{par}(R)\cup\tau)$ is associated where $|\bar{x}|$ equals the arity of $R$ and $\mathrm{par}(R)$ is the set of parents of $R$ in the DAG $\mathbb{G}$. We call $\theta_{R}$ the formula associated to $R$ by the $PLA^{*}(\sigma)$-network.
We will denote a $PLA^{*}(\sigma)$-network by the same symbol (usually $\mathbb{G}$, possibly with a sub or superscript) as its underlying DAG.
(ii) Let $\mathbb{G}$ denote a $PLA^{*}(\sigma)$-network based on $\tau$, let $\tau\subseteq\sigma^{\prime}\subseteq\sigma$, and suppose that for every $R\in\sigma^{\prime}$, $\mathrm{par}(R)\subseteq\sigma^{\prime}$. Then the $PLA^{*}(\sigma^{\prime})$-network specified by the induced subgraph of $\mathbb{G}$ with vertex set $\sigma^{\prime}\setminus\tau$ and the associated formulas $\theta_{R}$ for all $R\in\sigma^{\prime}\setminus\tau$ will be called the $PLA^{*}(\sigma^{\prime})$-subnetwork of $\mathbb{G}$ induced by $\sigma^{\prime}$.
Observe that if $\sigma=\tau$ then the DAG of a $PLA^{*}(\sigma)$-network have an empty vertex set and hence there is no probability formula associated to it. Example 6.11 gives an example of a $PLA^{*}(\sigma)$-network.
From now on let $\mathbb{G}$ be a $PLA^{*}(\sigma)$-network based on $\tau$.
Definition 6.9.
(i) If $\sigma=\tau$ then $\mathbb{P}_{n}$, the probability distribution on $\mathbf{W}_{n}$ induced by $\mathbb{G}$, is the unique probability distribution on (the singleton set) $\mathbf{W}_{n}$.
(ii) Now suppose that $\tau$ is a proper subset of $\sigma$. For each $R\in\sigma$ let $\nu_{R}$ denote its arity, and let $\theta_{R}(\bar{x})$, where $|\bar{x}|=\nu_{R}$, be the formula which $\mathbb{G}$ associates to $R$. Suppose that the underlying DAG of $\mathbb{G}$ has height $\rho$. For each $0\leq l\leq\rho$ let $\mathbb{G}_{l}$ be the subnetwork which is induced by $\sigma_{l}=\{R\in\sigma:\text{$R$ is on level $i$ and $i\leq l$}\}$ and note that $\sigma_{\rho}=\sigma$ and $\mathbb{G}_{\rho}=\mathbb{G}$. Also let $\sigma_{-1}=\tau$ and let $\mathbb{P}^{-1}_{n}$ be the unique probability distribution on $\mathbf{W}^{-1}_{n}=\{\mathcal{T}_{n}\}$. By induction on $r$ we define, for every $l=0,1,\ldots,\rho$, a probability distribution $\mathbb{P}^{l}_{n}$ on the set $\mathbf{W}^{l}_{n}=\{\mathcal{A}\negthickspace\upharpoonright\negthickspace\sigma_{l}:\mathcal{A}\in\mathbf{W}_{n}\}$ as follows:
For every $\mathcal{A}\in\mathbf{W}^{l}_{n}$, let $\mathcal{A}^{\prime}=\mathcal{A}\negthickspace\upharpoonright\negthickspace\sigma_{l-1}$ and
$\mathbb{P}^{l}_{n}(\mathcal{A})\ =\ \mathbb{P}^{l-1}_{n}(\mathcal{A}^{\prime})\prod_{R\in\sigma_{l}\setminus\sigma_{l-1}}\ \prod_{\bar{a}\in R^{\mathcal{A}}}\mathcal{A}^{\prime}\big{(}\theta_{R}(\bar{a})\big{)}\ \prod_{\bar{a}\in\ (T_{n})^{{}^{\nu_{R}}}\ \setminus\ R^{\mathcal{A}}}\big{(}1-\mathcal{A}^{\prime}\big{(}\theta_{R}(\bar{a})\big{)}\big{)}.$
Finally we let $\mathbb{P}_{n}=\mathbb{P}^{\rho}_{n}$ and note that $\mathbf{W}_{n}=\mathbf{W}^{\rho}_{n}$, so $\mathbb{P}_{n}$ is a probability distribution on $\mathbf{W}_{n}$ which we call the probability distribution on $\mathbf{W}_{n}$ induced by $\mathbb{G}$. We also call $\mathbb{P}=(\mathbb{P}_{n}:n\in\mathbb{N}^{+})$ the sequence of probability distributions induced by $\mathbb{G}$.
From Definition 6.9 of $\mathbb{P}^{l}_{n}$ and $\mathbb{P}_{n}$ we immediately get the following:
Lemma 6.10.
Let $\rho\in\mathbb{N}$, $l\in\{0,\ldots,\rho\}$, and let $\sigma_{l}$, $\mathbf{W}^{l}_{n}$, $\mathbb{P}^{l}_{n}$ and $\mathbb{P}_{n}$ be as in Definition 6.9.
(i) Let $R\in\sigma_{l}\setminus\sigma_{l-1}$, $n\in\mathbb{N}^{+}$, $\bar{a}\in(T_{n})^{\nu_{R}}$ (where $\nu_{R}$ is the arity of $R$), and $\mathcal{A}^{\prime}\in\mathbf{W}^{l-1}_{n}$. Then
$\mathbb{P}^{l}_{n}\big{(}\{\mathcal{A}\in\mathbf{W}^{l}_{n}:\mathcal{A}\models R(\bar{a})\}\ |\ \{\mathcal{A}\in\mathbf{W}_{n}:\mathcal{A}\negthickspace\upharpoonright\negthickspace\sigma_{l-1}=\mathcal{A}^{\prime}\}\big{)}\ =\ \mathcal{A}^{\prime}(\theta_{R}(\bar{a})).$
(ii) Let $R_{1},\ldots,R_{t}\in\sigma_{l}\setminus\sigma_{l-1}$ where we allow that $R_{i}=R_{j}$ even if $i\neq j$, $n\in\mathbb{N}^{+}$, $\bar{a}_{i}\in(T_{n})^{\nu_{i}}$ for $i=1,\ldots,t$, and $\mathcal{A}^{\prime}\in\mathbf{W}^{l-1}_{n}$. Suppose that for $i=1,\ldots,t$, $\varphi_{i}(\bar{x})$ is a literal in which $R_{i}$ occurs, and if $i\neq j$ then $\bar{a}_{i}\neq\bar{a}_{j}$ or $R_{i}\neq R_{j}$. Using the probability distribution $\mathbb{P}_{n}^{l}$, the event $\mathbf{E}_{n}^{\varphi_{1}(\bar{a})}=\{\mathcal{A}\in\mathbf{W}^{l}_{n}:\mathcal{A}\models\varphi_{1}(\bar{a})\}$ is independent of the event $\bigcap_{i=2}^{t}\mathbf{E}_{n}^{\varphi_{i}(\bar{a})}$ (where $\mathbf{E}_{n}^{\varphi_{i}(\bar{a})}$ is defined similarly), conditioned on the event $\{\mathcal{A}\in\mathbf{W}_{n}:\mathcal{A}\negthickspace\upharpoonright\negthickspace\sigma_{l-1}=\mathcal{A}^{\prime}\}$.
(iii) Suppose that $\mathbf{X}_{n}\subseteq\mathbf{W}_{n}^{l-1}$ and $\mathbf{Y}_{n}=\{\mathcal{A}\in\mathbf{W}_{n}^{l}:\mathcal{A}\negthickspace\upharpoonright\negthickspace\sigma_{l-1}\in\mathbf{X}_{n}\}$. Then $\mathbb{P}_{n}^{l}(\mathbf{Y}_{n})=\mathbb{P}_{n}^{l-1}(\mathbf{X}_{n})$.
Example 6.11.
Recall the very informal example in the introduction where $\sigma=\tau\cup\{P_{1},P_{2},P_{3},R\}$ where $P_{1},P_{2},P_{3}$ are unary and $R$ binary. A $PLA^{*}(\sigma)$-network based on $\tau$ can (for example) have the following underlying DAG:
The $PLA^{*}(\sigma)$-network must also associate a $PLA^{*}(\sigma)$-formula to each relation symbol $P_{1},P_{2},P_{3},R$. Examples of such formulas will be described informally, and for the descriptions to make sense we imagine that they refer to a (large) tree $\mathcal{T}_{n}$ of height 3 as in Assumption 6.3 (with $\Delta=3$). Moreover, we describe how the “output” of such formulas (a number in $[0,1]$) depends on the properties of its input (a vertex, or pair of vertices).
  1. $\theta_{P_{1}}(x)$ = “If $x$ is a child of the root then $1/3$, else 0.”
  2. $\theta_{P_{2}}(x)$ = “ If $x$ is not on level 2 then 0, else
    • if the parent of $x$ satisfies $P_{1}$ then $2/3$, else $1/3$.”
  3. $\theta_{P_{3}}(x)$ = “If $x$ is not on level 3, then 0, else
    • if the parent of $x$ satisfies $P_{2}$ then $1/3$, else $1/4$.”
  4. $\theta_{R}(x,y)$ = “If $x$ is not on level 3 then 0, else
    • if $y$ is on level 3 then
      • if $y$ belongs to the same subcommunity as $x$, then if $P_{3}(y)$ then $3/4$, else $1/2$, else
      • the proportion of $y$ on level 3 such that $P_{3}(y)$, divided by $3/4$, else
    • if $y$ is on level 2 then
      • if $y$ is a parent of $x$, then if $P_{2}(y)$ then $3/4$, else $1/2$, else
      • the proportion of $y$ on level 2 such that $P_{2}(y)$ among the $y$ that are not a parent of $x$, else
    • if $y$ is on level 1 then the proportion of children of $y$ such that $P_{1}(y)$, divided by $3/2$.”
7. Convergence, balance, and asymptotic elimination of aggregation functions
In this section we consider two technical notions, convergent pairs of formulas, and balanced triples of formulas, which are at the heart of the proofs of the main results. Intuitively speaking, a pair $(\varphi(\bar{x}),\psi(\bar{x}))$ of 0/1-valued formulas from $PLA^{*}(\sigma)$ converges to $\alpha$ if the probability that $\varphi(\bar{x})$ holds, conditioned on $\psi(\bar{x})$ being true, converges to $\alpha$ as $n\to\infty$. And, also intuitively speaking, a triple $(\varphi(\bar{x},\bar{y}),\psi(\bar{x},\bar{y}),\chi(\bar{x}))$ of 0/1-valued formulas is $\alpha$-balanced if with probability tending to 1 as $n\to\infty$, a random $\mathcal{A}\in\mathbf{W}_{n}$ such that $\mathcal{A}\models\chi(\bar{a})$ satisfies that $|\varphi(\bar{a},\mathcal{A})|/|\psi(\bar{a},\mathcal{A})|\approx\alpha$.
Recall the assumptions made on $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$, $\mathbf{W}_{n}$, $\mathbb{G}$, and $\mathbb{P}=(\mathbb{P}_{n}:n\in\mathbb{N}^{+})$ in the previous two sections, which we keep in this section.
7.1. Definitions and some immediate consequences
Definition 7.1.
Suppose that $\varphi(\bar{x})\in PLA^{*}(\sigma)$ and $\bar{a}\in(T_{n})^{|\bar{x}|}$ (for some $n$). Then
$\mathbf{E}_{n}^{\varphi(\bar{a})}=\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\mathcal{A}(\varphi(\bar{a}))=1\big{\}}.$
Definition 7.2.
Let $\varphi(\bar{x}),\psi(\bar{x})\in PLA^{*}(\sigma)$. We say that $\big{(}\varphi,\psi\big{)}$ converges (to $\alpha\in[0,1]$) with respect to $\mathbb{G}$ if, for all $\varepsilon>0$, there exists $n_{0}$, such that for all $n\geq n_{0}$ and all $\bar{a}\in T_{n}^{|\bar{x}|}$,
$\big{|}\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\bar{a})}\ \big{|}\ \mathbf{E}_{n}^{\psi(\bar{a})}\big{)}-\alpha\big{|}\leq\varepsilon\quad\text{if}\quad\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\psi(\bar{a})}\big{)}>0.$
If the stronger condition holds that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\bar{a})}\ \big{|}\ \mathbf{E}_{n}^{\psi(\bar{a})}\big{)}=\alpha$ whenever $n$ is large enough and $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\psi(\bar{a})}\big{)}>0$, then we say that $(\varphi,\psi)$ is eventually constant (with value $\alpha$) with respect to $\mathbb{G}$.
Remark 7.3.
Let $\varphi(\bar{x}),\psi(\bar{x})\in PLA^{*}(\sigma)$. It follows from the definition that if there are only finitely many $n$ such that there is $\bar{a}\in(T_{n})^{|\bar{x}|}$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\psi(\bar{a})}\big{)}>0$ then, for every $\alpha$, $(\varphi,\psi)$ converges to $\alpha$ with respect to $\mathbb{G}$. It also follows from the definition that if there are infinitely many $n$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\psi(\bar{a})}\big{)}>0$ for some $\bar{a}$, but there are only finitely many $n$ such that there is $\bar{a}\in(T_{n})^{|\bar{x}|}$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\bar{a})\wedge\psi(\bar{a})}\big{)}>0$ then $(\varphi,\psi)$ converges to 0 with respect to $\mathbb{G}$.
Therefore, when proving that a pair $(\varphi,\psi)$ converges we will assume that there are infinitely many $n$ such that there is $\bar{a}\in(T_{n})^{|\bar{x}|}$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\bar{a})\wedge\psi(\bar{a})}\big{)}>0$, where the positive probability implies that for some $\mathcal{A}\in\mathbf{W}_{n}$, $\mathcal{A}\models\varphi(\bar{a})\wedge\psi(\bar{a})$.
Definition 7.4.
Let $\varphi(\bar{x},\bar{y}),\psi(\bar{x},\bar{y}),\chi(\bar{x})\in PLA^{*}(\sigma)$.
(i) Let $\alpha\in[0,1]$, $\varepsilon>0$ and let $\mathcal{A}$ be a finite $\sigma$-structure. The triple $(\varphi,\psi,\chi)$ is called $(\alpha,\varepsilon)$-balanced in $\mathcal{A}$ if whenever $\bar{a}\in A^{|\bar{x}|}$ and $\mathcal{A}(\chi(\bar{a}))=1$, then
$(\alpha-\varepsilon)|\psi(\bar{a},\mathcal{A})|\leq|\varphi(\bar{a},\mathcal{A})\cap\psi(\bar{a},\mathcal{A})|\leq(\alpha+\varepsilon)|\psi(\bar{a},\mathcal{A})|.$
(ii) Let $\alpha\in[0,1]$. The triple $(\varphi,\psi,\chi)$ is $\alpha$-balanced with respect to $\mathbb{G}$ if for all $\varepsilon>0$, if
$\mathbf{X}_{n}^{\varepsilon}=\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\text{ $(\varphi,\psi,\chi)$ is $(\alpha,\varepsilon)$-balanced in $\mathcal{A}$}\big{\}}$
then $\lim_{n\to\infty}\mathbb{P}_{n}\big{(}\mathbf{X}_{n}^{\varepsilon}\big{)}=1$. The triple $(\varphi,\psi,\chi)$ is balanced with respect to $\mathbb{G}$ if, for some $\alpha\in[0,1]$, it is $\alpha$-balanced with respect to $\mathbb{G}$. If, in addition, $\alpha>0$ then we call $(\varphi,\psi,\chi)$ positively balanced (with respect to $\mathbb{G}$).
Remark 7.5.
Let $\varphi(\bar{x},\bar{y}),\psi(\bar{x},\bar{y}),\chi(\bar{x})\in PLA^{*}(\sigma)$ be $0/1$-valued. Suppose that $\varphi\wedge\psi\wedge\chi$ is not cofinally satisfiable. We claim that then $(\varphi,\psi,\chi)$ is 0-balanced with respect to $\mathbb{G}$. It suffices to show that, for all sufficiently large $n$, if $\mathcal{A}\in\mathbf{W}_{n}$, $\mathcal{A}\models q(\bar{a})$, and $\mathcal{A}\models\psi(\bar{a},\bar{b})$, then $\mathcal{A}\not\models\varphi(\bar{a},\bar{b})$. But this is immediate from the assumption that $\varphi\wedge\psi\wedge\chi$ is not cofinally satisfiable. For this reason we may assume, when proving results about balanced triples, that the conjunction of the involved formulas is cofinally satisfiable and in particular consistent.
Remark 7.6.
Let $p(\bar{x},\bar{y})$ be a closure type over $\sigma$, let $q(\bar{x})=p\negthickspace\upharpoonright\negthickspace\bar{x}$, and let $p_{\tau}(\bar{x},\bar{y})=p\negthickspace\upharpoonright\negthickspace\tau$ (so $p\models q$ and $p\models p_{\tau}$). Suppose that $\mathrm{rank}_{\bar{y}}(p_{\tau})=0$. We claim that $(p,p_{\tau},q)$ is 1-balanced with respect to $\mathbb{G}$. Since $\mathrm{rank}_{\bar{y}}(p_{\tau})=0$ we have $p_{\tau}(\bar{x},\bar{y})\models\mathrm{rng}(\bar{y})\subseteq\mathrm{cl}(\bar{x})$. (So if $\bar{x}$ is empty then $\bar{y}$ is a single variable, say $y$, and $p_{\tau}(y)$ expresses that $y$ is the root.) It follows that $q(\bar{x})\models\exists\bar{y}p(\bar{x},\bar{y})$ (so if $\bar{x}$ is empty then $q$ is a sentence which expresses which relations the root satisfies), and that if $\bar{a}\in(T_{n})^{|\bar{x}|}$ and $\bar{b}\in(T_{n})^{|\bar{y}|}$, $\mathcal{T}_{n}\models p_{\tau}(\bar{a},\bar{b})$, then $\bar{b}$ is the unique tuple that satisfies $p_{\tau}(\bar{a},\bar{y})$ in $\mathcal{T}_{n}$. So for every $n$, $\bar{a}\in(T_{n})^{|\bar{x}|}$, and $\mathcal{A}\in\mathbf{W}_{n}$, if $\mathcal{A}\models q(\bar{a})$ then $|p(\bar{a},\mathcal{A})|=|p_{\tau}(\bar{a},\mathcal{A})|=1$. Hence $(p,p_{\tau},q)$ is 1-balanced.
7.2. Proofs of results about convergence and balance
We will use induction on the height of the underlying DAG of $\mathbb{G}$ to prove that some pairs of formulas converge and that some triples of formulas are balanced. The base case will not be when the height is 0, but it will be when the height is $-1$, where we recall the convention (from Definition 5.2) that an empty $DAG$ (one without any element/vertex) has height $-1$. The assumption that $\tau=\sigma$ is equivalent to the assumption that the height of the underlying DAG of $\mathbb{G}$ is $-1$. So the base case considers the case when $\tau=\sigma$. The following lemma shows that the induction hypothesis (Assumption 7.8) that we will use holds in the base case when $\tau=\sigma$.
Lemma 7.7.
Suppose that $\tau=\sigma$, so $\mathbf{W}_{n}=\{\mathcal{T}_{n}\}$ for each $n$ and let $\mathbb{G}$ be the unique $PLA^{*}(\sigma)$-network over $\tau$. For each $n$ let $\mathbb{P}_{n}$ be the probability distribution on $\mathbf{W}_{n}$ which is induced by $\mathbb{G}$ (so $\mathbb{P}_{n}(\mathcal{T}_{n})=1$). Then:
  1. (1)
    If $p(\bar{x})$ and $q(\bar{x})$ are closure types over $\tau$ then $(p,q)$ converges with respect to $\mathbb{G}$.
  2. (2)
    Suppose that $p(\bar{x},\bar{y})$, $r(\bar{x},\bar{y})$ and $q(\bar{x})$ are closure types over $\tau$ and suppose that $p\wedge r\wedge q$ is cofinally satisfiable. Then there is $\alpha\in[0,1]$ such that for all $\varepsilon>0$ there is $c>0$ such that for all sufficiently large $n$, if $\bar{a}\in(T_{n})^{|\bar{x}|}$, $B\subseteq r(\bar{a},\mathcal{T}_{n})$, and $|B|\geq g_{3}(n)$, then
    $\mathbb{P}_{n}\big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\text{ if $\mathcal{A}\models q(\bar{a})$ then }(\alpha-\varepsilon)|B|\leq|p(\bar{a},\mathcal{A})\cap B|\leq(\alpha+\varepsilon)|B|\big{\}}\big{)}\ \geq$
    $1-e^{-cg_{3}(n)}.$
Proof. (1) Suppose that $p(\bar{x})$ and $q(\bar{x})$ are closure types over $\tau$. By Remark 7.3 we may assume that $p(\bar{x})\wedge q(\bar{x})$ is satisfiable in some $\mathcal{T}_{n}$. It follows from Definition 5.8 of closure type over $\tau$ that $p$ and $q$ are equivalent, so for all $n$ we have $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p(\bar{a})}\ \big{|}\ \mathbf{E}_{n}^{q(\bar{a})}\big{)}=1$ if $\mathcal{T}_{n}\models q(\bar{a})$, or equivalently, if $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{q(\bar{a})}\big{)}>0$. Hence $(p,q)$ converges to 1 with respect to $\mathbb{G}$. In fact $(p,q)$ is eventually constant with value 1.
(2) Suppose that $p(\bar{x},\bar{y})$, $r(\bar{x},\bar{y})$ and $q(\bar{x})$ are closure types over $\tau$ and suppose that $p\wedge r\wedge q$ is cofinally satisfiable. Then $p$ and $r$ are equivalent and $p$ implies $q$. Suppose that $B\subseteq r(\bar{a},\mathcal{T}_{n})$, and $|B|\geq g_{3}(n)$. As $\lim_{n\to\infty}g_{3}(n)$ this means that $B\neq\emptyset$ if $n$ is large enough. Suppose that $\mathcal{A}\in\mathbf{W}_{n}$, so $\mathcal{A}=\mathcal{T}_{n}$. Also suppose that $\mathcal{T}_{n}\models q(\bar{a})$. Let $\bar{b}\in B$, so $\mathcal{T}_{n}\models r(\bar{a},\bar{b})$ and hence $\mathcal{T}_{n}\models p(\bar{a},\bar{b})$ (as $p$ and $r$ are equivalent). Thus $B\subseteq p(\bar{a},\mathcal{T}_{n})$ and $|B\cap p(\bar{a},\mathcal{A})|=|B|$, so for every $\varepsilon>0$, $(1-\varepsilon)|B|\leq|p(\bar{a},\mathcal{T}_{n})\cap B|\leq(1+\varepsilon)|B|.$ Since $\mathbb{P}_{n}(\mathcal{T}_{n})=1$ we get the conclusion of part (2). $\square$
For the rest of Section 7 we assume that the height of the underlying DAG of $\mathbb{G}$ is $\rho+1$ where $\rho\geq-1$. We also let
$\sigma_{\rho}=\{R\in\sigma:\text{ $R$ is on level $i$ of the underlying graph of $\mathbb{G}$ where $i\leq\rho$}\},$
$\mathbf{W}_{n}^{\rho}=\{\mathcal{A}\negthickspace\upharpoonright\negthickspace\sigma_{\rho}:\mathcal{A}\in\mathbf{W}_{n}\},$
and we let $\mathbb{G}_{\rho}$ be the subnetwork of $\mathbb{G}$ which is induced by $\sigma_{\rho}$. Moreover, we assume the following:
Assumption 7.8.
(Induction hypothesis)
  1. (1)
    For every $R\in\sigma\setminus\tau$, there exists a closure-basic formula $\chi_{R}\in PLA^{*}(\sigma_{\rho})$, such that $\chi_{R}$ and $\theta_{R}$ are asymptotically equivalent with respect to $\mathbb{G}_{\rho}$, where $\theta_{R}$ is the formula of $\mathbb{G}$ associated to $R$.
  2. (2)
    If $p(\bar{x})$ is a self-contained closure type over $\sigma_{\rho}$ and $q(\bar{x})$ is a self-contained closure type over $\tau$, then $(p,q)$ converges with respect to $\mathbb{G}_{\rho}$.
  3. (3)
    Suppose that $p(\bar{x},y)$ and $q(\bar{x})$ are complete closure types over $\sigma_{\rho}$ and that $p_{\tau}(\bar{x},y)$ is a self-contained closure type over $\tau$ such that $\mathrm{rank}_{y}(p_{\tau})=1$. Also suppose that $p\wedge p_{\tau}\wedge q$ is cofinally satisfiable. Then there is $\alpha\in[0,1]$ such that for all $\varepsilon>0$ there is $c>0$ such that for all sufficiently large $n$, if $\bar{a}\in(T_{n})^{|\bar{x}|}$, $B\subseteq p_{\tau}(\bar{a},\mathcal{T}_{n})$, and $|B|\geq g_{3}(n)$, then
    $\mathbb{P}_{n}^{\rho}\big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}^{\rho}:\text{ if $\mathcal{A}\models q(\bar{a})$ then }(\alpha-\varepsilon)|B|\leq|p(\bar{a},\mathcal{A})\cap B|\leq(\alpha+\varepsilon)|B|\big{\}}\big{)}\ \geq$
    $1-e^{-cg_{3}(n)}.$
Remark 7.9.
Suppose (in this remark) that $\sigma_{\rho}=\tau$, or equivalently, that the height of the underlying DAG of $\mathbb{G}$ (with vertex set $\sigma$) is 00, so all $R\in\sigma$ are on level 0. Then part (1) of Assumption 7.8 holds vacuously and Lemma 7.7 implies that parts (2) and (3) hold. So if the height of the underlying DAG of $\mathbb{G}$ is 0 then Assumption 7.8 holds.
The goal of the rest of Section 7 is to prove, in the following order, that
  • (A)
    part (2) of Assumption 7.8 holds if $\sigma_{\rho}$ and $\mathbb{G}_{\rho}$ are replaced by $\sigma$ and $\mathbb{G}$, respectively,
  • (B)
    part (3) of Assumption 7.8 holds if $\sigma_{\rho}$, $\mathbb{P}_{n}^{\rho}$, and $\mathbf{W}_{n}^{\rho}$ are replaced by $\sigma$, $\mathbb{P}_{n}$, and $\mathbf{W}_{n}$, respectively, and
  • (C)
    if $\varphi(\bar{x})\in PLA^{*}(\sigma)$ satisfies certain conditions (stated in Proposition 7.26) then $\varphi$ is asymptotically equivalent (with respect to $\mathbb{G}$) to a closure-basic formula over $\sigma$. It will follow that if $\sigma\subset\sigma^{+}$ and $\mathbb{G}$ is an induced subnetwork of a $PLA^{*}(\sigma^{+})$-network $\mathbb{G}^{+}$ such that $\mathbb{G}^{+}$ has height $\rho+2$, then part (1) of Assumption 7.8 holds if $\sigma$ and $\mathbb{G}$ are replaced by $\sigma^{+}$ and $\mathbb{G}^{+}$, respectively, and $\sigma_{\rho}$ and $\mathbb{G}_{\rho}$ are replaced by $\sigma$ and $\mathbb{G}$, respectively.
Definition 7.10.
For every $n$ and $\mathcal{A}^{\prime}\in\mathbf{W}_{n}^{\rho}$ let $\mathbf{W}^{\mathcal{A}^{\prime}}=\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\mathcal{A}\negthickspace\upharpoonright\negthickspace\sigma_{\rho}=\mathcal{A}^{\prime}\big{\}}$.
7.3. Convergence
In this subsection we prove statement (B) above. We do it in three steps.
Lemma 7.11.
Let $p(\bar{x})$ be a self-contained closure type over $\sigma$ and $p_{\rho}(\bar{x})$ a self-contained complete closure type over $\sigma_{\rho}$. Then $(p,p_{\rho})$ converges with respect to $\mathbb{G}$.
Proof. According to Remark 7.3 we may assume that $p(\bar{x})\wedge p_{\rho}(\bar{x})$ is satisfiable in some $\mathcal{A}\in\mathbf{W}_{n}$ for infinitely many $n$. Let $\varepsilon>0$. For every $R\in\sigma\setminus\sigma_{\rho}$ let $\nu_{R}$ be the arity of $R$ and let $\theta_{R}(x_{1},\ldots,x_{\nu_{R}})$ be the $PLA^{*}(\sigma_{\rho})$-formula associated to $R$ by $\mathbb{G}$. By the first part of Assumption 7.8 there is a closure-basic formula $\chi_{R}(x_{1},\ldots,x_{\nu_{R}})$ over $\sigma_{\rho}$ which is asymptotically equivalent to $\theta_{R}$ with respect to $\mathbb{G}_{\rho}$. Let
$\mathbf{X}_{n}^{\rho,\varepsilon}=\big{\{}\mathcal{A}\in\mathbf{W}_{n}^{\rho}:\text{ for all $R\in\sigma\setminus\sigma_{\rho}$ and all $\bar{a}\in(T_{n})^{\nu_{R}}$, }\big{|}\mathcal{A}(\theta_{R}(\bar{a}))-\mathcal{A}(\chi_{R}(\bar{a}))\big{|}\leq\varepsilon\big{\}},$
and
$\mathbf{X}_{n}^{\varepsilon}=\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\text{ for all $R\in\sigma\setminus\sigma_{\rho}$ and all $\bar{a}\in(T_{n})^{\nu_{R}}$, }\big{|}\mathcal{A}(\theta_{R}(\bar{a}))-\mathcal{A}(\chi_{R}(\bar{a}))\big{|}\leq\varepsilon\big{\}}.$
As $\theta_{R}$ and $\chi_{R}$ are asymptotically equivalent with respect to $\mathbb{G}_{\rho}$ it follows that $\lim_{n\to\infty}$ $\mathbb{P}_{n}^{\rho}$ $\big{(}\mathbf{X}_{n}^{\rho,\varepsilon}\big{)}=1$. By Lemma 6.10 (iii) we also get $\lim_{n\to\infty}\mathbb{P}_{n}\big{(}\mathbf{X}_{n}^{\varepsilon}\big{)}=1$.
Thus, to show that $(p,p_{\rho})$ converges it suffices to show that there is $\alpha\in[0,1]$ such that for every $\varepsilon^{\prime}>0$ there is $\varepsilon>0$ such that for all sufficiently large $n$ and all $\bar{a}\in T_{n}^{|\bar{x}|}$,
$\Big{|}\mathbb{P}_{n}\Big{(}\mathbf{E}_{n}^{p(\bar{a})}\ \big{|}\ \mathbf{E}_{n}^{p_{\rho}(\bar{a})}\cap\mathbf{X}_{n}^{\varepsilon}\Big{)}-\alpha\Big{|}\leq\varepsilon^{\prime}\ \ \text{ if }\ \ \mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p_{\rho}(\bar{a})}\cap\mathbf{X}_{n}^{\varepsilon}\big{)}>0.$
Let $\bar{a}\in(T_{n})^{|\bar{x}|}$ and observe that $\mathbf{E}_{n}^{p_{\rho}(\bar{a})}\cap\mathbf{X}_{n}^{\varepsilon}$ is the disjoint union of sets $\mathbf{W}^{\mathcal{A}^{\prime}}$ as $\mathcal{A}^{\prime}$ ranges over structures in $\mathbf{X}_{n}^{\rho,\varepsilon}$ such that $\mathcal{A}^{\prime}\models p_{\rho}(\bar{a})$. By Lemma 2.3, it suffices to show that there is $\alpha$ such that for all $\varepsilon^{\prime}>0$ there is $\varepsilon>0$ such that for all sufficiently large $n$ and all $\mathcal{A}^{\prime}\in\mathbf{X}_{n}^{\rho,\varepsilon}$ such that $\mathcal{A}^{\prime}\models p_{\rho}(\bar{a})$ and $\mathbb{P}_{n}\big{(}\mathbf{W}^{\mathcal{A}^{\prime}}\big{)}>0$ we have
(7.1) $\Big{|}\mathbb{P}_{n}\Big{(}\mathbf{E}_{n}^{p(\bar{a})}\ \big{|}\ \mathbf{W}^{\mathcal{A}^{\prime}}\Big{)}-\alpha\Big{|}\leq\varepsilon^{\prime}.$
So suppose that $\mathcal{A}^{\prime}\in\mathbf{X}_{n}^{\rho,\varepsilon}$, $\mathcal{A}^{\prime}\models p_{\rho}(\bar{a})$ and $\mathbb{P}_{n}\big{(}\mathbf{W}^{\mathcal{A}^{\prime}}\big{)}>0$. Let $\bar{x}=(x_{1},\ldots,x_{m})$ and $\bar{a}=(a_{1},\ldots,a_{m})$.
Since we assume that $p(\bar{x})$ is a self-contained closure type over $\sigma$ it follows, by the definition of how $\mathbb{G}$ induces $\mathbb{P}_{n}$, that
$\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p(\bar{a})}\ \big{|}\ \mathbf{W}^{\mathcal{A}^{\prime}}\big{)}=$
$\prod_{\begin{subarray}{c}R\in\sigma\setminus\sigma_{\rho}\\ 1\leq i_{1}<\ldots<i_{\nu_{R}}\leq m\\ p(\bar{x})\models R(x_{i_{1}},...,x_{i_{\nu_{R}}})\end{subarray}}\mathcal{A}^{\prime}\big{(}\theta_{R}(a_{i_{1}},\ldots,a_{i_{\nu_{R}}})\big{)}\prod_{\begin{subarray}{c}R\in\sigma\setminus\sigma_{\rho}\\ 1\leq i_{1}<\ldots<i_{\nu_{R}}\leq m\\ p(\bar{x})\models\neg R(x_{i_{1}},...,x_{i_{\nu_{R}}})\end{subarray}}\Big{(}1-\mathcal{A}^{\prime}\big{(}\theta_{R}(a_{i_{1}},...,a_{i_{\nu_{R}}})\big{)}\Big{)}.$
Let $\Theta_{n}(\bar{a})=$
$\prod_{\begin{subarray}{c}R\in\sigma\setminus\sigma_{\rho}\\ 1\leq i_{1}<\ldots<i_{\nu_{R}}\leq m\\ p(\bar{x})\models R(x_{i_{1}},...,x_{i_{\nu_{R}}})\end{subarray}}\mathcal{A}^{\prime}\big{(}\chi_{R}(a_{i_{1}},\ldots,a_{i_{\nu_{R}}})\big{)}\prod_{\begin{subarray}{c}R\in\sigma\setminus\sigma_{\rho}\\ 1\leq i_{1}<\ldots<i_{\nu_{R}}\leq m\\ p(\bar{x})\models\neg R(x_{i_{1}},...,x_{i_{\nu_{R}}})\end{subarray}}\Big{(}1-\mathcal{A}^{\prime}\big{(}\chi_{R}(a_{i_{1}},...,a_{i_{\nu_{R}}})\big{)}\Big{)}.$
Let $\varepsilon^{\prime}>0$. Since $\mathcal{A}^{\prime}\in\mathbf{X}_{n}^{\rho,\varepsilon}$ it follows that if $\varepsilon>0$ is chosen small enough (and the choice depends only on $\varepsilon^{\prime}$, $p$ and $p_{\rho}$), then
$\big{|}\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p(\bar{a})}\ \big{|}\ \mathbf{W}^{\mathcal{A}^{\prime}}\big{)}-\Theta_{n}(\bar{a})\big{|}\ \leq\ \varepsilon^{\prime}.$
Recall that, for every $R\in\sigma\setminus\sigma_{\rho}$, $\chi_{R}(x_{1},\ldots,x_{\nu_{R}})$ is a closure-basic formula over $\sigma_{\rho}$, so it has the form
$\bigwedge_{j=1}^{\eta_{R}}\big{(}\psi_{R,j}(x_{1},\ldots,x_{\nu_{R}})\rightarrow c_{R,j}\big{)}\ \ \text{ where $c_{R,j}\in[0,1]$}$
and $\psi_{R,j}$ is a closure type over $\sigma_{\rho}$. Since $p_{\rho}(x_{1},\ldots,x_{m})$ is a complete closure type over $\sigma_{\rho}$ it follows that for all $R\in\sigma\setminus\sigma_{\rho}$ and all $1\leq i_{1}<\ldots<i_{\nu_{R}}\leq m$, either
$p_{\rho}(x_{1},\ldots,x_{m})\models\psi_{R,j}(x_{i_{1}},\ldots,x_{i_{\nu_{R}}})\ \text{ or }\ p_{\rho}(x_{1},\ldots,x_{m})\models\neg\psi_{R,j}(x_{i_{1}},\ldots,x_{i_{\nu_{R}}}).$
So for all $R\in\sigma\setminus\sigma_{\rho}$ and all $1\leq i_{1}<\ldots<i_{\nu_{R}}\leq m$, $\mathcal{A}\big{(}\chi_{R}(a_{i_{1}},\ldots,a_{i_{\nu_{R}}})\big{)}$ is determined only by $p_{\rho}$. Hence $\Theta_{n}(\bar{a})$ is determined only by $p_{\rho}$ and $p$ (and it is a product of numbers of the form $c_{R,j}$ and $(1-c_{R,j})$). Let $\alpha=\Theta_{n}(\bar{a})$. Then, if $\varepsilon>0$ is small enough,
$\big{|}\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p(\bar{a})}\ \big{|}\ \mathbf{W}^{\mathcal{A}^{\prime}}\big{)}-\alpha\big{|}\ \leq\ \varepsilon^{\prime}$
This completes the proof. $\square$
Proposition 7.12.
Suppose that $p(\bar{x})$ is a complete self-contained closure type over $\sigma$ and that $p_{\tau}(\bar{x})$ is a closure type over $\tau$. Then $(p,p_{\tau})$ converges with respect to $\mathbb{G}$.
Proof. As said in Remark 7.3 we may assume that there are infinitely many $n$ such that $p(\bar{x})\wedge p_{\tau}(\bar{x})$ is satisfiable in some $\mathcal{A}\in\mathbf{W}_{n}$, and it follows that $p\negthickspace\upharpoonright\negthickspace\tau$ is equivalent to $p_{\tau}$ (as every closure type over $\tau$ is a complete closure type over $\tau$, by the definition). Let $p_{\rho}=p\negthickspace\upharpoonright\negthickspace\sigma_{\rho}$, so $p_{\rho}$ is a complete closure type over $\sigma_{\rho}$. By Assumption 7.8 (2) there is $\alpha$ such that $(p_{\rho},p_{\tau})$ converges to $\alpha$ with respect to $\mathbb{G}$. By Lemma 7.11 there is $\beta$ such that $(p,p_{\rho})$ converges to $\beta$ with respect to $\mathbb{G}$. Hence, for all $\varepsilon>0$ there is $n_{0}$ such that for all $n\geq n_{0}$ and all $\bar{a}\in(T_{n})^{|\bar{x}|}$,
$\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p_{\rho}(\bar{a})}\ \big{|}\ \mathbf{E}_{n}^{p_{\tau}(\bar{a})}\big{)}\ \in\ [\alpha-\varepsilon,\alpha+\varepsilon]\quad\text{ if $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p_{\tau}(\bar{a})}\big{)}>0$ and}$
$\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p(\bar{a})}\ \big{|}\ \mathbf{E}_{n}^{p_{\rho}(\bar{a})}\big{)}\ \in\ [\beta-\varepsilon,\beta+\varepsilon]\quad\text{ if $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p_{\rho}(\bar{a})}\big{)}>0$}.$
Now we get
$\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p(\bar{a})}\ \big{|}\ \mathbf{E}_{n}^{p_{\tau}(\bar{a})}\big{)}=\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p(\bar{a})}\ \big{|}\ \mathbf{E}_{n}^{p_{\rho}(\bar{a})}\big{)}\cdot\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p_{\rho}(\bar{a})}\ \big{|}\ \mathbf{E}_{n}^{p_{\tau}(\bar{a})}\big{)}\ \in\ [\alpha\beta-3\varepsilon,\alpha\beta+3\varepsilon].$
Since $\varepsilon>0$ can be chosen as small as we like this completes the proof. $\square$
Corollary 7.13.
Suppose that $p(\bar{x})$ is a (not necessarily complete) self-contained closure type over $\sigma$ and that $p_{\tau}(\bar{x})$ is a closure type over $\tau$. Then $(p,p_{\tau})$ converges with respect to $\mathbb{G}$.
Proof. Let $p(\bar{x})$ and $p_{\tau}(\bar{x})$ be as assumed. Then there are complete self-contained closure types over $\sigma$ $p_{1}(\bar{x}),\ldots,p_{k}(\bar{x})$ such that $p(\bar{x})$ is equivalent to $\bigvee_{i=1}^{k}p_{i}(\bar{x})$ and if $i\neq j$ then $p_{i}(\bar{x})\wedge p_{j}(\bar{x})$ is inconsistent. Then for every $n$ and $\bar{a}\in(T_{n})^{|\bar{x}|}$ we have
$\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p(\bar{a})}\ \big{|}\ \mathbf{E}_{n}^{p_{\tau}(\bar{a})}\big{)}=\sum_{i=1}^{k}\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p_{i}(\bar{a})}\ \big{|}\ \mathbf{E}_{n}^{p_{\tau}(\bar{a})}\big{)}$
where, for each $i$, $(p_{i},p_{\tau})$ converges, by Proposition 7.12, to some $\alpha_{i}$. It follows that $(p,p_{\tau})$ converges to $\alpha=\alpha_{1}+\ldots+\alpha_{k}$. $\square$
Remark 7.14.
By Corollary 7.13, part (2) of Assumption 7.8 holds if $\sigma_{\rho}$ and $\mathbb{G}_{\rho}$ are replaced by $\sigma$ and $\mathbb{G}$, respectively. So the induction step for convergence is completed, that is, claim (B) above is proved.
Remark 7.15.
Let $p(\bar{x})$, $p_{\rho}(\bar{x})$, and $p_{\tau}(\bar{x})$ be complete closure types over $\sigma$, $\sigma_{\rho}$, and $\tau$, respectively (and recall that a closure type over $\tau$ is the same as a complete closure type over $\tau$). Suppose that $p\models p_{\rho}$ and $p_{\rho}\models p_{\tau}$. Suppose that, for every $R\in\sigma\setminus\sigma_{\rho}$, $\theta_{R}$ is a closure-basic formula over $\sigma_{\rho}$. In the proof of Lemma 7.11 we can then let $\chi_{R}$ be the same formula as $\theta_{R}$. Then, with the notation of that proof, we get $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p(\bar{a})}\ \big{|}\ \mathbf{E}_{n}^{p_{\rho}(\bar{a})}\big{)}=\Theta_{n}(\bar{a})=\alpha$ for all $n$ and $\bar{a}\in(T_{n})^{|\bar{x}|}$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p_{\rho}(\bar{a})}\big{)}>0$. Hence $(p,p_{\rho})$ is eventually constant with value $\alpha$.
Suppose, in addition, that $(p_{\rho},p_{\tau})$ is eventually constant with value $\beta$. It follows straightforwardly from the definition of ‘eventually constant’ that $(p,p_{\tau})$ is eventually constant with value $\alpha\beta$. By induction on the height of the underlying DAG of $\mathbb{G}$, it follows that if, for all $R\in\sigma\setminus\tau$, $\theta_{R}$ is a closure-basic formula, then $(p,p_{\tau})$ is eventually constant, for all closure types $p_{\tau}(\bar{x})$ over $\tau$ and all complete closure types $p(\bar{x})$ over $\sigma$. By the proof of Corollary 7.13 it follows that (under the same assumption) $(p,p_{\tau})$ is eventually constant also in the case when $p$ is not complete.
7.4. Balance
In this section we prove statement (C), in the first part of Lemma 7.17, but we also prove other results that will be used to prove statement (A).
Lemma 7.16.
Suppose that $p_{\rho}(\bar{x},y)$ is a complete self-contained closure type over $\sigma_{\rho}$ and that $p(\bar{x},y)$ is a self-contained closure type over $\sigma$ such that $p\models p_{\rho}$ (or equivalently, $p_{\rho}$ is equivalent to $p\negthickspace\upharpoonright\negthickspace\sigma_{\rho}$). Suppose that $\mathrm{rank}_{y}(p_{\rho})\ (=\mathrm{rank}_{y}(p))=1$ and that for every $R\in\sigma\setminus\sigma_{\rho}$, if $\bar{z}$ is a subsequence of $\bar{x}\bar{y}$ and $p\models R(\bar{z})$ or $p\models\neg R(\bar{z})$, then $\bar{z}$ contains $y$. Furthermore, suppose that, for some $n$, $\bar{a}\in(T_{n})^{|\bar{x}|}$, $\mathcal{A}^{\prime}\in\mathbf{W}_{n}^{\rho}$, $P\subseteq p_{\rho}(\bar{a},\mathcal{A}^{\prime})$, and $\mathbb{P}_{n}^{\rho}(\mathcal{A}^{\prime})>0$. There is $\alpha\in[0,1]$, depending only on $p_{\rho}$ and $p$, such that for every $\varepsilon>0$ there is $c>0$, depending only on $\alpha$ and $\varepsilon$, such that if $n$ and $|P|$ are large enough, then
[…]
Proof. We adopt all assumptions of the lemma. Since $\mathbb{P}_{n}^{\rho}(\mathcal{A}^{\prime})>0$ it follows from Lemma 6.10 that $\mathbb{P}_{n}\big{(}\mathbf{W}^{\mathcal{A}^{\prime}}\big{)}>0$. As $\mathcal{A}^{\prime}\models p_{\rho}(\bar{a},b)$ for all $b\in P$, it follows that for all $b\in P$, $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p_{\rho}(\bar{a},b)}\big{)}>0$. Lemma 7.11 tells that, for some $\alpha$, $(p,p_{\rho})$ converges to $\alpha$ with respect to $\mathbb{G}$. So for every $\varepsilon>0$, if $n$ is large enough we have
$\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p(\bar{a},b)}\ \big{|}\ \mathbf{E}_{n}^{p_{\rho}(\bar{a},b)}\big{)}\ \in\ [\alpha-\varepsilon,\alpha+\varepsilon]\quad\text{ for all }b\in P.$
Since, by assumption, $\mathcal{A}^{\prime}\models p_{\rho}(\bar{a},b)$ for all $b\in P$ we have $\mathbf{W}^{\mathcal{A}^{\prime}}\subseteq\mathbf{E}_{n}^{p_{\rho}(\bar{a},b)}$ for all $b\in P$. Hence
$\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p(\bar{a},b)}\ \big{|}\ \mathbf{W}^{\mathcal{A}^{\prime}}\big{)}=\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p(\bar{a},b)}\ \big{|}\ \mathbf{E}_{n}^{p_{\rho}(\bar{a},b)}\big{)}\ \in\ [\alpha-\varepsilon,\alpha+\varepsilon]\quad\text{ for all }b\in P.$
By the assumption that $p_{\rho}$ (and hence $p$) is self-contained, the assumption regarding the literals using a relation symbol from $\sigma\setminus\sigma_{\rho}$, and Lemma 6.10, if we condition on $\mathbf{W}^{\mathcal{A}^{\prime}}$, then for every $b\in P$ the event $\mathbf{E}_{n}^{p(\bar{a},b)}$ is independent from the events $\mathbf{E}_{n}^{p(\bar{a},b^{\prime})}$ as $b^{\prime}$ ranges over $P\setminus\{b\}$. Corollary 2.2 now implies that there is $c>0$, depending only on $\alpha$ and $\varepsilon$, such that if $n$ and $|P|$ are large enough then
[…]
so the proof is complete. $\square$
Lemma 7.17.
Let $p_{\tau}(\bar{x},y)$ be a self-contained closure type over $\tau$ and let $p(\bar{x},y)$ and $q(\bar{x})$ be complete closure types over $\sigma$. Suppose that $\mathrm{rank}_{y}(p_{\tau})=1$.
(i) Also suppose that $p\wedge p_{\tau}\wedge q$ is cofinally satisfiable. Then there is $\gamma\in[0,1]$ such that for all $\varepsilon>0$ there is $c>0$ such that for all sufficiently large $n$, if $\bar{a}\in(T_{n})^{|\bar{x}|}$, $B\subseteq p_{\tau}(\bar{a},\mathcal{T}_{n})$, and $|B|\geq g_{3}(n)$, then
$\mathbb{P}_{n}\big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\text{ if $\mathcal{A}\models q(\bar{a})$ then }(\gamma-\varepsilon)|B|\leq|p(\bar{a},\mathcal{A})\cap B|\leq(\gamma+\varepsilon)|B|\big{\}}\big{)}\ \geq$
$1-e^{-cg_{3}(n)}.$
(ii) $(p,p_{\tau},q)$ is balanced with respect to $\mathbb{G}$.
Proof. (i) The assumption that $p\wedge p_{\tau}\wedge q$ is cofinally satisfiable implies that $p\wedge p_{\tau}\wedge q$ is consistent and this implies that $q(\bar{x})$ is equivalent to $p\negthickspace\upharpoonright\negthickspace\bar{x}$ and $p_{\tau}(\bar{x},\bar{y})$ is equivalent to $p\negthickspace\upharpoonright\negthickspace\tau$ (so $\mathrm{rank}_{y}(p)=\mathrm{rank}_{y}(p_{\tau})=1$). Let $p_{\rho}(\bar{x},y)=p\negthickspace\upharpoonright\negthickspace\sigma_{\rho}$ and $q_{\rho}(\bar{x})=q\negthickspace\upharpoonright\negthickspace\sigma_{\rho}$, so $p_{\rho}(\bar{x},y)$ and $q_{\rho}(\bar{x})$ are complete closure types over $\sigma_{\rho}$. By the induction hypothesis, Assumption 7.8, there is $\alpha\in[0,1]$ such that for all $\varepsilon>0$ there is $d>0$ such that for all sufficiently large $n$, if $\bar{a}\in(T_{n})^{|\bar{x}|}$, $B\subseteq p_{\tau}(\bar{a},\mathcal{T}_{n})$, $|B|\geq g_{3}(n)$, and
(7.2) $\mathbf{X}_{n}^{\rho,\varepsilon}=$
$\big{\{}\mathcal{A}^{\prime}\in\mathbf{W}_{n}^{\rho}:\text{ if $\mathcal{A}^{\prime}\models q_{\rho}(\bar{a})$ then }(\alpha-\varepsilon)|B|\leq|p_{\rho}(\bar{a},\mathcal{A}^{\prime})\cap B|\leq(\alpha+\varepsilon)|B|\big{\}},$
then $\text{then }\mathbb{P}_{n}^{\rho}\big{(}\mathbf{X}_{n}^{\rho,\varepsilon}\big{)}\ \geq\ 1-e^{-dg_{3}(n)}.$
First suppose that $\alpha=0$. Let $\mathbf{X}_{n}^{\varepsilon}=\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\mathcal{A}\negthickspace\upharpoonright\negthickspace\sigma_{\rho}\in\mathbf{X}_{n}^{\rho,\varepsilon}\big{\}}$. By Lemma 6.10, $\mathbb{P}_{n}\big{(}\mathbf{X}_{n}^{\varepsilon}\big{)}=\mathbb{P}_{n}^{\rho}\big{(}\mathbf{X}_{n}^{\rho,\varepsilon}\big{)}$ so $\lim_{n\to\infty}\mathbb{P}_{n}\big{(}\mathbf{X}_{n}^{\varepsilon}\big{)}=1$. Since $p$ implies $p_{\rho}$ and $q$ implies $q_{\rho}$ it follows that if $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon}$ and $\mathcal{A}\models q(\bar{a})$, then $|p(\bar{a},\mathcal{A})|\leq\varepsilon|B|$. Hence the conclusion of part (i) of the lemma holds with $\gamma=0$.
Now suppose that $\alpha>0$. Let $\varepsilon>0$, $\bar{a}\in(T_{n})^{|\bar{x}|}$, $B\subseteq p_{\tau}(\bar{a},\mathcal{T}_{n})$, and $|B|\geq g_{3}(n)$. Without loss of generality we may assume that $\alpha>\varepsilon$. Suppose that $\mathcal{A}^{\prime}\in\mathbf{X}_{n}^{\rho,\varepsilon}$, $\mathcal{A}^{\prime}\models q_{\rho}(\bar{a})$, and $\mathbb{P}_{n}^{\rho}(\mathcal{A}^{\prime})>0$, hence $\mathbb{P}_{n}\big{(}\mathbf{W}^{\mathcal{A}^{\prime}}\big{)}>0$ (by Lemma 6.10). By (7.2)
(7.3) $(\alpha-\varepsilon)|B|\leq|p_{\rho}(\bar{a},\mathcal{A}^{\prime})\cap B|\leq(\alpha+\varepsilon)|B|.$
Let $\hat{p}(\bar{x},y)$ be the conjuction of all $(\sigma\setminus\sigma_{\rho})$-literals $\varphi(\bar{z})$ such that $p(\bar{x},y)\models\varphi(\bar{z})$ and $\bar{z}$ is a subsequence of $\bar{x}y$ such that $y$ occurs in $\bar{z}$.
By Lemma 7.16 (with $P=p_{\rho}(\bar{a},\mathcal{A}^{\prime})\cap B$), there is $\beta$, depending only on $p_{\rho}$ and $\hat{p}$, and $d^{\prime}>0$, depending only on $\beta$ and $\varepsilon$, such that if $n$ is large enough, then
(7.4) $\mathbb{P}_{n}\Big{(}\Big{\{}\mathcal{A}\in\mathbf{W}^{\mathcal{A}^{\prime}}:$ $(\beta-\varepsilon)|p_{\rho}(\bar{a},\mathcal{A}^{\prime})\cap B|\ \leq\ |p_{\rho}(\bar{a},\mathcal{A}^{\prime})\cap\hat{p}(\bar{a},\mathcal{A})\cap B|\ \leq$
$(\beta+\varepsilon)|p_{\rho}(\bar{a},\mathcal{A}^{\prime})\cap B|\Big{\}}\ \Big{|}\ \mathbf{W}^{\mathcal{A}^{\prime}}\Big{)}\ \geq$
[…]
From (7.4) and (7.3) we get
(7.5) $\mathbb{P}_{n}\Big{(}\Big{\{}\mathcal{A}\in\mathbf{W}^{\mathcal{A}^{\prime}}:$ $(\alpha\beta-3\varepsilon)|B|\ \leq\ |p_{\rho}(\bar{a},\mathcal{A}^{\prime})\cap\hat{p}(\bar{a},\mathcal{A})\cap B|\ \leq$
$(\alpha\beta+3\varepsilon)|B|\Big{\}}\ \Big{|}\ \mathbf{W}^{\mathcal{A}^{\prime}}\Big{)}\ \geq$
$1-e^{-d^{\prime}(\alpha-\varepsilon)g_{3}(n)}=1-e^{-d^{\prime\prime}g_{3}(n)}\text{ if }d^{\prime\prime}=d^{\prime}(\alpha-\varepsilon).$
Since $p_{\rho}\in PLA^{*}(\sigma_{\rho})$ we have $p_{\rho}(\bar{a},\mathcal{A})=p_{\rho}(\bar{a},\mathcal{A}^{\prime})$ for all $\mathcal{A}\in\mathbf{W}^{\mathcal{A}^{\prime}}$. Also, $q(\bar{x})\wedge p_{\rho}(\bar{x},\bar{y})\wedge\hat{p}(\bar{x},\bar{y})\models p(\bar{x},\bar{y})$. So (7.5) implies that
(7.6) $\mathbb{P}_{n}\Big{(}\Big{\{}\mathcal{A}\in\mathbf{W}_{n}:$ if $\mathcal{A}\models q(\bar{a})$ then
$(\alpha\beta-3\varepsilon)|B|\leq|p(\bar{a},\mathcal{A})\cap B|\leq(\alpha\beta+3\varepsilon)|B|\Big{\}}\ \Big{|}\ \mathbf{W}^{\mathcal{A}^{\prime}}\Big{)}$
$\geq\ 1-e^{-d^{\prime\prime}g_{3}(n)}.$
Define $\mathbf{X}_{n}^{\varepsilon}=\bigcup_{\mathcal{A}^{\prime}\in\mathbf{X}_{n}^{\rho,\varepsilon}}\mathbf{W}^{\mathcal{A}^{\prime}}$ and note that the union is disjoint. From (7.6) and Lemma 2.3 we now get
(7.7) $\mathbb{P}_{n}\Big{(}\Big{\{}\mathcal{A}\in\mathbf{W}_{n}:$ if $\mathcal{A}\models q(\bar{a})$ then
$(\alpha\beta-3\varepsilon)|B|\leq|p(\bar{a},\mathcal{A})\cap B|\leq(\alpha\beta+3\varepsilon)|B|\Big{\}}\ \Big{|}\ \mathbf{X}_{n}^{\varepsilon}\Big{)}$
$\geq\ 1-e^{-d^{\prime\prime}g_{3}(n)}.$
Let
$\mathbf{Y}_{n}^{\varepsilon}=\big{\{}\mathcal{A}\in\mathbf{W}_{n}:$ if $\mathcal{A}\models q(\bar{a})$ then
$(\alpha\beta-3\varepsilon)|B|\leq|p(\bar{a},\mathcal{A})\cap B|\leq(\alpha\beta+3\varepsilon)|B|\big{\}},$
so $\mathbb{P}_{n}\big{(}\mathbf{Y}_{n}^{\varepsilon}\big{|}\mathbf{X}_{n}^{\varepsilon}\big{)}\geq 1-e^{-d^{\prime\prime}g_{3}(n)}$. By Lemma 6.10 we have $\mathbb{P}_{n}\big{(}\mathbf{W}^{\mathcal{A}^{\prime}}\big{)}=\mathbb{P}_{n}^{\rho}(\mathcal{A}^{\prime})$ for all $\mathcal{A}^{\prime}\in\mathbf{W}_{n}^{\rho}$ and therefore $\mathbb{P}_{n}\big{(}\mathbf{X}_{n}^{\varepsilon}\big{)}=\mathbb{P}_{n}^{\rho}\big{(}\mathbf{X}_{n}^{\rho,\varepsilon}\big{)}$. By (7.2) we get $\mathbb{P}_{n}\big{(}\mathbf{X}_{n}^{\varepsilon}\big{)}\geq 1-e^{-dg_{3}(n)}$ for all sufficiently large $n$. Hence, for all large enough $n$,
$\mathbb{P}_{n}\big{(}\mathbf{Y}_{n}^{\varepsilon}\big{)}\geq\mathbb{P}_{n}\big{(}\mathbf{Y}_{n}^{\varepsilon}\ \big{|}\ \mathbf{X}_{n}^{\varepsilon}\big{)}\cdot\mathbb{P}_{n}\big{(}\mathbf{X}_{n}^{\varepsilon})\ \geq\ \big{(}1-e^{-d^{\prime\prime}g_{3}(n)}\big{)}\big{(}1-e^{-dg_{3}(n)}\big{)}\ \geq$
$1-e^{-cg_{3}(n)}\quad\text{ for an appropriate choice of $c>0$ which depends only on $d$ and $d^{\prime\prime}$.}$
Now the conclusion of part (i) follows if $\gamma=\alpha\beta$ because $\varepsilon>0$ can be chosen as small as we like.
(ii) By Remark 7.5 we may assume that $p\wedge p_{\tau}\wedge q$ is cofinally satisfiable. It follows that that $q(\bar{x})$ is equivalent to $p\negthickspace\upharpoonright\negthickspace\bar{x}$ and $p_{\tau}(\bar{x},y)$ is equivalent to $p\negthickspace\upharpoonright\negthickspace\tau$. Let $\bar{x}=(x_{1},\ldots,x_{k})$. Since $\mathrm{rank}_{y}(p_{\tau})=1$ it follows that for some $i\in\{1,\ldots,k\}$
$p_{\tau}(\bar{x},y)\models_{tree}\text{ ``$y$ is a child of $x_{i}$ and $y\neq x_{j}$ for all $j=1,\ldots,k$''.}$ ”.
For notational simplicity (and without loss of generality) let us assume that $i=1$.
Suppose that (for some $n$) $\bar{a}=(a_{1},\ldots,a_{k})\in(T_{n})^{k}$, $p_{\tau}(\bar{a},\mathcal{A})\neq\emptyset$ (so $\bar{a}$ satisfies the restriction of $p_{\tau}$ to $\bar{x}$). Then $p_{\tau}(\bar{a},\mathcal{A})$ is the set of all children of $a_{1}$ that do not belong to $\mathrm{rng}(\bar{a})$. By Assumption 6.3, $|p_{\tau}(\bar{a},\mathcal{A})|\geq g_{1}(n)-|\bar{x}|$. By the same assumption again, we have $\lim_{n\to\infty}(g_{1}(n)-g_{3}(n))=\infty$, so if $n$ is large enough then $|p_{\tau}(\bar{a},\mathcal{A})|\geq g_{3}(n)$.
It now follows from part (i), with $B=p_{\tau}(\bar{a},\mathcal{A})$, that there is $\gamma$, depending only on $p_{\tau}$ and $p$, such that for every $\varepsilon>0$ there is $c>0$, depending only on $\varepsilon$ and $\gamma$, such that if $n$ is sufficiently large
$\mathbb{P}_{n}\big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:$ if $\mathcal{A}\models q(\bar{a})$ then
$(\gamma-\varepsilon)|p_{\tau}(\bar{a},\mathcal{A})|\leq|p(\bar{a},\mathcal{A})|\leq(\gamma+\varepsilon)|p_{\tau}(\bar{a},\mathcal{A})|\big{\}}\big{)}\ \geq$
$1-e^{-cg_{3}(n)}.$
The above holds for all $\bar{a}\in(T_{n})^{|\bar{x}|}$. (If $p_{\tau}(\bar{a},\mathcal{A})=\emptyset$ then the inequalities that define the event above are trivial since all cardinalities involved are 0 in this case.) The number of $\bar{a}\in(T_{n})^{|\bar{x}|}$ is (by Assumption 6.3) bounded from above by $(g_{4}(n)^{\Delta})^{|\bar{x}|}$ which is bounded by $f(n)$ for some polynomial $f$. It follows that
$\mathbb{P}_{n}\big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:(p,p_{\tau},q)\text{ is {\bf not} $(\gamma,\varepsilon)$-balanced in $\mathcal{A}$}\big{\}}\big{)}\ \leq\ f(n)e^{-cg_{3}(n)}.$
By Lemma 6.5 , $\lim_{n\to\infty}f(n)e^{-cg_{3}(n)}=0$. As $\varepsilon>0$ was arbitrary it follows that $(p,p_{\tau},q)$ is $\gamma$-balanced with respect to $\mathbb{G}$. $\square$
Remark 7.18.
By part (i) of Lemma 7.17 we have proved statement (C) above (at the end of Section 7.2).
Although we have completed the inductive step for part (3) of Assumption 7.8 we will continue to prove results about balanced triples because we need them for proving statement (A) and the main results of this article.
Proposition 7.19.
Let $p_{\tau}(\bar{x},\bar{y})$ be a self-contained and $\bar{y}$-independent closure type over $\tau$ and let $p(\bar{x},\bar{y})$ and $q(\bar{x})$ be complete closure types over $\sigma$. Then $(p,p_{\tau},q)$ is balanced with respect to $\mathbb{G}$.
Proof. By Remark 7.5 we may assume that $p\wedge p_{\tau}\wedge q$ is cofinally satisfiable, hence consistent. Since $p$ and $q$ are complete closure types over $\sigma$, it follows that $q(\bar{x})$ is equivalent to $p\negthickspace\upharpoonright\negthickspace\bar{x}$ and $p_{\tau}(\bar{x},\bar{y})$ is equivalent to $p\negthickspace\upharpoonright\negthickspace\tau$. We use induction on $\mathrm{rank}_{\bar{y}}(p_{\tau})$ (which equals $\mathrm{rank}_{\bar{y}}(p)$). If $\mathrm{rank}_{\bar{y}}(p_{\tau})=0$ then $p$ is not $\bar{y}$-independent, contradicting the assumption, so $\mathrm{rank}_{\bar{y}}(p_{\tau})\geq 1$. If $\mathrm{rank}_{\bar{y}}(p_{\tau})=1$ then $(p,p_{\tau},q)$ is balanced with respect to $\mathbb{G}$, by Lemma 7.17.
So suppose that $\mathrm{rank}_{\bar{y}}(p_{\tau})=\kappa+1$ where $\kappa\geq 1$. The induction hypothesis is that the proposition holds if we add the assumption that $\mathrm{rank}_{\bar{y}}(p_{\tau})\leq\kappa$. By Remark 5.20, we may assume that $\bar{y}=z\bar{w}$, $\mathrm{rank}_{\bar{w}}(p_{\tau})=\kappa$, $\mathrm{rank}_{z}(r_{\tau})=1$, where $r(\bar{x},z)=p\negthickspace\upharpoonright\negthickspace\bar{x}z$ and $r_{\tau}(\bar{x},z)=p_{\tau}\negthickspace\upharpoonright\negthickspace\bar{x}z$. By the same remark we may also assume that $r_{\tau}$ is self-contained. As $\mathrm{rank}_{z}(r_{\tau})=1$ it follows that $r$ is $z$-independent. Since $\bar{w}$ is a subsequence of $\bar{y}$ and $p_{\tau}$ is $\bar{y}$-independent it follows thay $p_{\tau}$ is also $\bar{w}$-independent. Note that $r(\bar{x},z)$ is a complete closure type over $\sigma$. Since $\mathrm{rank}_{z}(r_{\tau})=1$ it follows that
$p(\bar{x},z,\bar{w})\models_{tree}\text{ ``$z$ is a child of a member of $\bar{x}$''.}$ ”.
By reordering $\bar{w}$ if necessary we may also assume that $\bar{w}=\bar{u}\bar{v}$ and
$p(\bar{x},z,\bar{u},\bar{v})\models_{tree}$ “all members of $\bar{u}$ are successors of $z$, and
no member of $\bar{v}$ is a successor of $z$”.
It follows that
$p(\bar{x},z,\bar{u},\bar{v})\models_{tree}\text{ ``every member of $\bar{v}$ is a successor of some member of $\bar{x}$''.}$ ”.
Let $s(\bar{x},z,\bar{u})=p\negthickspace\upharpoonright\negthickspace\bar{x}z\bar{u}$ and $t(\bar{x},\bar{v})=p\negthickspace\upharpoonright\negthickspace\bar{x}\bar{v}$ and note that $s$ and $t$ are complete closure types over $\sigma$. Let $s_{\tau}=s\negthickspace\upharpoonright\negthickspace\tau$ and $t_{\tau}=t\negthickspace\upharpoonright\negthickspace\tau$. Then $p_{\tau}\negthickspace\upharpoonright\negthickspace\bar{x}z\bar{u}$ is equivalent to $s_{\tau}(\bar{x},z,\bar{u})$ and $p_{\tau}\negthickspace\upharpoonright\negthickspace\bar{x}\bar{v}$ is equivalent to $t_{\tau}(\bar{x},\bar{v})$. By the choices of $z$, $\bar{u}$, and $\bar{v}$, $s(\bar{x},z,\bar{u})$ is self-contained and $z\bar{u}$-independent, and $t(\bar{x},\bar{v})$ is self-contained and $\bar{v}$-independent. Moreover, in every $\sigma$-structure $\mathcal{A}$ that expands a tree,
$p_{\tau}(\bar{x},z,\bar{u},\bar{v})$ is equivalent to $s_{\tau}(\bar{x},z,\bar{u})\wedge t_{\tau}(\bar{x},\bar{v})$.
It follows that
(7.8) for all n, all a¯(Tn)|x¯| and all $\text{for all $n$, all $\bar{a}\in(T_{n})^{|\bar{x}|}$ and all $\mathcal{A}\in\mathbf{W}_{n}$},|p_{\tau}(\bar{a},\mathcal{A})|=|s_{\tau}(\bar{a},\mathcal{A})|\cdot|t_{\tau}(\bar{a},\mathcal{A})|.$
Case 1. Suppose that $\bar{v}$ is not empty.
Then $1\leq\mathrm{rank}_{\bar{v}}(p)\leq\kappa$ and $1\leq\mathrm{rank}_{z\bar{u}}(s)\leq\kappa$. The induction hypothesis implies that $(s,s_{\tau},q)$ is $\alpha$-balanced and $(p,p_{\tau},s)$ is $\beta$-balanced (with respect to $\mathbb{G}$) for some $\alpha$ and $\beta$.
Let $\varepsilon>0$ and define
$\mathbf{X}_{n}^{\varepsilon}=\big{\{}\mathcal{A}\in\mathbf{W}_{n}:(s,s_{\tau},q)\text{ is $(\alpha,\varepsilon)$-balanced in $\mathcal{A}$}\big{\}}\ \text{ and}$
$\mathbf{Y}_{n}^{\varepsilon}=\big{\{}\mathcal{A}\in\mathbf{W}_{n}:(p,p_{\tau},s)\text{ is $(\beta,\varepsilon)$-balanced in $\mathcal{A}$}\big{\}}.$
Then $\lim_{n\to\infty}\mathbb{P}_{n}\big{(}\mathbf{X}_{n}^{\varepsilon}\cap\mathbf{Y}_{n}^{\varepsilon}\big{)}=1$.
Let $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon}\cap\mathbf{Y}_{n}^{\varepsilon}$ and $\bar{a}\in(T_{n})^{|\bar{x}|}$. Then
$|p(\bar{a},\mathcal{A})|\ =\ \sum_{b\bar{c}\in s(\bar{a},\mathcal{A})}|p(\bar{a},b,\bar{c},\mathcal{A})|\ \leq\ (\beta+\varepsilon)\sum_{b\bar{c}\in s(\bar{a},\mathcal{A})}|p_{\tau}(\bar{a},b,\bar{c},\mathcal{A})|$
[…]
$=\ (\beta+\varepsilon)|s(\bar{a},\mathcal{A})|\cdot|t_{\tau}(\bar{a},\mathcal{A})|\ \leq\ (\beta+\varepsilon)(\alpha+\varepsilon)|s_{\tau}(\bar{a},\mathcal{A})|\cdot|t_{\tau}(\bar{a},\mathcal{A})|$
[…]
$\leq\ (\alpha\beta+3\varepsilon)|p_{\tau}(\bar{a},\mathcal{A})|.$
In a similar way we can derive that $(\alpha\beta-3\varepsilon)|p_{\tau}(\bar{a},\mathcal{A})|\leq|p(\bar{a},\mathcal{A})|$. Hence $(p,p_{\tau},q)$ is $(\alpha\beta,3\varepsilon)$-balanced in every $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon}\cap\mathbf{Y}_{n}^{\varepsilon}$. As $\varepsilon>0$ can be taken as small as we like it follows that $(p,p_{\tau},q)$ is $\alpha\beta$-balanced with respect to $\mathbb{G}$.
Case 2. Suppose that $\bar{v}$ is empty.
Then $\bar{y}=z\bar{u}$ so $s(\bar{x},z,\bar{u})$ is the same as $p(\bar{x},z,\bar{u})$ (and $p_{\tau}$ is the same as $s_{\tau}$). Since $p_{\tau}(\bar{x},z,\bar{u})$ is self-contained it follows that if $\mathcal{T}_{n}\models p_{\tau}(\bar{a},b,\bar{c})$ then $\mathcal{T}^{\prime}:=\mathcal{T}_{n}\negthickspace\upharpoonright\negthickspace\mathrm{rng}(b\bar{c})$ is a subtree of $\mathcal{T}_{n}$ which is rooted in $b$ and the isomorphism type of $\mathcal{T}^{\prime}$ is determined by $p_{\tau}$ alone. It is easy to see that if $\mathcal{T}_{n}\models r_{\tau}(\bar{a},b)$ and $\mathcal{T}^{\prime\prime}$ is a subtree of $\mathcal{T}_{n}$ which is rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$, then the elements of $T^{\prime\prime}\setminus\{b\}$ can be ordered in a unique way as $\bar{c}$ so that $\mathcal{T}_{n}\models p_{\tau}(\bar{a},b,\bar{c})$. Therefore there is a bijection between $p_{\tau}(\bar{a},b,\mathcal{T}_{n})$ and the subtrees of $\mathcal{T}_{n}$ that are rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$. Recall that $\mathsf{N}_{n}(b,\mathcal{T}^{\prime}):=\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})$ is the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$. Now we have $|p_{\tau}(\bar{a},b,\mathcal{T}_{n})|=\mathsf{N}_{n}(b,\mathcal{T}^{\prime})$ if $p_{\tau}(\bar{a},b,\mathcal{T}_{n})\neq\emptyset$. Recall that, by Remark 6.4, we have $\mathsf{N}_{n}(b,\mathcal{T}^{\prime})\geq g_{1}(n)$ whenever $n$ is large enough and $\mathsf{N}_{n}(b,\mathcal{T}^{\prime})>0$.
Suppose that $\bar{x}=(x_{1},\ldots,x_{k})$. Without loss of generality we can assume that $p_{\tau}(\bar{x},z,\bar{u})\models$$z$ is a child of $x_{1}$”. It follows that for all $n$ and all $\bar{a}=(a_{1},\ldots,a_{k})\in(T_{n})^{k}$, if $r_{\tau}(\bar{a},\mathcal{T}_{n})\neq\emptyset$ then $r_{\tau}(\bar{a},\mathcal{T}_{n})$ is the set of all children of $a_{1}$ that do not belong to $\bar{a}$.
We now assume that $p_{\tau}(\bar{a},\mathcal{T}_{n})\neq\emptyset$; otherwise there is nothing to prove. Then $r_{\tau}(\bar{a},\mathcal{T}_{n})\neq\emptyset$ and for every $b\in r_{\tau}(\bar{a},\mathcal{A})$ we have $|p_{\tau}(\bar{a},b,\mathcal{T}_{n})|=\mathsf{N}_{n}(b,\mathcal{T}^{\prime})\geq g_{1}(n)$ if $n$ is large enough. We will now define a partition $P_{1},\ldots,P_{\mu}$ of $r_{\tau}(\bar{a},\mathcal{T}_{n})$. Recall the functions $g_{1},g_{2},g_{3}:\mathbb{N}\to\mathbb{R}^{+}$ from Assumption 6.3.
  1. (1)
    Let $\lambda\geq 1$ be the minimal integer such that there is $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $g_{1}(n)+(\lambda-1)g_{2}(n)\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<g_{1}(n)+\lambda g_{2}(n)$.
    1. (a)
      If $\lambda=1$ then let $P_{1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{1,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{1,2}$ where $m_{1,1}:=g_{1}(n)$ and $m_{1,2}:=m_{1,1}+(\lambda+3)g_{2}(n)$.
    2. (b)
      If $\lambda>1$ then let $P_{1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{1,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{1,2}$ where $m_{1,1}:=g_{1}(n)+(\lambda-1)g_{2}(n)$ and $m_{1,2}:=m_{1,1}+(\lambda+3)g_{2}(n)$.
    Assumption 6.3 implies that, in both cases, $|P_{1}|\geq g_{3}(n)$.
  2. (2)
    Now suppose that the parts $P_{1},\ldots,P_{l}$ have been defined such that $|P_{i}|\geq g_{3}(n)$ for all $i$ and that numbers $m_{i,j}$ (with $j\in\{1,2\}$) have been defined such that $P_{i}$ contains all $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{i,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{i,2}$, and $m_{i+1,1}=m_{i,2}$ if $i<l$.
    Let $\lambda\geq 1$ be the minimal integer (if it exists) such that there is $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{l,2}+\lambda g_{2}(n)\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{l,2}+(\lambda+1)g_{2}(n)$.
    1. (a)
      If $\lambda=1$ then let $P_{l+1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{l+1,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{l+1,2}$ where $m_{l+1,1}:=m_{l,2}$ and $m_{l+1,2}:=m_{l,2}+3g_{2}(n)$.
    2. (b)
      If $\lambda>1$ then
      1. (i)
        let $P_{l+1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{l+1,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{l+1,2}$ where $m_{l+1,1}:=m_{l,2}+(\lambda-1)g_{2}(n)$ and $m_{l+1,2}:=m_{l,2}+(\lambda+2)g_{2}(n)$, and
      2. (ii)
        redefine $P_{l}$ so that it contains contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{l,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{l,2}+g_{2}(n)$ and then redefine $m_{l,2}$ by $m_{l,2}:=m_{l,2}+g_{2}(n)$.
    In both cases we have $|P_{l+1}|\geq g_{3}(n)$ by Assumption 6.3. If no $\lambda$ as above exists then $r_{\tau}(\bar{a},\mathcal{T}_{n})=P_{1}\cup\ldots\cup P_{l}$ and we are done and let $\mu=l$.
Observe that when a part $P_{i}$ was first defined then, for some real number $m_{i}$, all $b\in P_{i}$ satisfy $m_{i}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{i}+3g_{2}(n)$. Also, each part $P_{i}$ was redefined at most once and in such a way that (with the same $m_{i}$) all $b\in P_{i}$ satisfy $m_{i}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{i}+4g_{2}(n)$. By the construction we have $m_{i}\geq g_{1}(n)$ for all $i$. It follows that
(7.9) for each part $P_{i}$, $|P_{i}|\geq g_{3}(n)$ and there is $m_{i}\geq g_{1}(n)$ such that
for all $b\in P_{i}$, $m_{i}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})\leq m_{i}+4g_{2}(n)$.
It follows from part (i) of Lemma 7.17 that there is $\alpha$, depending only on $r$, such that for every $\varepsilon>0$ there is $C>0$, depending only on $\alpha$ and $\varepsilon$, such that if $n$ is sufficiently large, then for every part $P_{i}\subseteq r_{\tau}(\bar{a},\mathcal{T}_{n})$,
[…]
$\mathbf{X}_{n}^{\varepsilon,i}(\bar{a})=\big{\{}\mathcal{A}\in\mathbf{W}_{n}:$ if $\mathcal{A}\models q(\bar{a})$ then
$(\alpha-\varepsilon)|P_{i}|\leq|r(\bar{a},\mathcal{A})\cap P_{i}|\leq(\alpha+\varepsilon)|P_{i}|\big{\}}.$
Recall that (by Assumption 6.3) every element of $\mathcal{T}_{n}$ has at most $f_{4}(n)$ children, where $f_{4}$ is a polynomial. Since all members of $r_{\tau}(\bar{a},\mathcal{T}_{n})$ are children of $a_{1}$ it follows that $|r_{\tau}(\bar{a},\mathcal{T}_{n})|\leq f_{4}(n)$. As $P_{1},\ldots,P_{\mu}$ is a partition of $r_{\tau}(\bar{a},\mathcal{T}_{n})$ it follows that $\mu\leq f_{4}(n)$. Let
$\mathbf{X}_{n}^{\varepsilon}(\bar{a})=\bigcap_{i=1}^{\mu}\mathbf{X}_{n}^{\varepsilon,i}(\bar{a}).$
[…]
$\mathbf{X}_{n}^{\varepsilon}=\bigcap_{\bar{a}\in(T_{n})^{|\bar{x}|}}\mathbf{X}_{n}^{\varepsilon}(\bar{a}).$
Since all elements of $\mathcal{T}_{n}$ have at most $f_{4}(n)$ children and $\mathcal{T}_{n}$ has height at most $\Delta$ (by Assumption 6.3) it follows that $\big{|}(T_{n})^{|\bar{x}|}\big{|}\leq\big{(}1+f_{4}(n)^{\Delta}\big{)}^{|\bar{x}|}=f(n)$ where $f$ is a polynomial. Thus $\mathbb{P}_{n}\big{(}\mathbf{X}_{n}^{\varepsilon}\big{)}\geq 1-f(n)e^{-C^{\prime}g_{3}(n)}$ and by Lemma 6.5 there is $C^{\prime\prime}>0$ such that $\mathbb{P}_{n}\big{(}\mathbf{X}_{n}^{\varepsilon}\big{)}\geq 1-e^{-C^{\prime\prime}g_{3}(n)}$ for all sufficiently large $n$.
Claim. Suppose that $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon}$. If $n$ is sufficiently large then, for all $\bar{a}\in(T_{n})^{|\bar{x}|}$,
$(\alpha-3\varepsilon)\sum_{b\in r_{\tau}(\bar{a},\mathcal{A})}|p_{\tau}(\bar{a},b,\mathcal{A})|\leq\sum_{b\in r(\bar{a},\mathcal{A})}|p_{\tau}(\bar{a},b,\mathcal{A})|\leq(\alpha+3\varepsilon)\sum_{b\in r_{\tau}(\bar{a},\mathcal{A})}|p_{\tau}(\bar{a},b,\mathcal{A})|.$
Proof. Let $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon}$, so $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon,i}$ for all $i=1,\ldots,\mu$, and let $\bar{a}\in(T_{n})^{|\bar{x}|}$. Suppose that $r_{\tau}(\bar{a},\mathcal{A})\neq\emptyset$, since otherwise the inequalities are trivial. Fix any $i\in\{1,\ldots,\mu\}$. By (7.9) there is a number $m_{i}\geq g_{1}(n)$ such that if $b\in P_{i}$ then $m_{i}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{i}+4g_{2}(n)$ and hence $m_{i}\leq|p_{\tau}(\bar{a},b,\mathcal{A})|<m_{i}+4g_{2}(n)$.
Since $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon,i}(\bar{a})$ we get, by the use of (7.10),
$\frac{\sum_{b\in r(\bar{a},\mathcal{A})\cap P_{i}}|p_{\tau}(\bar{a},b,\mathcal{A})|}{\sum_{b\in P_{i}}|p_{\tau}(\bar{a},b,\mathcal{A})|}\leq\frac{\sum_{b\in r(\bar{a},\mathcal{A})\cap P_{i}}(m_{i}+4g_{2}(n))}{\sum_{b\in P_{i}}m_{i}}\leq$
$\frac{|r(\bar{a},\mathcal{A})\cap P_{i}|(m_{i}+4g_{2}(n))}{|P_{i}|m_{i}}\leq(\alpha+\varepsilon)\frac{(m_{i}+4g_{2}(n))}{m_{i}}=$
$(\alpha+\varepsilon)\bigg{(}1+\frac{4g_{2}(n)}{m_{i}}\bigg{)}\leq$
$(\alpha+\varepsilon)(1+\varepsilon)\leq(\alpha+3\varepsilon)\ \text{ if $n$ is large enough }$
because $m_{i}\geq g_{1}(n)$ and by Assumption 6.3 $g_{2}(n)/g_{1}(n)\to 0$ as $n\to\infty$.
In a similar way we get
$(\alpha-3\varepsilon)\leq\frac{\sum_{b\in r(\bar{a},\mathcal{A})\cap P_{i}}|p_{\tau}(\bar{a},b,\mathcal{A})|}{\sum_{b\in P_{i}}|p_{\tau}(\bar{a},b,\mathcal{A})|}$
and hence
$(\alpha-3\varepsilon)\sum_{b\in P_{i}}|p_{\tau}(\bar{a},b,\mathcal{A})|\ \leq\ \sum_{b\in r(\bar{a},\mathcal{A})\cap P_{i}}|p_{\tau}(\bar{a},b,\mathcal{A})|\ \leq\ (\alpha+3\varepsilon)\sum_{b\in P_{i}}|p_{\tau}(\bar{a},b,\mathcal{A})|.$
Since
$\sum_{b\in r(\bar{a},\mathcal{A})}|p_{\tau}(\bar{a},b,\mathcal{A})|=\sum_{i=1}^{\mu}\ \sum_{b\in r(\bar{a},\mathcal{A})\cap P_{i}}|p_{\tau}(\bar{a},b,\mathcal{A})|\quad\text{ and}$
$\sum_{b\in r_{\tau}(\bar{a},\mathcal{A})}|p_{\tau}(\bar{a},b,\mathcal{A})|=\sum_{i=1}^{\mu}\ \sum_{b\in P_{i}}|p_{\tau}(\bar{a},b,\mathcal{A})|$
the claim follows. $\square$
Since $\mathrm{rank}_{z}(r)=1$ it follows that $\mathrm{rank}_{\bar{u}}(p)=\kappa$. By the induction hypothesis, $(p,p_{\tau},r)$ is $\beta$-balanced with respect to $\mathbb{G}$, for some $\beta$. Let
$\mathbf{Y}_{n}^{\varepsilon}=\big{\{}\mathcal{A}\in\mathbf{W}_{n}:(p,p_{\tau},r)\text{ is $(\beta,\varepsilon)$-balanced in $\mathcal{A}$}\big{\}},$
so $\lim_{n\to\infty}\mathbb{P}_{n}\big{(}\mathbf{Y}_{n}^{\varepsilon}\big{)}=1$. Suppose that $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon}\cap\mathbf{Y}_{n}^{\varepsilon}$ and $\bar{a}\in(T_{n})^{|\bar{x}|}$. Then
$|p(\bar{a},\mathcal{A})|\ =\ \sum_{b\in r(\bar{a},\mathcal{A})}|p(\bar{a},b,\mathcal{A})|\ \leq\ (\beta+\varepsilon)\sum_{b\in r(\bar{a},\mathcal{A})}|p_{\tau}(\bar{a},b,\mathcal{A})|$
$\leq\ (\beta+\varepsilon)(\alpha+\varepsilon)\sum_{b\in r_{\tau}(\bar{a},\mathcal{A})}|p_{\tau}(\bar{a},b,\mathcal{A})|\quad\quad\text{ by the claim}$
$=\ (\beta+\varepsilon)(\alpha+\varepsilon)|p_{\tau}(\bar{a},\mathcal{A})|\ \leq\ (\alpha\beta+3\varepsilon)|p_{\tau}(\bar{a},\mathcal{A})|.$
In a similar way we get $(\alpha\beta-3\varepsilon)|p_{\tau}(\bar{a},\mathcal{A})|\leq|p(\bar{a},\mathcal{A})|$. Hence, for every $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon}\cap\mathbf{Y}_{n}^{\varepsilon}$, $(p,p_{\tau},q)$ is is $(\alpha\beta,3\varepsilon)$-balanced in $\mathcal{A}$. As $\varepsilon>0$ can be chosen as small as we like it and $\lim_{n\to\infty}\mathbb{P}_{n}\big{(}\mathbf{X}_{n}^{\varepsilon}\cap\mathbf{Y}_{n}^{\varepsilon}\big{)}=1$ it follows that $(p,p_{\tau},q)$ is $\alpha\beta$-balanced with respect to $\mathbb{G}$. This completes the proof of Proposition 7.19. $\square$
Corollary 7.20.
Let $p_{\tau}(\bar{x},\bar{y})$ be a closure type over $\tau$ and let $p(\bar{x},\bar{y})$ and $q(\bar{x})$ be complete closure types over $\sigma$. Then $(p,p_{\tau},q)$ is balanced with respect to $\mathbb{G}$.
Proof. Let $p_{\tau}$, $p$ and $q$ be as assumed. By Remark 7.5 we may assume that $p\wedge p_{\tau}\wedge q$ is cofinally satisfiable, hence consistent. By Remark 7.6 we may also assume that $\mathrm{rank}_{\bar{y}}(p_{\tau})>0$. Since $p$ is a complete closure type over $\sigma$, $p(\bar{x},\bar{y})\models_{tree}p_{\tau}(\bar{x},\bar{y})\wedge q(\bar{x})$. By Remark 5.19, there are sequences of variables $\bar{u}$ and $\bar{v}$ such that $\bar{y}$ is a subsequence of $\bar{u}\bar{v}$, and a complete closure type over $\tau$, say $p^{*}_{\tau}(\bar{x},\bar{u},\bar{v})$, such that $p^{*}_{\tau}$ is self-contained, $\bar{v}$-independent, and for all $n$, all $\mathcal{A}\in\mathbf{W}_{n}$ and all $\bar{a}\in(T_{n})^{|\bar{x}|}$, if $p_{\tau}(\bar{a},\mathcal{A})\neq\emptyset$ then there is a unique $\bar{b}\in(T_{n})^{|\bar{u}|}$ such that $|p_{\tau}(\bar{a},\mathcal{A})|=|p^{*}_{\tau}(\bar{a},\bar{b},\mathcal{A})|$ (and we have $\bar{b}\in\mathrm{cl}_{\mathcal{T}_{n}}(\bar{a})$). By Remark 5.19 again there is a self-contained and $\bar{v}$-independent closure type $p^{*}(\bar{x},\bar{u},\bar{v})$ over $\sigma$ such that $p^{*}\models_{tree}p^{*}_{\tau}$ and if $p(\bar{a},\mathcal{A})\neq\emptyset$ then there is a unique $\bar{b}\in(T_{n})^{|\bar{u}|}$ such that $|p(\bar{a},\mathcal{A})|=|p^{*}(\bar{a},\bar{b},\mathcal{A})|$. Since $p^{*}\models_{tree}p^{*}_{\tau}$ this $\bar{b}$ must be the same tuple for which $|p_{\tau}(\bar{a},\mathcal{A})|=|p^{*}_{\tau}(\bar{a},\bar{b},\mathcal{A})|$. By Proposition 7.19, $(p^{*},p^{*}_{\tau},q^{*})$ is $\alpha$-balanced with respect to $\mathbb{G}$ for some $\alpha$. Because of the identities of cardinalities above it follows that also $(p,p_{\tau},q)$ is $\alpha$-balanced with respect to $\mathbb{G}$. $\square$
Corollary 7.21.
Let $p_{\tau}(\bar{x},\bar{y})$ be a closure type over $\tau$, let $p(\bar{x},\bar{y})$ be a (possibly not complete) closure type over $\sigma$, and $q(\bar{x})$ be a complete closure type over $\sigma$. Then $(p,p_{\tau},q)$ is balanced with respect to $\mathbb{G}$.
Proof. By Remark 7.5 we may assume that $p\wedge p_{\tau}\wedge q$ is cofinally satisfiable. Let $p_{1}(\bar{x},\bar{y}),\ldots p_{m}(\bar{x},\bar{y})$ be an enumeration of all, up to equivalence, complete closure types over $\sigma$ such that, for each $i$, $p_{i}(\bar{x},\bar{y})\models_{tree}p(\bar{x},\bar{y})\wedge q(\bar{x})$. So $p(\bar{x},\bar{y})\wedge q(\bar{x})$ is equivalent, in every $\sigma$-structure that expands a tree, to $\bigvee_{i=1}^{m}p_{i}(\bar{x},\bar{y})$ and we may assume that if $i\neq j$ then $p_{i}\wedge p_{j}$ is inconsistent. It follows that if $\mathcal{A}\in\mathbf{W}_{n}$, $\bar{a}\in(T_{n})^{|\bar{x}|}$ and $\mathcal{A}\models q(\bar{a})$, then $|p(\bar{a},\mathcal{A})|=\sum_{i=1}^{m}|p_{i}(\bar{a},\mathcal{A})|$. By Corollary 7.20, for each $i$, $(p_{i},p_{\tau},q)$ is $\alpha_{i}$-balanced with respect to $\mathbb{G}$ for some $\alpha_{i}$. It is now straightforward to verify that $(p,p_{\tau},q)$ is $(\alpha_{1}+\ldots+\alpha_{m})$-balanced with respect to $\mathbb{G}$. $\square$
Definition 7.22.
Let $p(\bar{x},\bar{y})$ be a closure type over $\sigma$. We call $p$ $\bar{y}$-positive (with respect to $\mathbb{G}$) if the following holds: if $q(\bar{x})$ is a complete closure type over $\sigma$, $p\wedge q$ is cofinally satisfiable, and $p_{\tau}(\bar{x},\bar{y})=p\negthickspace\upharpoonright\negthickspace\tau$, then there is $\alpha>0$ such that $(p,p_{\tau},q)$ is $\alpha$-balanced with respect to $\mathbb{G}$.
Remark 7.23.
(i) Let $p(\bar{x},\bar{y})$ be a closure type over $\tau$, so in particular $p$ is an incomplete closure type over $\sigma$ (if $\tau$ is a proper subset of $\sigma$). Then $p\negthickspace\upharpoonright\negthickspace\tau$ is the same as (or at least equivalent to) $p$ and therefore it is evident from the definitions that $(p,p\negthickspace\upharpoonright\negthickspace\tau,q)$ is 1-balanced with respect to $\mathbb{G}$ and hence $p$ is $\bar{y}$-positive.
(ii) Let $p(\bar{x},\bar{y})$ be a closure type over $\tau$ and suppose that $\mathrm{rank}_{\bar{y}}(p)=0$. The argument in Remark 7.6 shows that if $q(\bar{x})$ is a complete closure type over $\sigma$ and $p\wedge q$ is cofinally satisfiable (hence $q(\bar{x})\models\exists\bar{y}p(\bar{x},\bar{y})$), then $(p,p\negthickspace\upharpoonright\negthickspace\tau,q)$ is 1-balanced with respect to $\mathbb{G}$. Hence $p$ is $\bar{y}$-positive.
Proposition 7.24.
Let $p(\bar{x},\bar{y})$ and $r(\bar{x},\bar{y})$ be (possibly incomplete) closure types over $\sigma$, where $r$ is $\bar{y}$-positive with respect to $\mathbb{G}$, and let $q(\bar{x})$ be complete closure type over $\sigma$, Then $(p,r,q)$ is balanced with respect to $\mathbb{G}$.
Proof. By Remark 7.5 we may assume that $p\wedge r\wedge q$ is cofinally satisfiable. Let $p_{\tau}(\bar{x},\bar{y})=p\negthickspace\upharpoonright\negthickspace\tau$. Then $p_{\tau}$ is equivalent to $r\negthickspace\upharpoonright\negthickspace\tau$, $p\models p_{\tau}$ and $r\models p_{\tau}$. Since $r$ is $\bar{y}$-positive it follows that $(r,p_{\tau},q)$ is $\alpha$-balanced for some $\alpha>0$. By Corollary 7.21, $(p\wedge r,p_{\tau},q)$ is $\beta$-balanced for some $\beta$. Now it is straightforward to verify that $(p,r,q)$ is $\beta/\alpha$-balanced. $\square$
Remark 7.25.
Let $p(\bar{x},\bar{y})$, $r(\bar{x},\bar{y})$, and $q(\bar{x})$ be as in Proposition 7.24 and suppose that $p\wedge r\wedge q$ is cofinally satisfiable. By analysing the proofs of Section 7.4 (this section) we see that if $(p,r,q)$ is 00-balanced, then $(p,r)$ converges to 00. If $(p,r)$ is also eventually constant, then it must be eventually constant with value 00. This implies that for all sufficiently large $n$ and all $\bar{a}\in(T_{n})^{|\bar{x}|}$, $p(\bar{a},\mathcal{A})\cap r(\bar{a},\mathcal{A})=\emptyset$.
By involving the conclusions of Remark 7.15 we can now conclude the following, by induction on the height of $\mathbb{G}$: Suppose that for all $R\in\sigma\setminus\tau$, $\theta_{R}$ is a closure basic formula. If $p(\bar{x},\bar{y})$, $r(\bar{x},\bar{y})$, and $q(\bar{x})$ are as in Proposition 7.24 and $(p,r,q)$ is 00-balanced, then for all sufficiently large $n$ and all $\bar{a}\in(T_{n})^{|\bar{x}|}$, $p(\bar{a},\mathcal{A})\cap r(\bar{a},\mathcal{A})=\emptyset$.
7.5. Asymptotic elimination of aggregation functions
In this subsection we prove statement (A) above, that is, we prove that given assumptions 6.3 and 7.8 (the induction hypothesis) and the results proved earlier in Section7 we can asymptotically eliminate aggregation functions, provided that some conditions are satisfied. Some motivation for these conditions are given in the beginning of Section 8.
Proposition 7.26.
Suppose that assumptions 6.3 and 7.8 hold.
(i) Let $\varphi(\bar{x})\in PLA^{*}(\sigma)$ and suppose that for every subformula of $\varphi$ of the form
$F\big{(}\varphi_{1}(\bar{y},\bar{z}),\ldots,\varphi_{m}(\bar{y},\bar{z}):\bar{z}:\chi_{1}(\bar{y},\bar{z}),\ldots,\chi_{m}(\bar{y},\bar{z})\big{)},$
it holds that, for all $i=1,\ldots,m$, $\chi_{i}(\bar{y},\bar{z})$ is a $\bar{z}$-positive closure type over $\sigma$, and either $F$ is continuous, or $\mathrm{rank}_{\bar{z}}(\chi_{i})=0$ for all $i=1,\ldots,m$ and $F$ is admissible. Then $\varphi(\bar{x})$ is asymptotically equivalent to a closure-basic formula over $\sigma$ with respect to the sequence of probability distributions $\mathbb{P}=(\mathbb{P}_{n}:n\in\mathbb{N}^{+})$ induced by $\mathbb{G}$.
(ii) Suppose that for every $R\in\sigma\setminus\sigma_{\rho}$ the formula $\theta_{R}$ associated to $R$ (in $\mathbb{G}$) is a closure-basic formula. Then, in part (i), we can replace the occurence of “continuous” by the weaker condition “admissible”.
Proof. We prove part (i) by induction on the complexity of $\varphi(\bar{x})$. In the base case we assume that $\varphi(\bar{x})$ is aggregation-free. Then it follows from Lemma 5.22 that $\varphi(\bar{x})$ is equivalent, hence asymptotically equivalent, to a closure-basic formula over $\sigma$.
Next, suppose that $\varphi(\bar{x})$ has the form $\mathsf{C}\big{(}\psi_{1}(\bar{x}),\ldots,\psi_{k}(\bar{x})\big{)}$ where $\mathsf{C}:[0,1]^{k}\rightarrow[0,1]$ is continuous. By the induction hypothesis, each $\psi_{i}(\bar{x})$ is asymptotically equivalent to a closure-basic formula $\psi^{\prime}_{i}(\bar{x})$. By Lemma 5.22, $\mathsf{C}\big{(}\psi^{\prime}_{1}(\bar{x}),\ldots,\psi^{\prime}_{k}(\bar{x})\big{)}$ is asymptotically equivalent to a closure-basic formula $\chi(\bar{x})$. Since $\mathsf{C}$ is continuous it follows that $\mathsf{C}\big{(}\psi_{1}(\bar{x}),\ldots,\psi_{k}(\bar{x})\big{)}$ is asymptotically equivalent to $\mathsf{C}\big{(}\psi^{\prime}_{1}(\bar{x}),\ldots,\psi^{\prime}_{k}(\bar{x})\big{)}$. By transitivity of asymptotic equivalence, $\mathsf{C}\big{(}\psi_{1}(\bar{x}),\ldots,\psi_{k}(\bar{x})\big{)}$ is asymptotically equivalent to $\chi(\bar{x})$.
Finally, suppose that $\varphi(\bar{x})$ is of the form
$F\big{(}\varphi_{1}(\bar{x},\bar{y}),\ldots,\varphi_{m}(\bar{x},\bar{y}):\bar{y}:\chi_{1}(\bar{x},\bar{y}),\ldots,\chi_{m}(\bar{x},\bar{y})\big{)}$
where $F:\big{(}[0,1]^{<\omega}\big{)}^{m}\rightarrow[0,1]$ is continuous and, for all $i=1,\ldots,m$, $\chi_{i}(\bar{x},\bar{y})$ is a $\bar{y}$-positive closure type over $\sigma$. By the induction hypothesis, each $\varphi_{i}(\bar{x},\bar{y})$ is asymptotically equivalent to a closure-basic formula $\varphi^{\prime}_{i}(\bar{x},\bar{y})$ over $\sigma$. Then each $\varphi^{\prime}_{i}(\bar{x},\bar{y})$ has the form
$\bigwedge_{j=1}^{s_{i}}\big{(}\varphi_{i,j}(\bar{x},\bar{y})\rightarrow c_{i,j}\big{)}$
where each $\varphi_{i,j}(\bar{x},\bar{y})$ is a complete closure type over $\sigma$ and $c_{i,j}\in[0,1]$.
We will use Theorem 4.8 to show that $\varphi(\bar{x})$ is asymptotically equivalent to a closure-basic formula. We first assume that $F$ is continuous and show that $\varphi(\bar{x})$ is asymptotically equivalent to a closure basic formula. Then we make some observations from which we can make the other conclusions of part (i) of the proposition. In order to use Theorem 4.8 we need to define appropriate subsets $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ and show that Assumption 4.7 holds with this choice of $L_{0}$ and $L_{1}$. Let $L_{0}$ be the the set of all complete closure types over $\sigma$ and let $L_{1}$ be the set of all (not necessarily complete) closure types over $\sigma$. (Note that $L_{0}$ also contains closure types without free variables and recall that these express what relations the root satisfies. Thus the argument that follows makes sense also if $\bar{x}$ is empty.) Note that with this choice of $L_{0}$, each $\varphi^{\prime}_{i}$ is an $L_{0}$-basic formula. We now show that Assumption 4.7 holds for $L_{0}$ and $L_{1}$. Part (1) of Assumption 4.7 follows from Lemma 5.22. It remains to verify that part (2) of Assumption 4.7 holds.
For each $p(\bar{x},\bar{y})\in L_{0}$ define $L_{p(\bar{x},\bar{y})}$ to be the set of all $\bar{y}$-positive $\lambda(\bar{x},\bar{y})\in L_{1}$. (So $L_{p(\bar{x},\bar{y})}$ depends only on the free variables $\bar{x}$ and $\bar{y}$ but not on other properties of $p$.) Suppose that $p_{i}(\bar{x},\bar{y})\in L_{0}$ for $i=1,\ldots,k$ and let $\lambda_{i}(\bar{x},\bar{y})\in L_{p_{i}(\bar{x},\bar{y})}$ for $i=1,\ldots,k$, so $\lambda_{i}(\bar{x},\bar{y})$ is a $\bar{y}$-positive closure type over $\sigma$. Let $q_{i}(\bar{x})\in L_{0}$, $i=1,\ldots,s$, enumerate all, up to equivalence, complete closure types over $\sigma$ in the variables $\bar{x}$ and we may assume that $q_{i}(\bar{x})$ is not equivalent to $q_{j}(\bar{x})$ if $i\neq j$ (otherwise just remove some $q_{i}(\bar{x})$).
Observe that by Assumption 6.3, Remark 6.4 and the definition of $\mathbf{W}_{n}$, it follows that if $p(\bar{x},\bar{y})$ and $q(\bar{x})$ are complete closure types over $\sigma$ and $p(\bar{x},\bar{y})\wedge q(\bar{x})$ is satisfied in some $\mathcal{A}\in\mathbf{W}_{n}$ for some $n$, then for all sufficiently large $n$ the same formula is satisfied in some $\mathcal{A}\in\mathbf{W}_{n}$. Now let $\lambda^{\prime}_{1}(\bar{x}),\ldots,\lambda^{\prime}_{t}(\bar{x})$ be an enumeration of all $q_{i}(\bar{x})$ such that, for some $j$, $q_{i}(\bar{x})\wedge\lambda_{j}(\bar{x},\bar{y})$ is not satisfied in any $\mathcal{A}\in\mathbf{W}_{n}$ for any $n$. With these choices we have, for every $n$ and every $\mathcal{A}\in\mathbf{W}_{n}$, that
(7.11) $\mathcal{A}\models\forall\bar{x}\big{(}\bigvee_{i=1}^{s}q_{i}(\bar{x})\big{)},$
$\mathcal{A}\models\forall\bar{x}\neg\big{(}q_{i}(\bar{x})\wedge q_{j}(\bar{x})\big{)}\text{ if $i\neq j$, and }$
$\mathcal{A}\models\Big{(}\bigvee_{i=1}^{k}\neg\exists\bar{y}\lambda_{i}(\bar{x},\bar{y})\Big{)}\leftrightarrow\Big{(}\bigvee_{i=1}^{t}\lambda^{\prime}_{i}(\bar{x})\Big{)}.$
We now verify that condition (d) of part (2) of Assumption 4.7 holds. It will then be immediate from (7.11) that also conditions (a), (b) and (c) of part (2) of Assumption 4.7 hold. From Proposition 7.24 it follows that, for all $i=1,\ldots,k$ and all $j=1,\ldots,s$, the triple $(p_{i},\lambda_{i},q_{j})$ is $\alpha_{i,j}$-balanced for some $\alpha_{i,j}$. This means that for all $\varepsilon>0$ and all $n$, if
$\mathbf{Y}_{n,\varepsilon}^{i,j}=\big{\{}\mathcal{A}\in\mathbf{W}_{n}:(p_{i},\lambda_{i},q_{j})\text{ is $(\alpha_{i,j},\varepsilon)$-balanced in }\mathcal{A}\big{\}}$
then $\lim_{n\to\infty}\mathbb{P}_{n}(\mathbf{Y}_{n,\varepsilon}^{i,j})=1$. Let $\mathbf{Y}_{n}^{\varepsilon}=\bigcap_{i=1}^{k}\bigcap_{j=1}^{s}\mathbf{Y}_{n,\varepsilon}^{i,j}$. Then $\lim_{n\to\infty}\mathbb{P}_{n}(\mathbf{Y}_{n,\varepsilon}^{i,j})=1$ and, for every choice of $i$ and $j$, if $\bar{a}\in(T_{n})^{|\bar{x}|}$, $\mathcal{A}\in\mathbf{Y}_{n}^{\varepsilon}$ and $\mathcal{A}\models q_{j}(\bar{a})$, then
$(\alpha_{i,j}-\varepsilon)|\lambda_{i}(\bar{a},\mathcal{A})|\leq|p_{i}(\bar{a},\mathcal{A})\cap\lambda_{i}(\bar{a},\mathcal{A})|\leq(\alpha_{i,j}+\varepsilon)|\lambda_{i}(\bar{a},\mathcal{A})|.$
Hence condition (d) of part (2) of Assumption 4.7 holds. As (7.11) holds for all $\mathcal{A}\in\mathbf{W}_{n}$ it also holds for all $\mathcal{A}\in\mathbf{Y}_{n}^{\varepsilon}$. Hence conditions (a), (b) and (c) of part (2) of Assumption 4.7 hold. Now part (i) of Theorem 4.8 implies that $\varphi(\bar{x})$ is asymptotically equivalent to an $L_{0}$-basic formula, or equivalently, to a closure-basic formula over $\sigma$.
So far we assumed that $F$ is continuous. Now suppose that $F$ is admissible and that $\mathrm{rank}_{\bar{y}}(\lambda_{i})=0$ for all $i=1,\ldots,k$. Then for all $i=1,\ldots,k$ and $j=1,\ldots,s$, either $q_{j}(\bar{x})\wedge\lambda_{i}(\bar{x},\bar{y})\models_{tree}p_{i}(\bar{x},\bar{y})$, or $q_{j}(\bar{x})\wedge\lambda_{i}(\bar{x},\bar{y})\models_{tree}\neg p_{i}(\bar{x},\bar{y})$. So if $\alpha_{i,j}=0$ then, for all $n$, all $\bar{a}\in(T_{n})^{|\bar{x}|}$, and all $\mathcal{A}\in\mathbf{W}_{n}$, if $\mathcal{A}\models q_{j}(\bar{a})$ then $p_{i}(\bar{a},\mathcal{A})\cap\lambda_{i}(\bar{a},\mathcal{A})=\emptyset$. This means that the extra condition in part (ii) of Theorem 4.8 is satisfied and therefore $\varphi(\bar{x})$ is asymptotically equivalent to a closure-basic formula.
Part (ii) follows from the conclusion of Remark 7.25 and the above argument (where we use part (ii) of Theorem 4.8). $\square$
Remark 7.27.
Recall that we have assumed that $\mathbb{G}$ is a $PLA^{*}(\sigma)$-network of height $\rho+1$. Suppose that $\sigma\subset\sigma^{+}$ and that $\mathbb{G}^{+}$ is a $PLA^{*}(\sigma^{+})$-network of height $\rho+2$ such that $\mathbb{G}$ is the subnetwork of $\mathbb{G}^{+}$ which is induced by $\sigma$. For every $R\in\sigma^{+}\setminus\sigma$, let $\theta_{R}$ be the $PLA^{*}(\sigma)$-formula which $\mathbb{G}^{+}$ associates to $R$. Suppose that for each $R\in\sigma^{+}\setminus\sigma$, $\theta_{R}$ satisfies the assumptions on $\varphi$ in Proposition 7.26. Then Proposition 7.26 implies that part (1) of Assumption 7.8 holds if $\sigma_{\rho}$, $\sigma$ and $\mathbb{G}_{\rho}$ are replaced by $\sigma$, $\sigma^{+}$ and $\mathbb{G}$, respectively. Hence statement (A) is proved.
8. The main results
In this section we prove the main results. The reader may note that we only “asymptotically eliminate” aggregations that are conditioned on closure types. From the perspective of first-order logic this may seem like a strong constraint since first-order quantifications range over the whole domain. However, in applications one is often not interested in searching through the whole domain (of a database for example), but instead one may be interested in some part of the domain (defined by some constraints). In the current context, when conditioning aggregations on closure types we restrict the range of aggregations to vertices, or tuples of vertices, on specified levels and, possibly, with specified ancestors, in the underlying tree. Also note that, with the assumptions that we have adopted on the sequence $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ of underlying trees, the number of vertices on level $l+1$ divided by the number of vertices on level $l$ tends to infinity (unless level $l+1$ is empty). So if we use a continuous aggregation function in an aggregation, then the highest level (in the tree) that the aggregation ranges over will, up to a small discrepancy, determine the output.
Theorem 8.1.
Let $\tau\subset\sigma$, let $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ be a sequence of trees such that Assumption 6.3 holds, and let $\mathbb{G}$ be a $PLA^{*}(\sigma)$-network based on $\tau$ (and recall that $\mathbb{G}$ associates a formula $\theta_{R}\in PLA^{*}(\mathrm{par}(R))$ to every $R\in\sigma\setminus\tau$). Furthermore let $\mathbb{P}_{n}$ be the probability distribution that $\mathbb{G}$ induces on $\mathbf{W}_{n}$, the set of expansions of $\mathcal{T}_{n}$ to $\sigma$.
Suppose that for every $R\in\sigma\setminus\tau$ and every subformula of $\theta_{R}$ of the form
(8.1) $F\big{(}\varphi_{1}(\bar{y},\bar{z}),\ldots,\varphi_{m}(\bar{y},\bar{z}):\bar{z}:\chi_{1}(\bar{y},\bar{z}),\ldots,\chi_{m}(\bar{y},\bar{z})\big{)},$
it holds that, for all $i=1,\ldots,m$, $\chi_{i}(\bar{y},\bar{z})$ is a (not necessary complete) $\bar{z}$-positive closure type over $\sigma$, and either $F$ is continuous, or $\mathrm{rank}_{\bar{z}}(\chi_{i})=0$ for all $i=1,\ldots,m$ and $F$ is admissible.
Let $\varphi(\bar{x})\in PLA^{*}(\sigma)$ and suppose that for every subformula of $\varphi(\bar{x})$ of the form (8.1) the same condition holds. Then:
(i) $\varphi(\bar{x})$ is asymptotically equivalent to a closure-basic formula over $\sigma$ with respect to $(\mathbb{P}_{n}:n\in\mathbb{N}^{+})$.
(ii) For every closure type $p(\bar{x})$ over $\tau$ there are $k\in\mathbb{N}^{+}$, $c_{1},\ldots,c_{k}\in[0,1]$, and $\alpha_{1},\ldots,\alpha_{k}\in[0,1]$ such that for every $\varepsilon>0$ there is $n_{0}$ such that if $n\geq n_{0}$, $\bar{a}\in(T_{n})^{|\bar{x}|}$ and $\mathcal{T}_{n}\models p(\bar{a})$, then
$\mathbb{P}_{n}\Big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\mathcal{A}(\varphi(\bar{a}))\in\bigcup_{i=1}^{k}[c_{i}-\varepsilon,c_{i}+\varepsilon]\big{\}}\Big{)}\geq 1-\varepsilon\text{ and, for all $i=1,\ldots,k$,}$
$\mathbb{P}_{n}\big{(}\{\mathcal{A}\in\mathbf{W}_{n}:\mathcal{A}(\varphi(\bar{a}))\in[c_{i}-\varepsilon,c_{i}+\varepsilon]\}\big{)}\in[\alpha_{i}-\varepsilon,\alpha_{i}+\varepsilon].$
Proof. (i) Suppose that $\tau$ is a proper subset of $\sigma$. $\mathbb{G}$ be a $PLA^{*}(\sigma)$-network $\mathbb{G}$ based on $\tau$ with height $\rho+1$ (where $\rho\geq-1$). Let $\sigma_{\rho}=\{R\in\sigma\setminus\tau:\text{ $R$ is on level $l$ for some $l\leq\rho$}\}$. We use induction on the height $\rho+1$. The base case $\rho+1=0$ (i.e. $\rho=-1$) is equivalent to assuming that $\sigma_{\rho}=\tau$. It was pointed out in Remark 7.9 that, by Lemma 7.7, Assumption 7.8 holds if $\sigma_{\rho}=\tau$.
By the observations in remarks 7.14, 7.18 and 7.27 it follows that if Assumption 7.8 holds for all $\sigma\supseteq\tau$ and all $PLA^{*}(\sigma)$-networks based on $\tau$ with height $\rho+1$ such that, for all $R\in\sigma\setminus\tau$, $\theta_{R}$ satisfies the conditions of Theorem 8.1, then Assumption 7.8 also holds for all $\sigma\supseteq\tau$ and all $PLA^{*}(\sigma)$-networks based on $\tau$ with height $\rho+2$ such that, for all $R\in\sigma\setminus\tau$, $\theta_{R}$ satisfies the conditions of Theorem 8.1. Thus Assumption 7.8 holds for all $\sigma\supseteq\tau$ and all $PLA^{*}(\sigma)$-networks based on $\tau$ such that, for all $R\in\sigma\setminus\tau$, $\theta_{R}$ satisfies the conditions of Theorem 8.1. Proposition 7.26 was derived by using only parts (2) and (3) of Assumption 7.8, so it follows that if $\varphi(\bar{x})$ is as assumed in the theorem, then it is asymptotically equivalent to a closure-basic formula.
Since all results in Section 7 were derived from Assumption 7.8 we can, by induction, conclude that all results in that section hold for every $\sigma\supseteq\tau$ and every $PLA^{*}(\sigma)$-network such that, for all $R\in\sigma\setminus\tau$, $\theta_{R}$ satisfies the conditions of Theorem 8.1.
(ii) Let $p(\bar{x})$ be a closure type over $\tau$. By part (i), $\varphi(\bar{x})$ is asymptotically equivalent to a closure-basic formula
$\bigwedge_{i=1}^{m}(\varphi_{i}(\bar{x})\to c_{i}),$
where each $\varphi_{i}(\bar{x})$ is a complete closure type over $\sigma$. Without loss of generality we may assume that $\varphi_{i}(\bar{x})$, $i=1,\ldots,m$, enumerate all, up to equivalence, complete closure types over $\sigma$ that are cofinally satisfiable and that $\varphi_{i}\wedge\varphi_{j}$ is inconsistent if $i\neq j$. Let $p(\bar{x})$ be a closure type over $\tau$. By Proposition 7.12 (which as noted above holds under the assumptions of the theorem), for all $i$, $(\varphi_{i},p)$ converges to some $\alpha_{i}$. So for every $\varepsilon>0$, if $n$ is large enough, $\bar{a}\in(T_{n})^{|\bar{x}|}$ and $\mathcal{T}_{n}\models p(\bar{a})$, then
$\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi_{i}(\bar{a})}\ |\ \mathbf{E}_{n}^{p(\bar{a})}\big{)}\in[\alpha_{i}-\varepsilon,\alpha_{i}+\varepsilon].$
and (as $\varphi(\bar{x})$ and $\bigwedge_{i=1}^{m}(\varphi_{i}(\bar{x})\to c_{i})$ are asymptotically equivalent)
$\mathbb{P}_{n}\Big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:|\mathcal{A}(\varphi(\bar{a}))-\mathcal{A}(\bigwedge_{i=1}^{m}(\varphi_{i}(\bar{a})\to c_{i}))|\leq\varepsilon\big{\}}\Big{)}\geq 1-\varepsilon.$
Note that if $\mathcal{A}\models\varphi_{i}(\bar{a})$ then $\mathcal{A}(\bigwedge_{i=1}^{m}(\varphi_{i}(\bar{a})\to c_{i}))=c_{i}$. Since we assume that $\varphi_{i}(\bar{x})$, $i=1,\ldots,m$, enumerate all, up to equivalence, complete closure types over $\sigma$ that are cofinally satisfiable it follows that for all sufficiently large $n$
$\mathbb{P}_{n}\Big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\mathcal{A}(\varphi(\bar{a}))\in\bigcup_{i=1}^{k}[c_{i}-\varepsilon,c_{i}+\varepsilon]\big{\}}\Big{)}\geq 1-\varepsilon.$
Let $c\in\{c_{1},\ldots,c_{m}\}$ and for simplicity of notation suppose that, for some $1\leq s\leq m$, $c=c_{i}$ if $i\leq s$ and $c\neq c_{i}$ if $i>s$. Let $\beta_{c}=\alpha_{1}+\ldots+\alpha_{s}$. It now follows that if $n$ is large enough and $\mathcal{T}_{n}\models p(\bar{a})$ then
$\mathbb{P}_{n}\big{(}\{\mathcal{A}\in\mathbf{W}_{n}:\mathcal{A}(\varphi(\bar{a}))\in[c-\varepsilon,c+\varepsilon]\}\big{)}\in[\beta_{c}-s\varepsilon,\beta_{c}+s\varepsilon].$
The claim now follows since $\varepsilon>0$ can be chosen as small as we like. $\square$
Recall that the aggregation functions min and max are admissible, but not continuous. Therefore Theorem 8.1 only applies to min and max if they are used together with a conditioning closure type $\chi(\bar{y},z)$ with $z$-rank 0. The next corollary states that if we strengthen the assumptions on the $PLA(\sigma)$-network $\mathbb{G}$ (thus limiting the range of distributions that such $\mathbb{G}$ can induce), then the conclusions (i) and (ii) of Theorem 8.1 hold if the formula $\varphi(\bar{x})$ uses (only) admissible aggregation functions. In particular the conclusions of the theorem hold for first-order formulas.
Corollary 8.2.
Let $\tau$, $\sigma$, $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$, $\mathbf{W}_{n}$, $\mathbb{G}$ and $\mathbb{P}_{n}$ be as in Theorem 8.1. In particular Assumption 6.3 is adopted. Suppose that, for every $R\in\sigma\setminus\tau$, the formula $\theta_{R}$ that $\mathbb{G}$ associates to $R$ is a closure-basic formula over $\mathrm{par}(R)$ (the parents of $R$ in the underlying DAG of $\mathbb{G}$). Suppose that $\varphi(\bar{x})\in PLA^{*}(\sigma)$ and that if
$F\big{(}\varphi_{1}(\bar{y},\bar{z}),\ldots,\varphi_{m}(\bar{y},\bar{z}):\bar{z}:\chi_{1}(\bar{y},\bar{z}),\ldots,\chi_{m}(\bar{y},\bar{z})\big{)}$
is a subformula of $\varphi(\bar{x})$, then $F$ is admissible, and for all $i=1,\ldots,m$, $\chi_{i}(\bar{y},\bar{z})$ is a $\bar{z}$-positive closure type over $\sigma$. Then the conclusions (i) and (ii) of Theorem 8.1 hold.
Proof. Part (ii) follows from part (i) in exctly the same way as part (ii) of Theorem 8.1 follows from part (i) of that theorem.
Part (i) is proved in essentially the same way as part (i) of Theorem 8.1. But now we have a stronger assumption on all $\theta_{R}$, $R\in\sigma\setminus\tau$. As concluded in Remark 7.15 the stronger assumption implies the following: If $p(\bar{x},\bar{y})$ and $r(\bar{x},\bar{y})$ are closure types over $\sigma$, where $r$ is $\bar{y}$-positive, and $q(\bar{x})$ is a complete closure type over $\sigma$, and $(p,r,q)$ is 00-balanced, then for all sufficiently large $n$ and all $\bar{a}\in(T_{n})^{|\bar{x}|}$, $p(\bar{a},\mathcal{A})\cap r(\bar{a},\mathcal{A})=\emptyset$. This means that the extra condition in part (ii) of Theorem 4.8 is satisfied and therefore every admissible aggregation function can be asymptotically eliminated as in the proof of Proposition 7.26. It follows that every $\varphi(\bar{x})$ subject to the assumptions of the corollary is asymptotically equivalent to a closure-basic formula over $\sigma$. $\square$
Now we consider a lighter version of Assumption 6.3, where the functions $g_{2}$ and $g_{3}$ (in Assumption 6.3) have been removed. It turns out that “aggregations of dimension 1” can still be asymptotically eliminated if we only assume the following:
Assumption 8.3.
  1. (1)
    $\Delta\in\mathbb{N}^{+}$,
  2. (2)
    $g_{1}$ and $g_{4}$ are functions from $\mathbb{N}$ to the positive reals such that
    1. (a)
      $\lim_{n\to\infty}g_{1}(n)=\infty$,
    2. (b)
      $g_{4}$ is a polynomial,
  3. (3)
    $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ and each $\mathcal{T}_{n}$ is a tree such that
    1. (a)
      the height of $\mathcal{T}_{n}$ is $\Delta$ and all leaves of $\mathcal{T}$ are on level $\Delta$, and
    2. (b)
      every nonleaf has at least $g_{1}(n)$ children and at most $g_{4}(n)$ children.
Corollaries 8.4 and 8.6 below are analogoues of Theorem 8.1 and Corollary 8.2 that are adapted to Assumption 8.3.
Corollary 8.4.
Let $\tau\subset\sigma$, let $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ be a sequence of trees such that Assumption 8.3 holds, and let $\mathbb{G}$ be a $PLA^{*}(\sigma)$-network based on $\tau$ Furthermore let $\mathbb{P}_{n}$ be the probability distribution that $\mathbb{G}$ induces on $\mathbf{W}_{n}$, the set of expansions of $\mathcal{T}_{n}$ to $\sigma$.
Suppose that for every $R\in\sigma\setminus\tau$ and every subformula of $\theta_{R}$ of the form
(8.2) $F\big{(}\varphi_{1}(\bar{y},\bar{z}),\ldots,\varphi_{m}(\bar{y},\bar{z}):\bar{z}:\chi_{1}(\bar{y},\bar{z}),\ldots,\chi_{m}(\bar{y},\bar{z})\big{)},$
it holds that, for all $i=1,\ldots,m$, $\chi_{i}(\bar{y},\bar{z})$ is a (not necessary complete) $\bar{z}$-positive closure type over $\sigma$ and, either $F$ is continuous and $\mathrm{rank}_{\bar{z}}(\chi_{i})=1$ for all $i$, or $F$ is admissible and $\mathrm{rank}_{\bar{z}}(\chi_{i})=0$ for all $i$.
Let $\varphi(\bar{x})\in PLA^{*}(\sigma)$ and suppose that for every subformula of $\varphi(\bar{x})$ of the form (8.2) the same condition holds. Then conclusions (i) and (ii) of Theorem 8.1 hold.
Proof. The conditions of Assumption 6.3 that have been removed in Assumption 8.3 are only used in the proof of Proposition 7.19 when proving the inductive step, and in this step we considered a closure type of $\bar{y}$-rank at least 2. $\square$
Example 8.5.
We illustrate Corollary 8.4 and a contrast to Example 6.2. Let $\sigma=\tau\cup\{R\}$, where $R$ is unary, and let $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ be as in Example 6.2. Then $\mathbf{T}$ does not satisfy Assumption 6.3, but it does satisfy Assumption 8.3. We also let, as in Example 6.2, $\mathbf{W}_{n}$ be the set of all $\sigma$-structures that expand $\mathcal{T}_{n}$ and we let $\mathbb{P}_{n}$ be he uniform probability distribution on $\mathbf{W}_{n}$. ($\mathbb{P}_{n}$ is induced by a $PLA^{*}(\sigma)$-network based on $\tau$.) As in that example let $q(x)$ be a closure type over $\tau$ which expresses “$x$ is a child of the root”, and let $p(x,y)$ be a closure type over $\tau$ which expresses “$q(x)$ and $y$ is a child of $x$”. Then $\mathrm{rank}_{x}(q)=1$, $\mathrm{rank}_{(x,y)}(p)=2$, and $\exists xp(x,y)$ is a closure type over $\tau$ with $y$-rank 2. Let $\varphi$ be the sentence
$\mathrm{am}\big{(}\exists x(p(x,y)\wedge R(x)\wedge R(y)):y:\exists xp(x,y)\big{)}.$
We saw in Example 6.2 that for all sufficiently large $n$, the distribution of the values $\mathcal{A}(\varphi)$ for random $\mathcal{A}\in\mathbf{W}_{n}$ was quite different for odd $n$ compared to even $n$ (so the conclusions of Theorem 8.1 fail, showing that Assumption 6.3 is necessary for that theorem).
Let us now replace the aggregation that conditions on $\exists xp(x,y)$ with two aggregations that condition on closure types with $x$-rank, respectively $y$-rank, equal to 1, by letting $\psi$ be the sentence
$\mathrm{am}\big{(}R(x)\wedge\mathrm{am}\big{(}p(x,y)\wedge R(y):y:p(x,y)\big{)}:x:q(x)\big{)}.$
Suppose that $n$ is large. For every $a\in(T_{n})$ such that $\mathcal{T}_{n}\models q(a)$ (that is, such that $a$ is a child of the root) the probability, for a random $\mathcal{A}\in\mathbf{W}_{n}$, that roughly half of the children of $a$ satisfy $R(y)$ is close to 1. Hence the probability that
$\mathcal{A}\big{(}\mathrm{am}\big{(}p(a,y)\wedge R(y):y:p(a,y)\big{)}\big{)}\approx 1/2$
is close to 1. With probability close to 1, roughly half of all $a\in T_{n}$ such that $\mathcal{T}_{n}\models q(a)$ will satisfy $R(x)$, so for roughly half of such $a$, the value of $\mathcal{A}\big{(}\mathrm{am}\big{(}p(a,y)\wedge R(y):y:p(a,y)\big{)}\big{)}$ is close to $1/2$, and for the other such $a$ the value is 0. It follows that, with probability tending to 1 as $n\to\infty$, for a random $\mathcal{A}\in\mathbf{W}_{n}$, $\mathcal{A}(\psi)\approx 1/4$. Or more precisely, for all $\varepsilon>0$, if $n$ is large enough then $\mathbb{P}_{n}\big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\big{|}\mathcal{A}(\psi)-1/4\big{|}\leq\varepsilon\big{\}}\big{)}\geq 1-\varepsilon$, no matter if $n$ is even or odd (and Corollary 8.4 tells that we must have such kind of convergence). Hence $\psi$ and $\varphi$ are not asymptotically equivalent. This illustrates that it can matter if we condition an aggregation on a closure type of higher rank than 1, or if we instead “break up” such an aggregation and aggregate several times, each time conditioning the aggregation on a closure type of rank 1 (with respect to the variable that the aggregation binds).
Corollary 8.6.
Let $\tau$, $\sigma$, $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$, $\mathbf{W}_{n}$, $\mathbb{G}$ and $\mathbb{P}_{n}$ be as in Corollary 8.4. Suppose that, for every $R\in\sigma\setminus\tau$, the formula $\theta_{R}$ that $\mathbb{G}$ associates to $R$ is a closure-basic formula over $\mathrm{par}(R)$. Suppose that $\varphi(\bar{x})\in PLA^{*}(\sigma)$ and that if
$F\big{(}\varphi_{1}(\bar{y},\bar{z}),\ldots,\varphi_{m}(\bar{y},\bar{z}):\bar{z}:\chi_{1}(\bar{y},\bar{z}),\ldots,\chi_{m}(\bar{y},\bar{z})\big{)}$
is a subformula of $\varphi(\bar{x})$ then, $F$ is admissible, for all $i=1,\ldots,m$, $\chi_{i}(\bar{y},\bar{z})$ is a $\bar{z}$-positive closure type over $\sigma$ and the $\bar{z}$-rank of $\chi_{i}$ is 0 or 1. Then the conclusions (i) and (ii) of Theorem 8.1 hold.
Proof. This corollary follows by the observations made in the proofs of corollaries 8.2 and 8.4. $\square$
Remark 8.7.
Theorem 8.1 and Corollary 8.4 can not be generalized by replacing ‘continuous’ with ‘admissible’. The reason is as follows. Suppose that $\mathcal{T}_{n}$ is a tree of height 1 such that the root has exactly $n$ children. Let $\sigma=\tau\cup\{R\}$ where $R$ is a binary relation symbol. Let $\alpha\in(0,1)$ be rational. With the kind of $PLA^{*}(\sigma)$-network that is allowed in Theorem 8.1 and Corollary 8.4 we can induce a probability distribution $\mathbb{P}_{n}$ on $\mathbf{W}_{n}$ such that for all different $a,b\in T_{n}$, the probability that $R(a,b)$ holds is $n^{-\alpha}$ independently of what the case is for other pairs. (To get such $\mathbb{P}_{n}$ we can use the aggregation function $\mathrm{length}^{-\alpha}$ which is continuous.) It follows from a result by Shelah and Spencer [26] that there is a first-order sentence $\varphi$ using no other relation symbol than $R$ such that $\mathbb{P}_{n}\big{(}\{\mathcal{A}\in\mathbf{W}_{n}:\mathcal{A}\models\varphi\}\big{)}$ does not converge as $n\to\infty$. Without loss of generality we can assume that all quantifications in $\varphi$ are relativized to the set of children of the root (thus excluding the root from quantifications). Recall that the aggregation functions min and max are admissible. By letting $q(x)$ be a closure type over $\tau$ which expresses “$x$ is a child of the root” and replacing, in $\varphi$, every quantification ‘$\forall x\ldots$’, respectively ‘$\exists x\dots$’, by ‘$\min(\ldots:x:q(x))$’, respectively by ‘$\max(\ldots:x:q(x))$’, we get a 0/1-valued $PLA^{*}(\sigma)$-sentence, say $\varphi^{\prime}$, such that $\mathbb{P}_{n}\big{(}\{\mathcal{A}\in\mathbf{W}_{n}:\mathcal{A}\models\varphi^{\prime}\}\big{)}$ does not converge.
Remark 8.8.
Let $coPLA^{*}(\sigma)$ denote the set of formulas $\varphi(\bar{x})\in PLA^{*}(\sigma)$ such that every subformula of $\varphi$ of the form (8.1) satisfies the conditions stated in Theorem 8.1. In particular, the aggregation functions maximum and minimum may only be used when the closure types $\chi_{i}(\bar{y},\bar{z})$ that we condition on have $\bar{z}$-rank 0. This intuitively means that first-order quantification can only be done over a bounded subset of a tree (a subset which is included in the closure of a tuple of vertices). So $coPLA^{*}(\sigma)$ allows only first-order quantification in a very limited way.
However in some contexts one may be more interested in the proportion (rather than existence) of elements that satisfy some (two-valued) formula. For this we can use the continuous aggregation function am which returns this proportion. If we want to express that the proportion of elements that satisfy a formula is at least as large as the proportion that satisfies another formula, then we could, tentatively, use the connective $\leq$$(x,y)$ which returns 1 if $x\leq y$ and 0 otherwise. But since this connective is not continuous it cannot be used in $PLA^{*}(\sigma)$, so instead we consider a continuous approximation of $\leq$, say $\leq_{\varepsilon}(x,y)$ = 1 if $x\leq y$, 0 if $x\geq y+\varepsilon$, and $(y+\varepsilon-x)/\varepsilon$ if $y<x<y+\varepsilon$, where $\varepsilon>0$ is small. With $\leq_{\varepsilon}$ and am we can, with formulas in $coPLA^{*}$, express approximations of statements about proportions ranging over elements that satisfy a $\bar{z}$-positive closure type over $\sigma$ for some $\bar{z}$. The idea of considering continuous approximations of connectives or aggregation functions can be used more generally to get $coPLA^{*}(\sigma)$-formulas that estimate discrete/discontinuous properties.

### v2

@@
 1. Introduction
-When building a model for probabilistic inference we may not only want to take into account properties and relations which are uncertain and described by (conditional) probabilities, but also “background information” that is certain.
-A common kind of background information is information about a hierarchical structure, and such information can be described by a tree.
-As an example of a tree that represents a hierarchical structure, consider a tree in which the children of the root represent some communities, where the children of a community represent subcommunities, and the children of a subcommunity represent individuals.
-The height of this hierarchy, or tree, is 3 (or 4, depending on our conventions).
-But the number of individuals in even the smallest subcommunity that we consider may be much larger than 3.
-This article considers hierarchical background information represented by a tree such that the height of the tree is very much smaller than the minimal number of children of a nonleaf of the tree.
-On top of the background information, represented by the example tree described above, we may consider information that is given by (conditional) probabilities and is formally represented by (parametrized) 0/1-valued (i.e.
+Logical convergence laws have been proved or disproved in various contexts since the pioneering work of Glebskii et al [13] and, independently, Fagin [12].
+Since finite structures in the sense of first-order logic can represent, for example, relational databases, such convergence laws may have impact outside of mathematics itself.
+When a convergence law applies to a sentence of a formal logic one can use random sampling of structures on a sufficiently large domain to get an estimate of the probability of the sentence on all large enough domains and the estimate will, with high probability, be as close to the actual probability as we like (according to some predetermined error marginal).
+Moreover, sometimes a logical convergence law can be proved (as in [13]) by showing that quantifiers/aggregations can be removed from formulas step by step to produce a new simpler (e.g.
+quantifier-free) formula which is “asymptotically equivalent” to the original one, and the simpler formula can be evaluated in time which is independent of the domain size.
+In theory this gives a deterministic algorithm for estminating the probability of a sentence (or formula with some parameters/constants) which is independent of the domain size.
+In practice, properties of the (generalized) quantifiers or aggregation functions that are involved in the elimination will of course influence the computational complexity.
+Most studies of logical convergence laws have studied contexts in which all relations are modelled probabilistically; intuitively speaking they are “uncertain”.
+When building a model for inference we may not only want to take into account properties and relations which are uncertain, but also “background information” that is certain.
+Background information can have many different forms, but arguably some forms of background information are more common, or more important, than other forms.
+Information that categorizes objects into classes, subclasses and so on is prevalent both in theory and in the physical world.
+A related form of information is that of a hierarchical structure among objects.
+Both forms of information can be represented by trees: the root represents the class of all objects (or the top of the hierarchy), the children of the root the classes of the first subdivision into subclasses (or the next level of the hierarchy), and so on.
+This is a motivation to consider logical convergence laws for random structures that expand a tree, that is, the tree is fixed but in addition we have random relations and the probability that such a relation, say $R(a,b)$, holds may depend on the positions of $a$ and $b$ in the tree.
+As an example, consider a tree in which the children of the root represent some communities, where the children of a community represent subcommunities, and the children of a subcommunity represent individuals.
+On top of the background information, represented by the tree, we may consider information that is given by (conditional) probabilities and is formally represented by (parametrized) 0/1-valued (i.e.
 false/true-valued) random variables.
@@
 So for example, if $R$ is a unary relation symbol then there is a $PLA^{*}$-formula (without free variables) such that its value is the proportion of elements in the domain which satisfy $R(x)$.
-Since the aggregation functions maximum and minimum are available there is a $PLA^{*}$ formula which expresses that at least one (alternatively all) element in the domain satisfies $R(x)$.
 By using the aggregation functions maximum and minimum one can show that every query on finite structures which can be expressed by a first-order formula can be expressed by a $PLA^{*}$-formula (Lemma 3.13).
@@
 For a formula $\varphi(x_{1},\ldots,x_{k})$ of $PLA^{*}$, finite structure $\mathcal{A}$ (with matching signature/vocabulary), and $a_{1},\ldots,a_{k}\in A$, we let $\mathcal{A}(\varphi(a_{1},\ldots,a_{k}))$ denote the value of $\varphi(a_{1},\ldots,a_{k})$ in $\mathcal{A}$.
-The precise way that we will mathematically represent the kind of probabilistic (or uncertain) information described above together with some background information, in the form of a tree, is as follows.
-We let $\Delta$ be a fixed positive integer and we let $\mathcal{T}_{n}$, where $n$ is a positive integer, be a tree with maximum height $\Delta$ that represents the hierarchical background information.
-Since we are interested in what happens for (arbitrarily) large trees where nonleaves have many children, we assume that, for every positive integer $k$, if $n$ is large enough then every nonleaf in $\mathcal{T}_{n}$ has at least $k$ children.
+Since we are interested in large trees we let $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ be a sequence of trees where the number of vertices of $\mathcal{T}_{n}$ tends to infinity as $n\to\infty$.
 Each $\mathcal{T}_{n}$ can be represented by a first-order $\tau$-structure where $\tau=\{E\}$ and $E$ is a binary relation symbol of arity 2.
-Then let $\sigma$ be a finite relational signature that includes $\tau$.
+Let $\sigma$ be a finite relational signature that includes $\tau$.
 In the example above we have $\sigma=\{E,P_{1},P_{2},P_{3},R\}$.
 Then let $\mathbf{W}_{n}$ be the set of all expansions of $\mathcal{T}_{n}$ to $\sigma$, where we think of each such expansion $\mathcal{A}$ as a “possible world”, and let $\mathbb{P}_{n}$ be the probability distribution on $\mathbf{W}_{n}$ induced by a $PLA^{*}(\sigma)$-network $\mathbb{G}$ as defined in Definition 6.9.
-So given a sequence $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ of trees (of the described kind) and a $PLA^{*}(\sigma)$-network $\mathbb{G}$, we have a precise mathematical setting of computing the probability of a query (event) on $\mathbf{W}_{n}$ that can be expressed by a $PLA^{*}$-formula.
-The probability of a query depends on both on the hierarchical information given by $\mathcal{T}_{n}$ and by $\mathbb{G}$.
-However, if we use the brute force way of computing the probability of a query, namely that we compute the probability of every structure in $\mathbf{W}_{n}$ in which the query is true and then add all such probabilities, then the time needed for the computation will be exponential in the number of vertices of the tree $\mathcal{T}_{n}$.
+Now we have a precise mathematical setting in which we can speak about the probability of a query (event) on $\mathbf{W}_{n}$ that can be expressed by a $PLA^{*}$-formula.
+The probability of a query depends (in general) on both the hierarchical information given by $\mathcal{T}_{n}$ and by $\mathbb{G}$.
+If we use the brute force way of computing the probability of a query, namely that we compute the probability of every structure in $\mathbf{W}_{n}$ in which the query is true and then add all such probabilities, then the time needed for the computation will be exponential in the number of vertices of the tree $\mathcal{T}_{n}$.
 In other words, the brute force method of computing the probability of a query does not scale well to large trees $\mathcal{T}_{n}$.
 This motivates the search for other methods of computing, or estimating, the probability of a query in the described context.
+We do not expect to get a general convergence (or nonconvergence) result that covers every kind of sequence $\mathbf{T}$, so additional assumptions on $\mathbf{T}$ will be imposed.
+Koponen’s results in [20] apply to sequences $\mathbf{T}$ where for some $\Delta$ and all $n$ each vertex in $\mathcal{T}_{n}$ has at most $\Delta$ children.
+In this work we instead assume that for some $\Delta$ and all $n$ the height of $\mathcal{T}_{n}$ is at most $\Delta$.
+It will also be assumed that for every $k$ and all large enough $n$, every nonleaf in $\mathcal{T}_{n}$ has at least $k$ children.
+Such trees can (for large $n$) describe hierarchies (categorizations) in with few levels compared to the number of objects of each type in the hierarchy (or few partitions into subcategories compared to the number of objects in each category).
+In the above example the height of the tree, is 3 (or 4, depending on our conventions), but the number of individuals in even the smallest subcommunity that we consider may be much larger than 3.
 In this article we identify
   1. (1)
-    certain conditions on the sequence of trees $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ (the stronger Assumption 6.3 and the milder Assumption 8.3),
+    certain additional conditions on the sequence of trees $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ (the stronger Assumption 6.3 and the milder Assumption 8.3),
   2. (2)
@@
 All three conditions (1)–(3) are necessary for the conclusions described to hold, as we will show.
+Although condition (1) in the form of Assumption 6.3 has a rather technical statement, the general intuition is just that each tree $\mathcal{T}_{n}$ is “sufficiently homogeneous” in the sense that for every nonroot vertex $a$ and subtree $\mathcal{T}^{\prime}$ of $\mathcal{T}_{n}$ rooted in $a$ there are “sufficiently many” siblings $b$ of $a$ such that the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$ is rougly the same as the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $a$ and isomorphic to $\mathcal{T}^{\prime}$.
+For example, suppose that, in a clinical test say, a group of $n$ persons, where $n$ is large, is first divided into “many” groups of roughly the same (large) size, then each group is subdivided into roughly equally many subclasses of roughly the same (still large) size, and so on for, say 5, subdivisions.
+(The divisions may be based on putting people with similar values with respect to various measurements into the same group.) Given that $n$, “many”, and “large” are large enough, the tree $\mathcal{T}_{n}$ representing the subdivisions will be of the kind considered in (1).
+More precisely (according to Assumption 6.3), it suffices that “many” and “large” means “growing faster” than $c\ln n$ for every positive contant $c$.
 The main results, stated in Section 8.1, have the general form described above.
@@
 The essence of these corollaries is (i) that if we consider stronger assumptions on the $PLA^{*}(\sigma)$-network $\mathbb{G}$ (i.e.
-less general probability distributions), then we get results that apply to more queries, and (ii) if we loosen the assumptions on the sequence of trees $\mathbf{T}$ then we get results that apply to less, but still interesting, queries.
+less general probability distributions), then we get results that apply to more queries, and (ii) if we loosen the assumptions on the sequence of trees $\mathbf{T}$ then we get results that apply to fewer, but still interesting, queries.
 Related work.
 Instead of bounding the height of the trees, as we do in this article, one could consider trees with a bound on the number of children that vertices may have.
-Such sequences of trees are covered by the (more general) context considered by Koponen in [20], where probability distributions where defined by $PLA^{*}(\sigma)$-networks and queries by $PLA^{*}$-formulas.
+Such sequences of trees are covered by the (more general) context considered by Koponen in [20], where probability distributions were defined by $PLA^{*}(\sigma)$-networks and queries by $PLA^{*}$-formulas.
 Other work that we are aware of on logical convergence laws, and related issues, in the setting of expansions of “base structures” uses first-order logic (or $L_{\infty,\omega}^{\omega}$) as the query language.
@@
 Lynch [25] formulated a condition ($k$-extendibility) that makes sense for base sequences $\mathbf{B}$ in general and which guarantees that a convergence law holds for first-order logic and the uniform probability distribution.
-Besides this article and [20], other studies that use a a PGM for generating a probability distribution and that prove some result about asymptic probabilities do not consider background information in the form of a sequence of nontrivial base structures.
+Besides this article and [20], other studies that use a PGM for generating a probability distribution and that prove some result about asymptotic probabilities do not consider background information in the form of a sequence of nontrivial base structures.
 But in the context of no background information, when we let $\sigma$ be a finite and relational signature, $\mathbf{W}_{n}$ the set of all $\sigma$-structures with domain $\{1,\ldots,n\}$, $\mathbb{P}_{n}$ a probability distribution induced by a PGM, there are some studies.
@@
 [13] and was used to prove the classical zero-one law for first-order logic, also proved independently by somewhat different methods by Fagin [12].
-Since the pioneering work in [13] and [12] a large number of studies on logical convergence laws and related issues have been conducted in the fields of finite model theory and probabilistic combinatorics, but those studes consider less flexible ways of generating probability distributions than PGMs (most often the uniform distribution).
+Since the pioneering work in [13] and [12] a large number of studies on logical convergence laws and related issues have been conducted in the fields of finite model theory and probabilistic combinatorics, but those studies consider less flexible ways of generating probability distributions than PGMs (most often the uniform distribution).
+The present work and [21, 22, 20] consider similar formalisms for inducing probability and for defining queries/events, so we point out some (unavoidable) differences.
+In [21, 22] all relations are uncertain (probabilistically modelled) and every formula to which the main results(s) applies is asymptotically equivalent to a formula without aggregation functions and without quantifiers.
+In the present work and in [20] which also considers deterministic background information, represented by “base structures”, it is only shown that every formula to which some of the main results applies is asymptotically equivalent to a formula which may use (only) “local” aggregations/quantifications where the notion of “locality” is adapted to the kind of base structures considered, which in the present work are trees.
+It is not possible to remove the use of local aggregations/quantifications in this work (or in [20]) because aggregation/quantifier-free formulas cannot describe the structure of trees (or the base structures considered in [20]).
+Besides this difference compared with [21, 22], another difference is that in the present work (and [20]) we only asymptotically eliminate aggregations that range over elements, or tuples of elements, that satisfy some constraints with respect to the underlying base structure.
+It seems that one can not in general remove this assumption because probabilities may depend on the base structure.
 Organization.
@@
 In section 7 we prove technical results that imply that the assumptions of Theorem 4.8 in Section 4 are satisfied.
-Therefore we can, in the last subsection of Section 7, use Theorem 4.8 and the results from earlier subsections of Section 7 to prove a result about “asymptotic elimination of aggregation functions”.
+Therefore, we can, in the last subsection of Section 7, use Theorem 4.8 and the results from earlier subsections of Section 7 to prove a result about “asymptotic elimination of aggregation functions”.
 In Section 8 we use induction on the height of the $PLA^{*}$-network and the results from Section 7 to prove our main results.
@@
 Let $X\subseteq\Omega$ and $Y\subseteq\Omega$ be measurable.
-Also suppose that $Y=Y_{1}\cup\ldots\cup Y_{k}$, $Y_{i}\cap Y_{j}$ if $i\neq j$, and that each $Y_{i}$ is measurable.
+Also suppose that $Y=Y_{1}\cup\ldots\cup Y_{k}$, $Y_{i}\cap Y_{j}=\emptyset$ if $i\neq j$, and that each $Y_{i}$ is measurable.
 If $\alpha\in[0,1]$, $\varepsilon>0$, and $\mathbb{P}(X\ |\ Y_{i})\in[\alpha-\varepsilon,\alpha+\varepsilon]$ for all $i=1,\ldots,k$, then $\mathbb{P}(X\ |\ Y)\in[\alpha-\varepsilon,\alpha+\varepsilon]$.
 3. Probability logic with aggregation functions
-We consider a logic that we call probability logic with aggregation functions, or $PLA^{*}$, where $PLA^{*}$ is a more general version of $PLA$ and $PLA^{+}$ which where considered in [21] and [22], respectively, and of the probability logic considered by Jaeger in [14, 15].
+We consider a logic that we call probability logic with aggregation functions, or $PLA^{*}$, where $PLA^{*}$ is a more general version of $PLA$ and $PLA^{+}$ which were considered in [21] and [22], respectively, and of the probability logic considered by Jaeger in [14, 15].
 $PLA^{*}$ is a logic with (truth) values in the unit interval $[0,1]$.
@@
 We exemplify what can be expressed with $PLA^{*}(\sigma)$, provided that it contains a binary relation symbol, with the notion of PageRank [6].
-The PageRank of an internet site can be approximated in “stages” as follows (if we supress the “damping factor” for simplicity), where $IN_{x}$ is the set of sites that link to $x$, and $OUT_{y}$ is the set of sites that $y$ link to:
+The PageRank of an internet site can be approximated in “stages” as follows (if we supress the “damping factor” for simplicity), where $IN_{x}$ is the set of sites that link to $x$, and $OUT_{y}$ is the set of sites that $y$ links to:
 $PR_{0}(x)=1/N\text{ where $N$ is the number of sites,}$,
@@
 (Using $\exists$ and $\forall$ as abbreviations) Due to Lemma 3.13, if $\varphi(\bar{x},\bar{y})\in PLA^{*}(\sigma)$ is a 0/1-valued formula then we will often write ‘$\exists\bar{y}\varphi(\bar{x},\bar{y})$’ to mean the same as ‘$\max(\varphi(\bar{x},\bar{y}):\bar{y}:\top)$’, and ‘$\forall\bar{y}\varphi(\bar{x},\bar{y})$’ to mean the same as ‘$\min(\varphi(\bar{x},\bar{y}):\bar{y}:\top)$’.
-The next basic lemma has analogue in first-order logic and is proved straightforwardly by induction on the complexity of $PLA^{*}$-formulas.
+The next basic lemma has an analog in first-order logic and is proved straightforwardly by induction on the complexity of $PLA^{*}$-formulas.
 Lemma 3.15.
@@
 In all of this section let $D_{n}$, $n\in\mathbb{N}^{+}$, be finite sets such that $\lim_{n\to\infty}|D_{n}|=\infty$, and let $\mathbf{W}_{n}$ be a set of $\sigma$-structures with domain $D_{n}$.
-We will describe a method for “asymptotically eliminating aggregation functions” if
-  • the aggregation functions in question are “sufficiently” continuous, and
-  • if there are 0/1-valued formulas with certain properties, including that the proportions of some definable sets “stabilize” as the domain size tends to infinity.
-This method is studied in more detail by Koponen and Weitkämper in [23] and the definitions below and Theorem 4.8 come from [23].
-Theorem 4.8 will be used later (in Section 7.5) to prove our main results.
+We will describe a method for “asymptotically eliminating aggregation functions” which is studied in more detail by Koponen and Weitkämper in [23].
+The definitions below and Theorem 4.8 come from [23] where Theorem 4.8 will be used later (in Section 7.5) to prove our main results.
 Definition 4.1.
@@
 We say that $\varphi(\bar{x})$ and $\psi(\bar{x})$ are asymptotically equivalent (with respect to $\mathbb{P}$) if for all $\varepsilon>0$
-$\lim_{n\to\infty}\mathbb{P}_{n}\Big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\text{ for all $\bar{a}\in(B_{n})^{|\bar{x}|}$, $|\mathcal{A}(\varphi(\bar{a}))-\mathcal{A}(\psi(\bar{a}))|\leq\varepsilon$}\big{\}}\Big{)}\ =\ 1.$
+$\lim_{n\to\infty}\mathbb{P}_{n}\Big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:\text{ for all $\bar{a}\in(D_{n})^{|\bar{x}|}$, $|\mathcal{A}(\varphi(\bar{a}))-\mathcal{A}(\psi(\bar{a}))|\leq\varepsilon$}\big{\}}\Big{)}\ =\ 1.$
 For the rest of this section we fix a sequence $\mathbb{P}=(\mathbb{P}_{n}:n\in\mathbb{N}^{+})$ of probability distributions.
-To define the notions of continuity that we will use we need the notion of convergence testing sequence.
-It generalizes a similar notion used by Jaeger in [14].
-The intuition is that a sequence $\bar{p}_{n}\in[0,1]^{<\omega}$, $n\in\mathbb{N}^{+}$, is convergence testing for parameters $c_{1},\ldots,c_{k}\in[0,1]$ and $\alpha_{1},\ldots\alpha_{k}\in[0,1]$ if the length of $\bar{p}_{n}$ tends to infinity as $n\to\infty$ and, as $n\to\infty$, all entries in $\bar{p}_{n}$ congregate ever closer to the “convergence points” in $\{c_{1},\ldots,c_{k}\}$, and the proportion of entries in $\bar{p}$ which congregate to $c_{i}$ converges to $\alpha_{i}$.
+To define the notions of continuity that we will use we need the notion of convergence testing sequence which generalizes a similar notion used by Jaeger in [14].
 Definition 4.3.
@@
 (i) We call $F$ continuous (or strongly admissible) if $F$ is ct-continuous with respect to every choice of parameters $c_{i,j},\alpha_{i,j}\in[0,1]$, $i=1,\ldots,m$ and $j=1,\ldots,k_{i}$ (for arbitrary $m$ and $k_{i}$).
-(ii) We call $F$ admissible if $F$ is ct-continuous with respect to every choice of parameters $c_{i,j},\alpha_{i,j}\in[0,1]$, $i=1,\ldots,m$ and $j=1,\ldots,k_{i}$ (for arbitrary $m$ and $k_{i}$) such that $\alpha_{i,j}>0$ for all $i$ and $j$.
+(ii) We call $F$ admissible if $F$ is ct-continuous with respect to every choice of parameters $c_{i,j},\alpha_{i,j}\in[0,1]$, $i=1,\ldots,m$ and $j=1,\ldots,k_{i}$ (for arbitrary $k_{i}$) such that $\alpha_{i,j}>0$ for all $i$ and $j$.
 Example 4.6.
+The aggregation functions am, gm and $\mathrm{length}^{-\beta}$ are continuous, which is proved in [21] in the case of am and gm.
+In the case of $\mathrm{length}^{-\beta}$ the claim is easy to prove.
+The aggregation functions max and min are admissible (which is proved in [21]) but not continuous.
+To see that max is not continuous, consider, for $n\in\mathbb{N}$, $\bar{p}_{n}$ and $\bar{q}_{n}$, both of length $n+1$, where all entries of $\bar{p}_{n}$ are 0, the first entry of $\bar{q}_{n}$ is 1 and the rest of the entries are 0.
+Let $c_{1}=0$, $c_{2}=1$, $\alpha_{1}=1$, and $\alpha_{2}=0$.
+It is straightforward to verify that both $(\bar{p}_{n}:n\in\mathbb{N})$ and $(\bar{q}_{n}:n\in\mathbb{N})$ are convergence testing with parameters $c_{1},c_{2}$ and $\alpha_{1},\alpha_{2}$.
+But clearly $\max(\bar{p}_{n})=0$ and $\max(\bar{q}_{n})=1$ for all $n$, so $|\max(\bar{p}_{n})-\max(\bar{q}_{n})|$ does not tend to 0 as $n\to\infty$.
 […]
-The method that we consider for asymptotically eliminating continuous or admissible aggregation functions can be applied if one can find sets $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ of 0/1-valued formulas that satisfy the conditions of the following assumption:
+The method that we consider for asymptotically eliminating continuous or admissible aggregation functions can be applied if one can find sets $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ of 0/1-valued formulas that satisfy the conditions of Assumption 4.7 below (which comes from [23]).
+The intuition behind the technical part (2) of the assumption is that for the sets $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ and every $\varphi(\bar{x},\bar{y})\in L_{0}$ there is a set $L_{\varphi(\bar{x},\bar{y})}\subseteq L_{1}$ of formulas defining some “allowed” conditions (with respect to $\varphi(\bar{x},\bar{y})$) and there are some $\varphi^{\prime}_{1}(\bar{x}),\ldots,\varphi^{\prime}_{s}(\bar{x})\in L_{0}$ such that if $\mathcal{A}\models\varphi^{\prime}_{i}(\bar{a})$ and $\chi(\bar{x},\bar{y})\in L_{\varphi(\bar{x},\bar{y})}$, then the fraction $|\varphi(\bar{a},\mathcal{A})\cap\chi(\bar{a},\mathcal{A})|/|\chi(\bar{a},\mathcal{A})|$ is with high probability close to a number $\alpha_{i}$ that depends only on $\varphi(\bar{x},\bar{y}),\chi(\bar{x},\bar{y}),\varphi_{i}(\bar{x})$ and the sequence of probability distributions $\mathbb{P}$.
+As we allow aggregation functions with arity $m>1$, part (2) needs to simultaneously speak of a sequence $\varphi_{1}(\bar{x},\bar{y}),\ldots,\varphi_{m}(\bar{x},\bar{y})\in L_{0}$.
 Assumption 4.7.
@@
 (i) If $F$ is continuous then $\varphi(\bar{x})$ is asymptotically equivalent to an $L_{0}$-basic formula with respect to $\mathbb{P}$.
-(ii) Suppose, in addition, that if $\varphi_{j}(\bar{x},\bar{y})$, $\chi_{j}(\bar{x},\bar{y})$, $\varphi^{\prime}_{i}(\bar{x})$, $\mathbf{Y}^{\varepsilon}_{n}$ and $\alpha_{i,j}$ are like in part 2 of Asumption 4.7 and $\alpha_{i,j}=0$, then, for all sufficiently large $n$, all $\bar{a}\in(D_{n})^{|\bar{x}|}$ and all $\mathcal{A}\in\mathbf{Y}^{\varepsilon}_{n}$, if $\mathcal{A}\models\varphi^{\prime}_{i}(\bar{a})$ then $\varphi_{j}(\bar{a},\mathcal{A})\cap\chi_{j}(\bar{a},\mathcal{A})=\emptyset$.
-If $F$ is admissible then $\varphi(\bar{x})$ is asymptotically equivalent to an $L_{0}$-basic formula with respect to $\mathbb{P}$.
+(ii) Suppose, in addition, that the following holds if $\varphi_{j}(\bar{x},\bar{y})\in L_{0}$, $\chi_{j}(\bar{x},\bar{y})\in L_{1}$, $\varphi^{\prime}_{i}(\bar{x})\in L_{0}$, $\mathbf{Y}^{\varepsilon}_{n}$ and $\alpha_{i,j}$ are like in part 2 of Asumption 4.7: If $\alpha_{i,j}=0$ then, for all sufficiently large $n$, all $\bar{a}\in(D_{n})^{|\bar{x}|}$ and all $\mathcal{A}\in\mathbf{Y}^{\varepsilon}_{n}$, if $\mathcal{A}\models\varphi^{\prime}_{i}(\bar{a})$ then $\varphi_{j}(\bar{a},\mathcal{A})\cap\chi_{j}(\bar{a},\mathcal{A})=\emptyset$.
+Then it follows that if $F$ is admissible then $\varphi(\bar{x})$ is asymptotically equivalent to an $L_{0}$-basic formula with respect to $\mathbb{P}$.
 5. Directed acyclic graphs, trees and closure types
-For the rest of this article we let $\tau=\{E\}$ where $E$ is a binary relation symbol.
-Moreover, $\sigma$ will always denote a finite relational signature such that $\tau\subseteq\sigma$, although this later assumption will be repeated.
+For the rest of this article we let $\tau=\{E\}$ where $E$ is a binary relation symbol. Moreover, $\sigma$ will always denote a finite relational signature such that $\tau\subseteq\sigma$.
 Definition 5.1.
@@
 (iii) Let $a,b\in G$.
-A directed path (of length $l$) from $a$ to $b$ is a sequence $c_{0},c_{1},\ldots,c_{l}\in T$ such that $l\geq 1$, $\mathcal{T}\models E(c_{i},c_{i+1})$ for all $i=0,\ldots,l-1$, $a=c_{0}$ and $b=c_{l}$.
+A directed path (of length $l$) from $a$ to $b$ is a sequence $c_{0},c_{1},\ldots,c_{l}\in T$ such that $l\geq 1$, $\mathcal{G}\models E(c_{i},c_{i+1})$ for all $i=0,\ldots,l-1$, $a=c_{0}$ and $b=c_{l}$.
 If there is a directed path from $a$ to $b$ then $a$ is called an ancestor of $b$ and $b$ is called a successor of $a$.
@@
   2. (2)
-    there is a unique element $a\in T$, called the root, such that $\mathcal{T}\models\neg\exists xE(x,a)$, and
+    there is a unique element $a\in T$, called the root, such that $\mathcal{T}\models\neg\exists xE(x,a)$,
   3. (3)
-    for all $a\in T$, if $a$ is not the root then there is a unique $b\in T$, called the parent of $a$, such that $\mathcal{T}\models E(b,a)$, and in this case $a$ is called a child of $b$.
+    for all $a\in T$, if $a$ is not the root then there is a unique $b\in T$, called the parent of $a$, such that $\mathcal{T}\models E(b,a)$, and in this case $a$ is called a child of $b$, and
+  4. (4)
+    there do not exist $n\in\mathbb{N}^{+}$ and distinct $a_{0},\ldots,a_{n}\in T$ such that $\mathcal{T}\models E(a_{n},a_{0})$ and $\mathcal{T}\models E(a_{k},a_{k+1})$ for all $k=0,\ldots,n-1$.
 Note that every tree is a DAG.
@@
 Then either $a=b_{i}$ for some $i$, or for some $l\in\mathbb{N}^{+}$ and $i$ there is a directed path from $a$ to some $b_{i}$ of length $l$.
-The statement “there is a directed path from $x$ to $y$ of length $l$” can be expressed by a first-order formula, say $\xi_{l}(x,y)$.
-We will use the expression ‘$x\in\mathrm{cl}(y)$’ to denote a formula $\xi_{l}(x,y)$ for some unspecified $l$, and more generally, the expression ‘$x\in\mathrm{cl}(y_{1},\ldots,y_{k})$’ means $\bigvee_{i=1}^{k}x\in\mathrm{cl}(y_{i})$.
+The statement “there is a directed path from $x$ to $y$ of length at most $l$” can be expressed by a first-order formula, say $\xi_{l}(x,y)$.
+It follows that in all trees with height at most $l$ the property ‘$x\in\mathrm{cl}(y)$’ is expressed by the formula $\xi_{l}(x,y)$.
 Finally, observe that if $\mathcal{T}$ is a tree and $B\subseteq T$ is closed, then the substructure $\mathcal{T}\negthickspace\upharpoonright\negthickspace B$ is also a tree.
+For the rest of this section we fix some $\Delta\in\mathbb{N}^{+}$ and we assume that all trees mentioned have height at most $\Delta$, so “tree” will mean “tree of height at most $\Delta$”.
+With this assumption the property ‘$x\in\mathrm{cl}(y)$’ is expressed by the formula $\xi_{\Delta}(x,y)$ from Remark 5.6 and we will use the more intuitive expression ‘$x\in\mathrm{cl}(y)$’ to denote that formula.
+More generally, the expression ‘$x\in\mathrm{cl}(y_{1},\ldots,y_{k})$’ will mean ‘$\bigvee_{i=1}^{k}x\in\mathrm{cl}(y_{i})$’.
 Definition 5.7.
@@
     for all different $i$ and $j$ in $\{1,\ldots,k\}$, $x_{i}\neq x_{j}$ is a conjunct of $\varphi(x_{1},\ldots,x_{k})$.
-If, moreover, for every $R\in\sigma$, or arity $r$ say, and all $i_{1},\ldots,i_{r}\in\{1,\ldots,k\}$, either $R(x_{i_{1}},\ldots,x_{i_{r}})$ or $\neg R(x_{i_{1}},\ldots,x_{i_{r}})$ is a conjunct of $\varphi(x_{1},\ldots,x_{k})$, then we call $\varphi(x_{1},\ldots,x_{k})$ a complete atomic type over $\sigma$ with respect to $\mathcal{A}$.
+If, moreover, for every $R\in\sigma$, of arity $r$ say, and all $i_{1},\ldots,i_{r}\in\{1,\ldots,k\}$, either $R(x_{i_{1}},\ldots,x_{i_{r}})$ or $\neg R(x_{i_{1}},\ldots,x_{i_{r}})$ is a conjunct of $\varphi(x_{1},\ldots,x_{k})$, then we call $\varphi(x_{1},\ldots,x_{k})$ a complete atomic type over $\sigma$ with respect to $\mathcal{A}$.
 If the particular tree $\mathcal{T}$ is not important in the context we may omit the phrase ‘with respect to $\mathcal{T}$’.
@@
 (i) The restriction of $p(\bar{x})$ to $\sigma^{\prime}$, denoted $p\negthickspace\upharpoonright\negthickspace\sigma^{\prime}$, is a closure-type $p^{\prime}(\bar{x})$ over $\sigma^{\prime}$ (with respect to $\mathcal{T}$) such that $p(\bar{x})\models p^{\prime}(\bar{x})$, and for every closure-type $p^{*}(\bar{x})$ over $\sigma^{\prime}$, if $p(\bar{x})\models p^{*}(\bar{x})$ then $p^{\prime}(\bar{x})\models p^{*}(\bar{x})$.
+(The restriction is not syntactically unique, but it is unique up to logical equivalence, so technically we just choose one of the formulas that satisfy the condition of being a restriction.)
 (ii) The restriction of $p(\bar{x})$ to $\bar{y}$, denoted $p\negthickspace\upharpoonright\negthickspace\bar{y}$, is a closure type over $\sigma$ (with respect to $\mathcal{T}$) $p^{\prime}(\bar{y})$ such that $p(\bar{x})\models p^{\prime}(\bar{y})$, and for every closure-type $p^{*}(\bar{y})$ over $\sigma$, if $p(\bar{x})\models p^{*}(\bar{y})$ then $p^{\prime}(\bar{y})\models p^{*}(\bar{y})$.
-It is not hard to see that, under the same assumptions as in the above definition, if $p(\bar{x})$ is a closure type over $\sigma$ with respect to the tree $\mathcal{T}$, then $p\negthickspace\upharpoonright\negthickspace\sigma^{\prime}$ is a closure-type over $\sigma^{\prime}$ with respect to $\mathcal{T}$, and $p\negthickspace\upharpoonright\negthickspace\bar{y}$ are closure type over $\sigma$ with respect to $\mathcal{T}$.
+(Again, the restriction is unique up to logical equivalence.)
+It is not hard to see that, under the same assumptions as in the above definition, if $p(\bar{x})$ is a closure type over $\sigma$ with respect to the tree $\mathcal{T}$, then $p\negthickspace\upharpoonright\negthickspace\sigma^{\prime}$ is a closure-type over $\sigma^{\prime}$ with respect to $\mathcal{T}$, and $p\negthickspace\upharpoonright\negthickspace\bar{y}$ is a closure type over $\sigma$ with respect to $\mathcal{T}$.
 It will be convenient to use ‘$\models_{tree}$’ to denote consequence restricted to expansions of trees, as defined below.
 Definition 5.10.
-Suppose that $\tau\subseteq\sigma$ and let $\varphi(\bar{x}),\psi(\bar{x})\in PLA^{*}(\sigma)$ are $0/1$-valued formulas.
-By the expression $\varphi(\bar{x})\models_{tree}\psi(\bar{x})$ we mean that if $\mathcal{A}$ is a $\sigma$-structure that expands a tree, $\bar{a}\in(A)^{|\bar{x}|}$, and $\mathcal{A}\models\varphi(\bar{a})$, then $\mathcal{A}\models\psi(\bar{a})$.
+Suppose that $\tau\subseteq\sigma$ and let $\varphi(\bar{x}),\psi(\bar{x})\in PLA^{*}(\sigma)$ be $0/1$-valued formulas.
+By the expression $\varphi(\bar{x})\models_{tree}\psi(\bar{x})$ we mean that if $\mathcal{A}$ is a $\sigma$-structure that expands a tree (of height at most $\Delta$, $\bar{a}\in(A)^{|\bar{x}|}$, and $\mathcal{A}\models\varphi(\bar{a})$, then $\mathcal{A}\models\psi(\bar{a})$.
 The expression $\models_{tree}\psi(\bar{x})$ means the same as $\top\models_{tree}\psi(\bar{x})$, or equivalently, that $\forall\bar{x}\psi(\bar{x})$ is true in all $\sigma$-structures that expand a tree.
@@
 Then, for every $b\in\mathrm{rng}(\bar{b})$, either $b\in\mathrm{rng}(\bar{a})$ or there is a directed path of a particular length, say $l_{b}$, from $b$ to some $a_{b}\in\mathrm{rng}(\bar{a})$.
-In the later case, since $\mathcal{A}\negthickspace\upharpoonright\negthickspace\tau$ is a tree, it follows that $b$ is the unique element such that there is a directed path from $b$ to $a_{b}$ of length $l_{b}$.
+In the latter case, since $\mathcal{A}\negthickspace\upharpoonright\negthickspace\tau$ is a tree, it follows that $b$ is the unique element such that there is a directed path from $b$ to $a_{b}$ of length $l_{b}$.
 Therefore $\bar{b}$ is the unique tuple such that $\mathcal{A}\models p^{*}(\bar{a},\bar{b})$.
@@
 Lemma 5.13.
-Suppose that $\tau\subseteq\sigma$ and $p(\bar{x})$ is a closure-type over $\sigma$ where $\bar{x}=(x_{1},\ldots,k)$.
+Suppose that $\tau\subseteq\sigma$ and $p(\bar{x})$ is a closure-type over $\sigma$ where $\bar{x}=(x_{1},\ldots,x_{k})$.
 (i) For all distinct $i,j\in\{1,\ldots,k\}$ either $p(\bar{x})\models_{tree}$$x_{j}$ is an ancestor of $x_{i}$”, or $p(\bar{x})\models_{tree}$$x_{j}$ is not an ancestor of $x_{i}$”.
@@
 We call $p(\bar{x},\bar{y})$ $\bar{y}$-independent if, for all $i=1,\ldots,k$, $p(\bar{x},\bar{y})\models_{tree}y_{i}\notin\mathrm{cl}(\bar{x})$.
-In the next lemma, and later, the expression ‘$\mathrm{rng}(\bar{y})\subseteq\mathrm{cl}(\bar{x})$’ means that ‘for all $y_{i}\in\mathrm{rng}(\bar{y})$, $y_{i}\in\mathrm{cl}(\bar{x})$’.
+In the next lemma, and later, if $\bar{y}=(y_{1},\ldots,y_{k})$ the expression ‘$\mathrm{rng}(\bar{y})\subseteq\mathrm{cl}(\bar{x})$’ denotes the formula ‘$\bigwedge_{i=1}^{k}y_{i}\in\mathrm{cl}(\bar{x})$’.
 Lemma 5.15.
@@
 Definition 5.21.
-Let $\tau\subseteq\sigma$. A closure-basic formula over $\sigma$ is a formula of the form
-$\bigwedge_{i=1}^{k}(\varphi_{i}(\bar{x})\rightarrow c_{i})$ where for each $i=1,\ldots,k$, $\varphi_{i}(\bar{x})$ is a complete closure type over $\sigma$ and $c_{i}\in[0,1]$.
+Let $\tau\subseteq\sigma$.
+A closure-basic formula over $\sigma$ is a formula of the form $\bigwedge_{i=1}^{k}(\varphi_{i}(\bar{x})\rightarrow c_{i})$ where for each $i=1,\ldots,k$, $\varphi_{i}(\bar{x})$ is a complete closure type over $\sigma$ and $c_{i}\in[0,1]$.
 At one point of the proof of the main results we will use induction on the complexity on formulas, and the base case of the induction uses the following result.
@@
 (i) Suppose that $\varphi_{i}(\bar{x})$, $i=1,\ldots,k$, is a closure-basic formula over $\sigma$.
-Let $q_{1}(\bar{x}),\ldots,q_{m}(\bar{x})$ enumerate, up to logical equivalence, all complete closure-basic types over $\sigma$ in the free variables $\bar{x}$.
-Suppose that $\mathcal{A}$ is a finite $\sigma$-structure abd $\bar{a}\in A^{|\bar{x}|}$.
+Let $q_{1}(\bar{x}),\ldots,q_{m}(\bar{x})$ enumerate, up to logical equivalence, all complete closure types over $\sigma$ in the free variables $\bar{x}$.
+Suppose that $\mathcal{A}$ is a finite $\sigma$-structure and $\bar{a}\in A^{|\bar{x}|}$.
 Observe that for each $i$ the value $\mathcal{A}(\varphi_{i}(\bar{a}))$ depends only on which $q_{j}(\bar{x})$ the sequence $\bar{a}$ satisfies.
@@
 Then let $d_{j}=\mathsf{C}(c_{1,j},\ldots,c_{k,j})$ for $j=1,\ldots,m$.
-Now $\varphi(\bar{x})$ is equivalent in to the closure-basic formula $\bigwedge_{j=1}^{m}(q_{j}(\bar{x})\to d_{j})$.
+Now $\varphi(\bar{x})$ is equivalent to the closure-basic formula $\bigwedge_{j=1}^{m}(q_{j}(\bar{x})\to d_{j})$.
 (ii) Let $\varphi(\bar{x})\in PLA^{*}(\sigma)$ be aggregation-free.
@@
 We consider a base sequence $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ where each $\mathcal{T}_{n}$ is a tree, so in particular $\mathcal{T}_{n}$ is a $\tau$-structure, and $|T_{n}|\to\infty$ as $n\to\infty$.
-Then $\mathbf{W}_{n}$ will be the set of all expansions of $\mathcal{T}_{n}$ to $\sigma$, and a probability distribution $\mathbb{P}_{n}$ will be defined on $\mathbf{W}_{n}$, via the concept of $PLA^{*}(\sigma)$-network.
+Then $\mathbf{W}_{n}$ will be the set of all expansions of $\mathcal{T}_{n}$ to $\sigma$, and a probability distribution $\mathbb{P}_{n}$ will be defined on $\mathbf{W}_{n}$ via a $PLA^{*}(\sigma)$-network.
 Our goal is to identify some constraints on $\mathbf{T}$ and on the $PLA^{*}(\sigma)$-network such that under these constraints we can prove results about asymptotic equivalence of complex formulas to simpler (closure basic) formulas, and derive convergence results.
@@
 So, in addition, we will assume that there is an integer $\Delta\geq 1$ such that, for all $n$, the height of $\mathcal{T}_{n}$ is at most $\Delta$.
-We justify this condition by the fact that such trees can represent hierarchies consisting of a fixed number of levels and where each “group” of objects in such a hierarchy has many more members than the number of levels.
-Moreover, such hierarchies are present in many situations and may influence relations which are modelled/expressed by probabilistic methods.
 The perhaps simplest sequence $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ for which the main results of this article hold is obtained by fixing some integer $\Delta\geq 1$ and letting $\mathcal{T}_{n}$ be a tree of height $\Delta$ such that every leaf is on level $\Delta$ and all nonleaves have exactly $n$ children.
-But in order to increase the applicability of our results we wish prove results for more general sequences $\mathbf{T}$.
+But in order to increase the applicability, we prove results for more general sequences $\mathbf{T}$.
 For example, we want to allow different elements (that is, vertices) of $\mathcal{T}_{n}$ to have different numbers of children.
@@
 By Lemma 2.1 and since $|T_{n}|$ is bounded by a polynomial in $n$ we can argue (somewhat informally) like this.
-Let $a\in T_{n}$ be a child of the root such that such that $a$ is a child of the root with a child, hence $n$ children.
+Let $a\in T_{n}$ be a child of the root such that $a$ is a child of the root with a child, hence $n$ children.
 With high probability (approaching 1 as $n\to\infty$), if $\mathcal{A}\in\mathbf{W}_{n}$ is chosen at random and if we condition on $\mathcal{A}\models R(a)$, then $\mathcal{A}(\varphi(a))\approx 1/2$.
@@
 For odd $n\in\mathbb{N}^{+}$ let $\mathcal{T}_{n}$ be a tree of height $2$ such that the root has $n$ children, one child of the root has $2n^{3}$ children, and all other children of the root have $n$ children.
-For even $n\in\mathbb{N}^{+}$ let $\mathcal{T}_{n}$ be a tree of height $2$ such that the root has $n$ children, two children of the root has $n^{3}$ children, and all other children of the root have $n$ children.
+For even $n\in\mathbb{N}^{+}$ let $\mathcal{T}_{n}$ be a tree of height $2$ such that the root has $n$ children, two children of the root have $n^{3}$ children, and all other children of the root have $n$ children.
 Let $\mathbf{W}_{n}$ be the set of all expansions of $\mathcal{T}_{n}$ to $\sigma$ and let $\mathbb{P}_{n}$ be the uniform probability distribution on $\mathbf{W}_{n}$.
@@
 But we actually need, for every fixed tree $\mathcal{T}^{\prime}$, to put a (not too large) lower bound on the number of siblings $a^{\prime}$ of (any vertex) $a$ such that $\mathsf{N}_{\mathcal{T}_{n}}(a^{\prime},\mathcal{T}^{\prime})$ is not too different from $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})$, where we recall that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})$ is the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $a$ and isomorphic to $\mathcal{T}^{\prime}$.
-(It is not hard to modify Example 6.2, by considering trees in which all leaves are on level 3, so that it shows that this stronger constraint is necessary.) Also, we will assume that there is a polynomial upper bound to the number of children that any member of $\mathcal{T}_{n}$ can have.
+(It is not hard to modify Example 6.2, by considering trees in which all leaves are on level 3, so that it shows that this stronger constraint is necessary.) Also, we will assume that there is a polynomial upper bound on the number of children that any member of $\mathcal{T}_{n}$ can have.
 This will allow us to use Corollary 2.2 in such a way that the main results follow.
@@
       for every tree $\mathcal{T}^{\prime}$ of height at least 1, if $n$ is sufficently large then the following holds: if $a\in T_{n}$ is not the root and $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})>0$, then $a$ has at least $g_{3}(n)$ siblings $b$ such that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})-g_{2}(n)<\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})+g_{2}(n)$.
+Although the above assumption may look complicated, the requirement on $g_{1},g_{2}$ and $g_{3}$ is just that they grow faster than every logarithm and slower that some polynomial, and that $g_{1}$ grows faster than both $g_{2}$ and $g_{3}$.
+So there are many possible choices of such functions, for example $g_{1}(n)=(\ln n)^{3}$, $g_{2}(n)=\ln n$, and $g_{3}(n)=(\ln n)^{2}$ for all $n>1$.
 From now until and including Section 7 let $\mathbf{T}=(\mathcal{T}_{n}:n\in\mathbb{N}^{+})$ be a sequence of trees that satisfies Assumption 6.3.
@@
 (i) Note that (by Assumption 6.3) $|T_{n}|\leq g_{4}(n)^{\Delta}$ for all $n$, where $g_{4}(x)^{\Delta}$ is a polynomial function.
-(ii) Suppose that $\mathcal{T}^{\prime}$ is a tree of heigt $\geq 1$.
-Let $a\in T_{n}$ and suppose that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})>0$.
-Since every nonleaf of $\mathcal{T}_{n}$ has at least $g_{1}(n)$ children it follows that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})\geq g_{1}(n)$ if $n$ is large enough.
+(ii) Suppose that $\mathcal{T}^{\prime}$ is a tree of height $\geq 1$.
+Let $a\in T_{n}$.
+For all sufficiently large $n$ we have that if $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})>0$, then $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})\geq g_{1}(n)$ as we now demonstrate.
+Suppose that $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})>0$.
+Then there is a subtree $\mathcal{T}^{\prime\prime}$ of $\mathcal{T}_{n}$ that is isomorphic to $\mathcal{T}^{\prime}$ and rooted in $a$.
+Let $b$ be a nonleaf vertex of $\mathcal{T}^{\prime\prime}$ such that all children of $b$ are leaves.
+Let $b_{1},\ldots,b_{m}$ be all children of $b$ in $\mathcal{T}_{n}$, so $m\geq g_{1}(n)$, and let $b^{\prime}_{1},\ldots,b^{\prime}_{k}$ be all children of $b$ in $\mathcal{T}^{\prime\prime}$.
+Assuming that $n$ is large enough, $\{b^{\prime}_{1},\ldots,b^{\prime}_{k}\}$ is a proper subset of $\{b_{1},\ldots,b_{m}\}$.
+For every choice of $B\subseteq\{b_{1},\ldots,b_{m}\}$ such that $|B|=k$ we get a subtree $\mathcal{T}_{B}$ of $\mathcal{T}_{n}$ that is rooted in $a$ and isomorphic to $\mathcal{T}^{\prime\prime}$ (hence to $\mathcal{T}^{\prime}$) by replacing $b^{\prime}_{1},\ldots,b^{\prime}_{k}$ in $\mathcal{T}^{\prime\prime}$ by the vertices in $B$.
+The subtree $\mathcal{T}_{B}$ can also be described as the subtree of $\mathcal{T}_{n}$ generated by $(T^{\prime\prime}\setminus\{b^{\prime}_{1},\ldots,b^{\prime}_{k}\})\cup B$.
+There are $\binom{m}{k}\geq\binom{g_{1}(n)}{k}\geq g_{1}(n)$ choices of $B$ and therefore $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})\geq g_{1}(n)$.
 Moreover, if $n$ is large enough, then whether $\mathsf{N}_{\mathcal{T}_{n}}(a,\mathcal{T}^{\prime})>0$ or not depends only on which level $a$ belongs to, because all leaves of $\mathcal{T}_{n}$ are on level $\Delta$ and all nonleaves have at least $g_{1}(n)$ children where $g_{1}(n)\to\infty$.
@@
 Then the $PLA^{*}(\sigma^{\prime})$-network specified by the induced subgraph of $\mathbb{G}$ with vertex set $\sigma^{\prime}\setminus\tau$ and the associated formulas $\theta_{R}$ for all $R\in\sigma^{\prime}\setminus\tau$ will be called the $PLA^{*}(\sigma^{\prime})$-subnetwork of $\mathbb{G}$ induced by $\sigma^{\prime}$.
-Observe that if $\sigma=\tau$ then the DAG of a $PLA^{*}(\sigma)$-network have an empty vertex set and hence there is no probability formula associated to it.
+Observe that if $\sigma=\tau$ then the DAG of a $PLA^{*}(\sigma)$-network has an empty vertex set and hence there is no probability formula associated to it.
 Example 6.11 gives an example of a $PLA^{*}(\sigma)$-network.
@@
     • if $y$ is on level 3 then
-      • if $y$ belongs to the same subcommunity as $x$, then if $P_{3}(y)$ then $3/4$, else $1/2$, else
-      • the proportion of $y$ on level 3 such that $P_{3}(y)$, divided by $3/4$, else
+      • if $y$ is a sibling of $x$, then if $P_{3}(y)$ then $3/4$, else $1/2$, else
+      • the proportion of $y$ on level 3 such that $P_{3}(y)$, multiplied by $3/4$, else
     • if $y$ is on level 2 then
@@
       • the proportion of $y$ on level 2 such that $P_{2}(y)$ among the $y$ that are not a parent of $x$, else
-    • if $y$ is on level 1 then the proportion of children of $y$ such that $P_{1}(y)$, divided by $3/2$.”
+    • if $y$ is on level 1 then the proportion of children of $y$ such that $P_{1}(y)$, multiplied by $2/3$.”
 7. Convergence, balance, and asymptotic elimination of aggregation functions
@@
 Let $\varphi(\bar{x}),\psi(\bar{x})\in PLA^{*}(\sigma)$.
-It follows from the definition that if there are only finitely many $n$ such that there is $\bar{a}\in(T_{n})^{|\bar{x}|}$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\psi(\bar{a})}\big{)}>0$ then, for every $\alpha$, $(\varphi,\psi)$ converges to $\alpha$ with respect to $\mathbb{G}$.
-It also follows from the definition that if there are infinitely many $n$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\psi(\bar{a})}\big{)}>0$ for some $\bar{a}$, but there are only finitely many $n$ such that there is $\bar{a}\in(T_{n})^{|\bar{x}|}$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\bar{a})\wedge\psi(\bar{a})}\big{)}>0$ then $(\varphi,\psi)$ converges to 0 with respect to $\mathbb{G}$.
+It follows from the definition that if there are only finitely many $n$ such that there is $\bar{a}\in(T_{n})^{|\bar{x}|}$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\psi(\bar{a})}\big{)}>0$ then, for every $\alpha$, $(\varphi,\psi)$ converges to $\alpha$ with respect to $\mathbb{G}$.
+It also follows from the definition that if there are infinitely many $n$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\psi(\bar{a})}\big{)}>0$ for some $\bar{a}$, but there are only finitely many $n$ such that there is $\bar{a}\in(T_{n})^{|\bar{x}|}$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\bar{a})\wedge\psi(\bar{a})}\big{)}>0$ then $(\varphi,\psi)$ converges to 0 with respect to $\mathbb{G}$.
 Therefore, when proving that a pair $(\varphi,\psi)$ converges we will assume that there are infinitely many $n$ such that there is $\bar{a}\in(T_{n})^{|\bar{x}|}$ such that $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{\varphi(\bar{a})\wedge\psi(\bar{a})}\big{)}>0$, where the positive probability implies that for some $\mathcal{A}\in\mathbf{W}_{n}$, $\mathcal{A}\models\varphi(\bar{a})\wedge\psi(\bar{a})$.
@@
 We claim that then $(\varphi,\psi,\chi)$ is 0-balanced with respect to $\mathbb{G}$.
-It suffices to show that, for all sufficiently large $n$, if $\mathcal{A}\in\mathbf{W}_{n}$, $\mathcal{A}\models q(\bar{a})$, and $\mathcal{A}\models\psi(\bar{a},\bar{b})$, then $\mathcal{A}\not\models\varphi(\bar{a},\bar{b})$.
+It suffices to show that, for all sufficiently large $n$, if $\mathcal{A}\in\mathbf{W}_{n}$, $\mathcal{A}\models\chi(\bar{a})$, and $\mathcal{A}\models\psi(\bar{a},\bar{b})$, then $\mathcal{A}\not\models\varphi(\bar{a},\bar{b})$.
 But this is immediate from the assumption that $\varphi\wedge\psi\wedge\chi$ is not cofinally satisfiable.
@@
 Suppose that $B\subseteq r(\bar{a},\mathcal{T}_{n})$, and $|B|\geq g_{3}(n)$.
-As $\lim_{n\to\infty}g_{3}(n)$ this means that $B\neq\emptyset$ if $n$ is large enough.
+As $\lim_{n\to\infty}g_{3}(n)=\infty$ this means that $B\neq\emptyset$ if $n$ is large enough.
 Suppose that $\mathcal{A}\in\mathbf{W}_{n}$, so $\mathcal{A}=\mathcal{T}_{n}$.
@@
 For the rest of Section 7 we assume that the height of the underlying DAG of $\mathbb{G}$ is $\rho+1$ where $\rho\geq-1$. We also let
-$\sigma_{\rho}=\{R\in\sigma:\text{ $R$ is on level $i$ of the underlying graph of $\mathbb{G}$ where $i\leq\rho$}\},$
+$\sigma_{\rho}=\tau\cup\{R\in\sigma\setminus\tau:\text{ $R$ is on level $i$ of the underlying graph of $\mathbb{G}$ where $i\leq\rho$}\},$
 $\mathbf{W}_{n}^{\rho}=\{\mathcal{A}\negthickspace\upharpoonright\negthickspace\sigma_{\rho}:\mathcal{A}\in\mathbf{W}_{n}\},$
@@
 7.3. Convergence
-In this subsection we prove statement (B) above. We do it in three steps.
+In this subsection we prove statement (A) above. We do it in three steps.
 Lemma 7.11.
@@
 Let $p_{\rho}=p\negthickspace\upharpoonright\negthickspace\sigma_{\rho}$, so $p_{\rho}$ is a complete closure type over $\sigma_{\rho}$.
-By Assumption 7.8 (2) there is $\alpha$ such that $(p_{\rho},p_{\tau})$ converges to $\alpha$ with respect to $\mathbb{G}$.
+By Assumption 7.8 (2) there is $\alpha$ such that $(p_{\rho},p_{\tau})$ converges to $\alpha$ with respect to $\mathbb{G}_{\rho}$.
 By Lemma 7.11 there is $\beta$ such that $(p,p_{\rho})$ converges to $\beta$ with respect to $\mathbb{G}$.
@@
 By Corollary 7.13, part (2) of Assumption 7.8 holds if $\sigma_{\rho}$ and $\mathbb{G}_{\rho}$ are replaced by $\sigma$ and $\mathbb{G}$, respectively.
-So the induction step for convergence is completed, that is, claim (B) above is proved.
+So the induction step for convergence is completed, that is, claim (A) above is proved.
 Remark 7.15.
@@
 7.4. Balance
-In this section we prove statement (C), in the first part of Lemma 7.17, but we also prove other results that will be used to prove statement (A).
+In this section we prove statement (B), in the first part of Lemma 7.17, but we also prove other results that will be used to prove statement (C).
 Lemma 7.16.
 Suppose that $p_{\rho}(\bar{x},y)$ is a complete self-contained closure type over $\sigma_{\rho}$ and that $p(\bar{x},y)$ is a self-contained closure type over $\sigma$ such that $p\models p_{\rho}$ (or equivalently, $p_{\rho}$ is equivalent to $p\negthickspace\upharpoonright\negthickspace\sigma_{\rho}$).
-Suppose that $\mathrm{rank}_{y}(p_{\rho})\ (=\mathrm{rank}_{y}(p))=1$ and that for every $R\in\sigma\setminus\sigma_{\rho}$, if $\bar{z}$ is a subsequence of $\bar{x}\bar{y}$ and $p\models R(\bar{z})$ or $p\models\neg R(\bar{z})$, then $\bar{z}$ contains $y$.
-Furthermore, suppose that, for some $n$, $\bar{a}\in(T_{n})^{|\bar{x}|}$, $\mathcal{A}^{\prime}\in\mathbf{W}_{n}^{\rho}$, $P\subseteq p_{\rho}(\bar{a},\mathcal{A}^{\prime})$, and $\mathbb{P}_{n}^{\rho}(\mathcal{A}^{\prime})>0$.
+Suppose that $\mathrm{rank}_{y}(p_{\rho})\ (=\mathrm{rank}_{y}(p))=1$ and that for every $R\in\sigma\setminus\sigma_{\rho}$, if $\bar{z}$ is a subsequence of $\bar{x}y$ and $p\models R(\bar{z})$ or $p\models\neg R(\bar{z})$, then $\bar{z}$ contains $y$.
+Furthermore, suppose that $n\in\mathbb{N}^{+}$, $\bar{a}\in(T_{n})^{|\bar{x}|}$, $\mathcal{A}^{\prime}\in\mathbf{W}_{n}^{\rho}$, $P\subseteq p_{\rho}(\bar{a},\mathcal{A}^{\prime})$ and $\mathbb{P}_{n}^{\rho}(\mathcal{A}^{\prime})>0$.
 There is $\alpha\in[0,1]$, depending only on $p_{\rho}$ and $p$, such that for every $\varepsilon>0$ there is $c>0$, depending only on $\alpha$ and $\varepsilon$, such that if $n$ and $|P|$ are large enough, then
@@
 We adopt all assumptions of the lemma.
-Since $\mathbb{P}_{n}^{\rho}(\mathcal{A}^{\prime})>0$ it follows from Lemma 6.10 that $\mathbb{P}_{n}\big{(}\mathbf{W}^{\mathcal{A}^{\prime}}\big{)}>0$.
+Since $\mathbb{P}_{n}^{\rho}(\mathcal{A}^{\prime})>0$, it follows from Lemma 6.10 that $\mathbb{P}_{n}\big{(}\mathbf{W}^{\mathcal{A}^{\prime}}\big{)}>0$.
 As $\mathcal{A}^{\prime}\models p_{\rho}(\bar{a},b)$ for all $b\in P$, it follows that for all $b\in P$, $\mathbb{P}_{n}\big{(}\mathbf{E}_{n}^{p_{\rho}(\bar{a},b)}\big{)}>0$.
@@
 Remark 7.18.
-By part (i) of Lemma 7.17 we have proved statement (C) above (at the end of Section 7.2).
-Although we have completed the inductive step for part (3) of Assumption 7.8 we will continue to prove results about balanced triples because we need them for proving statement (A) and the main results of this article.
+By part (i) of Lemma 7.17 we have proved statement (B) above (at the end of Section 7.2).
+Although we have completed the inductive step for part (3) of Assumption 7.8 we will continue to prove results about balanced triples because we need them for proving statement (C) and the main results of this article.
 Proposition 7.19.
@@
 Case 2. Suppose that $\bar{v}$ is empty.
-Then $\bar{y}=z\bar{u}$ so $s(\bar{x},z,\bar{u})$ is the same as $p(\bar{x},z,\bar{u})$ (and $p_{\tau}$ is the same as $s_{\tau}$).
+Then $\bar{y}=z\bar{u}$ so $s(\bar{x},z,\bar{u})$ is equivalent to $p(\bar{x},z,\bar{u})$ (and $p_{\tau}$ is equivalent to $s_{\tau}$).
 Since $p_{\tau}(\bar{x},z,\bar{u})$ is self-contained it follows that if $\mathcal{T}_{n}\models p_{\tau}(\bar{a},b,\bar{c})$ then $\mathcal{T}^{\prime}:=\mathcal{T}_{n}\negthickspace\upharpoonright\negthickspace\mathrm{rng}(b\bar{c})$ is a subtree of $\mathcal{T}_{n}$ which is rooted in $b$ and the isomorphism type of $\mathcal{T}^{\prime}$ is determined by $p_{\tau}$ alone.
-It is easy to see that if $\mathcal{T}_{n}\models r_{\tau}(\bar{a},b)$ and $\mathcal{T}^{\prime\prime}$ is a subtree of $\mathcal{T}_{n}$ which is rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$, then the elements of $T^{\prime\prime}\setminus\{b\}$ can be ordered in a unique way as $\bar{c}$ so that $\mathcal{T}_{n}\models p_{\tau}(\bar{a},b,\bar{c})$.
-Therefore there is a bijection between $p_{\tau}(\bar{a},b,\mathcal{T}_{n})$ and the subtrees of $\mathcal{T}_{n}$ that are rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$.
-Recall that $\mathsf{N}_{n}(b,\mathcal{T}^{\prime}):=\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})$ is the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$.
-Now we have $|p_{\tau}(\bar{a},b,\mathcal{T}_{n})|=\mathsf{N}_{n}(b,\mathcal{T}^{\prime})$ if $p_{\tau}(\bar{a},b,\mathcal{T}_{n})\neq\emptyset$.
-Recall that, by Remark 6.4, we have $\mathsf{N}_{n}(b,\mathcal{T}^{\prime})\geq g_{1}(n)$ whenever $n$ is large enough and $\mathsf{N}_{n}(b,\mathcal{T}^{\prime})>0$.
+It is easy to see that if $\mathcal{T}_{n}\models r_{\tau}(\bar{a},b)$ and $\mathcal{T}^{\prime\prime}$ is a subtree of $\mathcal{T}_{n}$ which is rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$, then the elements of $T^{\prime\prime}\setminus\{b\}$ can be ordered as $\bar{c}$ so that $\mathcal{T}_{n}\models p_{\tau}(\bar{a},b,\bar{c})$.
+Also, note that if $\mathcal{T}^{*}$ is a different (from $\mathcal{T}^{\prime\prime}$) subtree of $\mathcal{T}_{n}$ that is rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$, then $T^{\prime\prime}\setminus\{b\}\neq T^{*}\setminus\{b\}$.
+Therefore, $|p_{\tau}(\bar{a},b,\mathcal{T}_{n})|$ is at least as large as the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$.
+Recall that $\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})$ is the number of subtrees of $\mathcal{T}_{n}$ that are rooted in $b$ and isomorphic to $\mathcal{T}^{\prime}$.
+Now we have $|p_{\tau}(\bar{a},b,\mathcal{T}_{n})|\geq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})$ if $p_{\tau}(\bar{a},b,\mathcal{T}_{n})\neq\emptyset$.
+Recall that, by Remark 6.4, we have $\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})\geq g_{1}(n)$ whenever $n$ is large enough and $\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})>0$.
 Suppose that $\bar{x}=(x_{1},\ldots,x_{k})$.
@@
 We now assume that $p_{\tau}(\bar{a},\mathcal{T}_{n})\neq\emptyset$; otherwise there is nothing to prove.
-Then $r_{\tau}(\bar{a},\mathcal{T}_{n})\neq\emptyset$ and for every $b\in r_{\tau}(\bar{a},\mathcal{A})$ we have $|p_{\tau}(\bar{a},b,\mathcal{T}_{n})|=\mathsf{N}_{n}(b,\mathcal{T}^{\prime})\geq g_{1}(n)$ if $n$ is large enough.
+Then $r_{\tau}(\bar{a},\mathcal{T}_{n})\neq\emptyset$ and for every $b\in r_{\tau}(\bar{a},\mathcal{A})$ we have $|p_{\tau}(\bar{a},b,\mathcal{T}_{n})|\geq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})\geq g_{1}(n)$ if $n$ is large enough.
 We will now define a partition $P_{1},\ldots,P_{\mu}$ of $r_{\tau}(\bar{a},\mathcal{T}_{n})$.
@@
   1. (1)
-    Let $\lambda\geq 1$ be the minimal integer such that there is $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $g_{1}(n)+(\lambda-1)g_{2}(n)\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<g_{1}(n)+\lambda g_{2}(n)$.
+    Let $\lambda\geq 1$ be the minimal integer such that there is $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $g_{1}(n)+(\lambda-1)g_{2}(n)\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<g_{1}(n)+\lambda g_{2}(n)$.
     1. (a)
-      If $\lambda=1$ then let $P_{1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{1,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{1,2}$ where $m_{1,1}:=g_{1}(n)$ and $m_{1,2}:=m_{1,1}+(\lambda+3)g_{2}(n)$.
+      If $\lambda=1$ then let $P_{1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{1,1}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{1,2}$ where $m_{1,1}:=g_{1}(n)$ and $m_{1,2}:=m_{1,1}+(\lambda+3)g_{2}(n)$.
     2. (b)
-      If $\lambda>1$ then let $P_{1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{1,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{1,2}$ where $m_{1,1}:=g_{1}(n)+(\lambda-1)g_{2}(n)$ and $m_{1,2}:=m_{1,1}+(\lambda+3)g_{2}(n)$.
+      If $\lambda>1$ then let $P_{1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{1,1}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{1,2}$ where $m_{1,1}:=g_{1}(n)+(\lambda-1)g_{2}(n)$ and $m_{1,2}:=m_{1,1}+(\lambda+3)g_{2}(n)$.
     Assumption 6.3 implies that, in both cases, $|P_{1}|\geq g_{3}(n)$.
   2. (2)
-    Now suppose that the parts $P_{1},\ldots,P_{l}$ have been defined such that $|P_{i}|\geq g_{3}(n)$ for all $i$ and that numbers $m_{i,j}$ (with $j\in\{1,2\}$) have been defined such that $P_{i}$ contains all $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{i,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{i,2}$, and $m_{i+1,1}=m_{i,2}$ if $i<l$.
-    Let $\lambda\geq 1$ be the minimal integer (if it exists) such that there is $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{l,2}+\lambda g_{2}(n)\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{l,2}+(\lambda+1)g_{2}(n)$.
+    Now suppose that the parts $P_{1},\ldots,P_{l}$ have been defined such that $|P_{i}|\geq g_{3}(n)$ for all $i$ and that numbers $m_{i,j}$ (with $j\in\{1,2\}$) have been defined such that $P_{i}$ contains all $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{i,1}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{i,2}$, and $m_{i+1,1}=m_{i,2}$ if $i<l$.
+    Let $\lambda\geq 1$ be the minimal integer (if it exists) such that there is $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{l,2}+\lambda g_{2}(n)\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{l,2}+(\lambda+1)g_{2}(n)$.
     1. (a)
-      If $\lambda=1$ then let $P_{l+1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{l+1,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{l+1,2}$ where $m_{l+1,1}:=m_{l,2}$ and $m_{l+1,2}:=m_{l,2}+3g_{2}(n)$.
+      If $\lambda=1$ then let $P_{l+1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{A})$ such that $m_{l+1,1}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{l+1,2}$ where $m_{l+1,1}:=m_{l,2}$ and $m_{l+1,2}:=m_{l,2}+3g_{2}(n)$.
     2. (b)
@@
       1. (i)
-        let $P_{l+1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{l+1,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{l+1,2}$ where $m_{l+1,1}:=m_{l,2}+(\lambda-1)g_{2}(n)$ and $m_{l+1,2}:=m_{l,2}+(\lambda+2)g_{2}(n)$, and
+        let $P_{l+1}$ contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{l+1,1}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{l+1,2}$ where $m_{l+1,1}:=m_{l,2}+(\lambda-1)g_{2}(n)$ and $m_{l+1,2}:=m_{l,2}+(\lambda+2)g_{2}(n)$, and
       2. (ii)
-        redefine $P_{l}$ so that it contains contain all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{l,1}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{l,2}+g_{2}(n)$ and then redefine $m_{l,2}$ by $m_{l,2}:=m_{l,2}+g_{2}(n)$.
+        redefine $P_{l}$ so that it contains all $b\in r_{\tau}(\bar{a},\mathcal{T}_{n})$ such that $m_{l,1}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{l,2}+g_{2}(n)$ and then redefine $m_{l,2}$ by $m_{l,2}:=m_{l,2}+g_{2}(n)$.
     In both cases we have $|P_{l+1}|\geq g_{3}(n)$ by Assumption 6.3.
 If no $\lambda$ as above exists then $r_{\tau}(\bar{a},\mathcal{T}_{n})=P_{1}\cup\ldots\cup P_{l}$ and we are done and let $\mu=l$.
-Observe that when a part $P_{i}$ was first defined then, for some real number $m_{i}$, all $b\in P_{i}$ satisfy $m_{i}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{i}+3g_{2}(n)$.
-Also, each part $P_{i}$ was redefined at most once and in such a way that (with the same $m_{i}$) all $b\in P_{i}$ satisfy $m_{i}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{i}+4g_{2}(n)$.
+Observe that when a part $P_{i}$ was first defined then, for some real number $m_{i}$, all $b\in P_{i}$ satisfy $m_{i}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{i}+3g_{2}(n)$.
+Also, each part $P_{i}$ was redefined at most once and in such a way that (with the same $m_{i}$) all $b\in P_{i}$ satisfy $m_{i}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{i}+4g_{2}(n)$.
 By the construction we have $m_{i}\geq g_{1}(n)$ for all $i$.
@@
 (7.9) for each part $P_{i}$, $|P_{i}|\geq g_{3}(n)$ and there is $m_{i}\geq g_{1}(n)$ such that
-for all $b\in P_{i}$, $m_{i}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})\leq m_{i}+4g_{2}(n)$.
+for all $b\in P_{i}$, $m_{i}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})\leq m_{i}+4g_{2}(n)$.
 It follows from part (i) of Lemma 7.17 that there is $\alpha$, depending only on $r$, such that for every $\varepsilon>0$ there is $C>0$, depending only on $\alpha$ and $\varepsilon$, such that if $n$ is sufficiently large, then for every part $P_{i}\subseteq r_{\tau}(\bar{a},\mathcal{T}_{n})$,
@@
 $(\alpha-\varepsilon)|P_{i}|\leq|r(\bar{a},\mathcal{A})\cap P_{i}|\leq(\alpha+\varepsilon)|P_{i}|\big{\}}.$
-Recall that (by Assumption 6.3) every element of $\mathcal{T}_{n}$ has at most $f_{4}(n)$ children, where $f_{4}$ is a polynomial.
-Since all members of $r_{\tau}(\bar{a},\mathcal{T}_{n})$ are children of $a_{1}$ it follows that $|r_{\tau}(\bar{a},\mathcal{T}_{n})|\leq f_{4}(n)$.
-As $P_{1},\ldots,P_{\mu}$ is a partition of $r_{\tau}(\bar{a},\mathcal{T}_{n})$ it follows that $\mu\leq f_{4}(n)$.
+Recall that (by Assumption 6.3) every element of $\mathcal{T}_{n}$ has at most $g_{4}(n)$ children, where $g_{4}$ is a polynomial.
+Since all members of $r_{\tau}(\bar{a},\mathcal{T}_{n})$ are children of $a_{1}$ it follows that $|r_{\tau}(\bar{a},\mathcal{T}_{n})|\leq g_{4}(n)$.
+As $P_{1},\ldots,P_{\mu}$ is a partition of $r_{\tau}(\bar{a},\mathcal{T}_{n})$ it follows that $\mu\leq g_{4}(n)$.
 Let
@@
 Fix any $i\in\{1,\ldots,\mu\}$.
-By (7.9) there is a number $m_{i}\geq g_{1}(n)$ such that if $b\in P_{i}$ then $m_{i}\leq\mathsf{N}_{n}(b,\mathcal{T}^{\prime})<m_{i}+4g_{2}(n)$ and hence $m_{i}\leq|p_{\tau}(\bar{a},b,\mathcal{A})|<m_{i}+4g_{2}(n)$.
+By (7.9) there is a number $m_{i}\geq g_{1}(n)$ such that if $b\in P_{i}$ then $m_{i}\leq\mathsf{N}_{\mathcal{T}_{n}}(b,\mathcal{T}^{\prime})<m_{i}+4g_{2}(n)$ and hence $m_{i}\leq|p_{\tau}(\bar{a},b,\mathcal{A})|<m_{i}+4g_{2}(n)$.
 Since $\mathcal{A}\in\mathbf{X}_{n}^{\varepsilon,i}(\bar{a})$ we get, by the use of (7.10),
@@
 Proposition 7.24.
-Let $p(\bar{x},\bar{y})$ and $r(\bar{x},\bar{y})$ be (possibly incomplete) closure types over $\sigma$, where $r$ is $\bar{y}$-positive with respect to $\mathbb{G}$, and let $q(\bar{x})$ be complete closure type over $\sigma$, Then $(p,r,q)$ is balanced with respect to $\mathbb{G}$.
+Let $p(\bar{x},\bar{y})$ and $r(\bar{x},\bar{y})$ be (possibly incomplete) closure types over $\sigma$, where $r$ is $\bar{y}$-positive with respect to $\mathbb{G}$, and let $q(\bar{x})$ be a complete closure type over $\sigma$, Then $(p,r,q)$ is balanced with respect to $\mathbb{G}$.
 Proof.
@@
 If $(p,r)$ is also eventually constant, then it must be eventually constant with value 00.
-This implies that for all sufficiently large $n$ and all $\bar{a}\in(T_{n})^{|\bar{x}|}$, $p(\bar{a},\mathcal{A})\cap r(\bar{a},\mathcal{A})=\emptyset$.
+This implies that for all sufficiently large $n$, all $\bar{a}\in(T_{n})^{|\bar{x}|}$, and all $\mathcal{A}\in\mathbf{W}_{n}$, we have $p(\bar{a},\mathcal{A})\cap r(\bar{a},\mathcal{A})=\emptyset$.
 By involving the conclusions of Remark 7.15 we can now conclude the following, by induction on the height of $\mathbb{G}$: Suppose that for all $R\in\sigma\setminus\tau$, $\theta_{R}$ is a closure basic formula.
@@
 7.5. Asymptotic elimination of aggregation functions
-In this subsection we prove statement (A) above, that is, we prove that given assumptions 6.3 and 7.8 (the induction hypothesis) and the results proved earlier in Section7 we can asymptotically eliminate aggregation functions, provided that some conditions are satisfied.
+In this subsection we prove statement (C) above, that is, we prove that given assumptions 6.3 and 7.8 (the induction hypothesis) and the results proved earlier in Section 7 we can asymptotically eliminate aggregation functions, provided that some conditions are satisfied.
 Some motivation for these conditions are given in the beginning of Section 8.
@@
 (So $L_{p(\bar{x},\bar{y})}$ depends only on the free variables $\bar{x}$ and $\bar{y}$ but not on other properties of $p$.) Suppose that $p_{i}(\bar{x},\bar{y})\in L_{0}$ for $i=1,\ldots,k$ and let $\lambda_{i}(\bar{x},\bar{y})\in L_{p_{i}(\bar{x},\bar{y})}$ for $i=1,\ldots,k$, so $\lambda_{i}(\bar{x},\bar{y})$ is a $\bar{y}$-positive closure type over $\sigma$.
-Let $q_{i}(\bar{x})\in L_{0}$, $i=1,\ldots,s$, enumerate all, up to equivalence, complete closure types over $\sigma$ in the variables $\bar{x}$ and we may assume that $q_{i}(\bar{x})$ is not equivalent to $q_{j}(\bar{x})$ if $i\neq j$ (otherwise just remove some $q_{i}(\bar{x})$).
+Let $q_{i}(\bar{x})\in L_{0}$, $i=1,\ldots,s$, enumerate all, up to equivalence, complete closure types over $\sigma$ in the variables $\bar{x}$.
+We may assume that $q_{i}(\bar{x})$ is not equivalent to $q_{j}(\bar{x})$ if $i\neq j$.
 Observe that by Assumption 6.3, Remark 6.4 and the definition of $\mathbf{W}_{n}$, it follows that if $p(\bar{x},\bar{y})$ and $q(\bar{x})$ are complete closure types over $\sigma$ and $p(\bar{x},\bar{y})\wedge q(\bar{x})$ is satisfied in some $\mathcal{A}\in\mathbf{W}_{n}$ for some $n$, then for all sufficiently large $n$ the same formula is satisfied in some $\mathcal{A}\in\mathbf{W}_{n}$.
@@
 Then Proposition 7.26 implies that part (1) of Assumption 7.8 holds if $\sigma_{\rho}$, $\sigma$ and $\mathbb{G}_{\rho}$ are replaced by $\sigma$, $\sigma^{+}$ and $\mathbb{G}$, respectively.
-Hence statement (A) is proved.
+Hence statement (C) is proved.
 8. The main results
@@
 and (as $\varphi(\bar{x})$ and $\bigwedge_{i=1}^{m}(\varphi_{i}(\bar{x})\to c_{i})$ are asymptotically equivalent)
-$\mathbb{P}_{n}\Big{(}\big{\{}\mathcal{A}\in\mathbf{W}_{n}:|\mathcal{A}(\varphi(\bar{a}))-\mathcal{A}(\bigwedge_{i=1}^{m}(\varphi_{i}(\bar{a})\to c_{i}))|\leq\varepsilon\big{\}}\Big{)}\geq 1-\varepsilon.$
+$\mathbb{P}_{n}\Big{(}\Big{\{}\mathcal{A}\in\mathbf{W}_{n}:\Big{|}\mathcal{A}(\varphi(\bar{a}))-\mathcal{A}\Big{(}\bigwedge_{i=1}^{m}\big{(}\varphi_{i}(\bar{a})\to c_{i}\big{)}\Big{)}\Big{|}\leq\varepsilon\Big{\}}\Big{)}\geq 1-\varepsilon.$
 Note that if $\mathcal{A}\models\varphi_{i}(\bar{a})$ then $\mathcal{A}(\bigwedge_{i=1}^{m}(\varphi_{i}(\bar{a})\to c_{i}))=c_{i}$.
@@
 Therefore Theorem 8.1 only applies to min and max if they are used together with a conditioning closure type $\chi(\bar{y},z)$ with $z$-rank 0.
-The next corollary states that if we strengthen the assumptions on the $PLA(\sigma)$-network $\mathbb{G}$ (thus limiting the range of distributions that such $\mathbb{G}$ can induce), then the conclusions (i) and (ii) of Theorem 8.1 hold if the formula $\varphi(\bar{x})$ uses (only) admissible aggregation functions.
+The next corollary states that if we strengthen the assumptions on the $PLA^{*}(\sigma)$-network $\mathbb{G}$ (thus limiting the range of distributions that such $\mathbb{G}$ can induce), then the conclusions (i) and (ii) of Theorem 8.1 hold if the formula $\varphi(\bar{x})$ uses (only) admissible aggregation functions.
 In particular the conclusions of the theorem hold for first-order formulas.
@@
 In particular Assumption 6.3 is adopted.
-Suppose that, for every $R\in\sigma\setminus\tau$, the formula $\theta_{R}$ that $\mathbb{G}$ associates to $R$ is a closure-basic formula over $\mathrm{par}(R)$ (the parents of $R$ in the underlying DAG of $\mathbb{G}$).
+Suppose that, for every $R\in\sigma\setminus\tau$, the formula $\theta_{R}$ that $\mathbb{G}$ associates to $R$ is a closure-basic formula over $\mathrm{par}(R)$.
 Suppose that $\varphi(\bar{x})\in PLA^{*}(\sigma)$ and that if
@@
 Proof.
-The conditions of Assumption 6.3 that have been removed in Assumption 8.3 are only used in the proof of Proposition 7.19 when proving the inductive step, and in this step we considered a closure type of $\bar{y}$-rank at least 2.
+The conditions of Assumption 6.3 that have been removed in Assumption 8.3 are only used in the proof of Proposition 7.19 when proving the inductive step, and the inductive step is only necessary if we only consider $p_{\tau}(\bar{x},\bar{y})$ in that proposition with $\mathrm{rank}_{\bar{y}}(p_{\tau})\geq 2$.
+As a consequence, all results in Section 7.4 after Proposition 7.19 restricted to closure types $p(\bar{x},\bar{y})$ with $\mathrm{rank}_{\bar{y}}(p)\leq 1$ follow from results before Proposition 7.19.
+Hence, Proposition 7.26 restricted to $\chi_{i}(\bar{y},\bar{z})$ with $\bar{z}$-rank at most 1 does not need Proposition 7.19.
+Consequently, the same holds for Corollary 8.4.
 $\square$
@@
 By letting $q(x)$ be a closure type over $\tau$ which expresses “$x$ is a child of the root” and replacing, in $\varphi$, every quantification ‘$\forall x\ldots$’, respectively ‘$\exists x\dots$’, by ‘$\min(\ldots:x:q(x))$’, respectively by ‘$\max(\ldots:x:q(x))$’, we get a 0/1-valued $PLA^{*}(\sigma)$-sentence, say $\varphi^{\prime}$, such that $\mathbb{P}_{n}\big{(}\{\mathcal{A}\in\mathbf{W}_{n}:\mathcal{A}\models\varphi^{\prime}\}\big{)}$ does not converge.
-Remark 8.8.
-Let $coPLA^{*}(\sigma)$ denote the set of formulas $\varphi(\bar{x})\in PLA^{*}(\sigma)$ such that every subformula of $\varphi$ of the form (8.1) satisfies the conditions stated in Theorem 8.1.
-In particular, the aggregation functions maximum and minimum may only be used when the closure types $\chi_{i}(\bar{y},\bar{z})$ that we condition on have $\bar{z}$-rank 0.
-This intuitively means that first-order quantification can only be done over a bounded subset of a tree (a subset which is included in the closure of a tuple of vertices).
-So $coPLA^{*}(\sigma)$ allows only first-order quantification in a very limited way.
-However in some contexts one may be more interested in the proportion (rather than existence) of elements that satisfy some (two-valued) formula.
-For this we can use the continuous aggregation function am which returns this proportion.
-If we want to express that the proportion of elements that satisfy a formula is at least as large as the proportion that satisfies another formula, then we could, tentatively, use the connective $\leq$$(x,y)$ which returns 1 if $x\leq y$ and 0 otherwise.
-But since this connective is not continuous it cannot be used in $PLA^{*}(\sigma)$, so instead we consider a continuous approximation of $\leq$, say $\leq_{\varepsilon}(x,y)$ = 1 if $x\leq y$, 0 if $x\geq y+\varepsilon$, and $(y+\varepsilon-x)/\varepsilon$ if $y<x<y+\varepsilon$, where $\varepsilon>0$ is small.
-With $\leq_{\varepsilon}$ and am we can, with formulas in $coPLA^{*}$, express approximations of statements about proportions ranging over elements that satisfy a $\bar{z}$-positive closure type over $\sigma$ for some $\bar{z}$.
-The idea of considering continuous approximations of connectives or aggregation functions can be used more generally to get $coPLA^{*}(\sigma)$-formulas that estimate discrete/discontinuous properties.
+Acknowledgements.
+The authors thank the anonymous referees for their insightful remarks and suggestions which have improved this article.
+Vera Koponen was partially supported by the Swedish Research Council, grant 2023-05238_VR.
+Vera Koponen and Yasmin Tousinejad were partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.